### v2

# Global weak solutions for the inverse mean curvature flow in the Heisenberg group

Abstract.
We consider the inverse mean curvature flow (IMCF) in the Heisenberg group $(\mathbb{H}^{n},d_{\varepsilon})$, where $d_{\varepsilon}$ is distance associated to either $|\cdot|_{\varepsilon}$, $\varepsilon>0$, the natural family of left-invariant Riemannian metrics, or with their sub-Riemannian counterparts for $\varepsilon=0$. For $\Omega\subseteq\mathbb{H}^{n}$ an open set with smooth boundary $\Sigma_{0}=\partial\Omega$ satisfying a uniform exterior gauge-ball condition and bounded complement we show existence of a global weak IMCF of generalized hypersurfaces $\{\Sigma^{\varepsilon}_{s}\}_{s\geq 0}\subseteq\mathbb{H}^{n}$ which is proper. Here, both in the Riemannian and in the sub-Riemannian setting, we adopt the weak formulation introduced by Huisken and Ilmanen in [28], following the approach in [34] due to Moser and based on the the link between IMCF and $p$-harmonic functions.
Key words and phrases:
Inverse mean curvature flow, Heisenberg group, $p$-capacitary potential
2020 Mathematics Subject Classification:
Primary 35R03, 53E10; Secondary 35B45, 35B50, 35J92
E.V. is a member of Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM). A.P. is partially funded by PRIN project 2022PJ9EFL “Geometric Measure Theory: Structure of Singular Measures, Regularity Theory and Applications in the Calculus of Variations”. E.V. is partially funded by PRIN project 2022R537CS “$NO^{3}$ Nodal Optimization, NOnlinear elliptic equations, NOnlocal geometric problems, with a focus on regularity” and by the Indam-GNAMPA project “Esistenza, unicità e regolarità di soluzioni per problemi singolari”.
1. Introduction
Let $\mathbb{H}^{n}$ be the $(2n+1)$-dimensional Heisenberg group and let $\Omega\subset\mathbb{H}^{n}$ be an open set with $C^{2}$-smooth boundary $\Sigma_{0}=\partial\Omega$ such that its complement [^1] We will use both notations for complement of sets, depending on the graphical convenience along the text. $\Omega^{c}:=\mathbb{H}^{n}\setminus\Omega$ is bounded. Throughout the paper we assume that $\Omega$ satisfies an exterior uniform gauge-ball condition ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}$ (see Definition 2.1).
In the present paper we are mainly interested in positive solutions to the boundary value problem
(1.1) $\left\{\begin{array}[]{rl}{\rm div}_{0}\left(\dfrac{\nabla_{0}u}{|\nabla_{0}u|_{0}}\right)=|\nabla_{0}u|_{0}&\textrm{in }\Omega,\\ u=0&\textrm{on }\partial\Omega,\end{array}\right.$
[…]
(1.2) $U(g)=(Q-1)\log\frac{\|g_{0}^{-1}*g\|}{R_{0}}\,,$
where $Q=2n+2$ being from now on the homogeneous dimension of $\mathbb{H}^{n}$.
Problem (1.1) can be considered as the natural sub-Riemannian counterpart of the level set formulation of the classical Inverse Mean Curvature Flow (IMCF) used by Huisken and Ilmanen in their celebrated proof of the Riemannian Penrose inequality [28]. Indeed, a smooth embedded horizontal IMCF for a manifold $\Sigma$ of dimension $2n$ inside $\Omega$ starting from $\Sigma_{0}\subseteq\overline{\Omega}$ is a smooth map $F:\Sigma\times[0,S)\to\mathbb{H}^{n}$ such that (i) $\Sigma_{0}=F(\Sigma,0)$ and $\Sigma_{s}=F(\Sigma,s)\subseteq\Omega$ for $s\in(0,S)$, (ii) $\{F(\cdot,s)\}_{s\in[0,S)}$ are embeddings whose time [^2] Although counterintuitive, in the present paper we call $s$ the time variable in the IMCF, keeping the usual notation $t$ for the variable along the center of the group $\mathbb{H}^{n}$.derivative for $s\in(0,S)$ is given by
$\frac{\partial F}{\partial s}=-\frac{\mathcal{H}_{H}}{|\mathcal{H}_{H}|_{0}^{2}}\,,$
where $\mathcal{H}_{H}\neq 0$ at a point $(p,s)\in\Sigma\times(0,S)$ is the horizontal mean curvature vector of the hypersurface $\Sigma_{s}=F(\Sigma,s)$ at the point $F(p,s)$. In case of regular hypersurfaces, if a smooth function $u:\overline{\Omega}\to[0,\infty)$ with $\nabla_{0}u\neq 0$ in $\overline{\Omega}$ satisfies (1.1) then the family $\{\Sigma^{\varepsilon}_{s}\}_{s\in[0,S)}$, with $\Sigma_{s}:=\{u=s\}$, $s\in[0,S)$ and $S>0$ small enough, is a smooth horizontal IMCF, with $\Sigma_{s}=F(\Sigma,s)$ and the map $F$ being the flow map associated to $\frac{\nabla_{0}u}{|\nabla_{0}u|_{0}^{2}}$. In particular, when the initial surface is $\Sigma_{0}=\partial B_{R_{0}}(g_{0})$ then according to (1.2) we have $\Sigma_{s}=\partial B_{R_{s}}(g_{0})$, where ${R_{s}=R_{0}e^{\frac{s}{Q-1}}}$ and $s\in[0,\infty)$, and indeed (1.1) is satisfied at least wherever $\nabla_{0}U\neq 0$, i.e., away from the center of the group.
Thus, together with [15], our paper provides a first attempt in understanding weak solutions for the horizontal IMCF in the sub-Riemannian Heisenberg group $\mathbb{H}^{n}$, in analogy with similar results for the horizontal MCF (see [7], [8], [20] and [18]) and the horizontal Gauss flow (see [26]).
Together with (1.1), we also consider for $0<\varepsilon\leq 1$ positive solutions to the family of problems
(1.3) $\left\{\begin{array}[]{rl}{\rm div}_{\varepsilon}\left(\dfrac{\nabla_{\varepsilon}u}{|\nabla_{\varepsilon}u|_{\varepsilon}}\right)=|\nabla_{\varepsilon}u|_{\varepsilon}&\textrm{in }\Omega,\\ u=0&\textrm{on }\partial\Omega,\end{array}\right.$
where $\rm{div}_{\varepsilon}$ and $\nabla_{\varepsilon}u$ are computed w.r.to the standard left invariant Riemannian metrics $|\cdot|_{\varepsilon}$ and the equation (1.3) give the analogous level set formulation for the corresponding IMCFs.
As first pointed out in [28], for $0\leq\varepsilon\leq 1$ equations (1.1)-(1.3) formally correspond to the Euler-Lagrange equations for minimizers of the (non differentiable) energy functionals
(1.4) $J^{\varepsilon}_{u}(w;K)=\int_{K}|\nabla_{\varepsilon}w|_{\varepsilon}+w|\nabla_{\varepsilon}u|_{\varepsilon}\,,$
among functions $w$ which are compactly supported perturbations of $u$, i.e., $\overline{\{u\neq w\}}\subset K$ and $K\subset\Omega$ any compact set. Here and throughout the paper, all the volume integrals are computed w.r.to the Lebesgue measure $\mathcal{L}^{2n+1}$ in $\mathbb{R}^{2n+1}\simeq\mathbb{H}^{n}$ (i.e., up to a factor the unique Haar measure on $\mathbb{H}^{n}$), also denoted sometimes by ||$\cdot$|| ⋅ | in the sequel, and it will be deliberately omitted in all the integrals.
In the present paper, following [28] we consider solutions to (1.1) in the sense of the following definition.
Definition 1.1.
A locally $d_{\varepsilon}$-Lipschitz function $u:\overline{\Omega}\to\mathbb{R}$ is a weak solution of (1.1) or (1.3) if $u\equiv 0$ on $\partial\Omega$ and for every compact set $K\subset\Omega$ and every locally $d_{\varepsilon}$-Lipschitz function $w:\overline{\Omega}\to\mathbb{R}$ such that $u\equiv w$ in $\Omega\setminus K$ the inequality
(1.5) $J^{\varepsilon}_{u}(u;K)\leq J^{\varepsilon}_{u}(w;K)\,$
holds. A weak solution is proper if ${\lim_{\|g\|\to\infty}u(g)=+\infty}$.
In order to deal with (1.1)-(1.3) we adapt the very clever approach introduced in the important paper [34] for the Euclidean case, see also [35], and then fruitfully exploited in the Riemannian and Finsler case, see e.g. [22, 23, 32, 40, 16, 6]. Thus, for fixed $\varepsilon\in(0,1]$ we are are going to construct solutions $u^{\varepsilon}:\overline{\Omega}\subseteq\mathbb{H}^{n}\to{\mathbb{R}}$ of (1.3) as limits as $p\downarrow 1$ of solutions $u^{\varepsilon}_{p}:\overline{\Omega}\subseteq\mathbb{H}^{n}\to{\mathbb{R}}$ to the regularized problems
(1.6a) ${\rm div}_{\varepsilon}\left(|\nabla_{\varepsilon}u^{\varepsilon}_{p}|_{\varepsilon}^{p-2}\nabla_{\varepsilon}{u^{\varepsilon}_{p}}\right)=|\nabla_{\varepsilon}u^{\varepsilon}_{p}|_{\varepsilon}^{p}$ in $\quad\textrm{in }\Omega,$
(1.6b) $u^{\varepsilon}_{p}=0$ on $\quad\textrm{on }\partial\Omega.$
On the other hand, concerning (1.1) we construct a solution $u:\overline{\Omega}\subseteq\mathbb{H}^{n}\to{\mathbb{R}}$ as $u=\lim u^{\varepsilon}_{p}$ where we first let $\varepsilon\downarrow 0$ and then $p\downarrow 1$ or we take both limits simultaneously.
As it will be shown below, for $0\leq\varepsilon\leq 1$ and $1<p<Q$ problems (1.6a) admit solutions $u_{p}^{\varepsilon}$ in $C(\overline{\Omega})\cap C^{1,\beta}(\Omega)$ where the validity of the equations is a consequence of (but actually equivalent to, see Lemma 7.3) the local minimality property analogous to (1.5) for the energy functionals
(1.7) $J^{p,\varepsilon}_{u_{p}^{\varepsilon}}(w;K)=\int_{K}\dfrac{|\nabla_{\varepsilon}w|_{\varepsilon}^{p}}{p}+w|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}\,,$
where $K\subset\Omega$ is any fixed compact and $\overline{\{w\neq u_{p}^{\varepsilon}\}}\subseteq K$.
As observed in [34], for $p>1$ and $0\leq\varepsilon\leq 1$ the substitution
(1.8) $u=(1-p)\log v\ ,$
turns solutions $u_{p}^{\varepsilon}$ of (1.6a)-(1.6b) into $p$-harmonic functions $v_{p}^{\varepsilon}:\overline{\Omega}\subseteq\mathbb{H}^{n}\to\mathbb{R}$ satisfying
(1.9a) ${\rm div}_{\varepsilon}\left(|\nabla_{\varepsilon}v^{\varepsilon}_{p}|_{\varepsilon}^{p-2}\nabla_{\varepsilon}v^{\varepsilon}_{p}\right)=0$ in $\quad\textrm{in }\Omega,$
(1.9b) $v^{\varepsilon}_{p}=1$ on $\quad\textrm{on }\partial\Omega.$
Thus, on the Heisenberg group both in the Riemannian and in the sub-Riemannian case equations (1.1), (1.3), (1.6a)-(1.6b), (1.8), and (1.9a)-(1.9b) give a link between IMCF and asymptotic behaviour of $p$-harmonic functions as $p\downarrow 1$.
Concerning problem (1.9a)-(1.9b) we will only consider for $1<p<Q$ finite energy solutions, i.e., those minimizing the energy functionals
(1.10) $E_{p}^{\varepsilon}(v)=\frac{1}{p}\int_{\Omega}|\nabla_{\varepsilon}v|_{\varepsilon}^{p}\,,$
among Sobolev functions in
(1.11) $\dot{W}_{1,\varepsilon}^{1,p}(\Omega):=\{v\in\dot{W}_{\varepsilon}^{1,p}(\mathbb{H}^{n})\quad s.t.\,\,v\equiv 1\quad\hbox{a.e. in }\Omega^{c}\}\,.$
Here and in the sequel $\dot{W}_{\varepsilon}^{1,p}(\mathbb{H}^{n})$ denotes for any $\varepsilon\in[0,1]$ the homogeneous Sobolev space of functions in $u\in L^{p^{*}}(\mathbb{H}^{n})$, $p^{*}=\frac{pQ}{Q-p}$, with weak gradient $\nabla_{\varepsilon}u$ such that $|\nabla_{\varepsilon}u|_{\varepsilon}\in L^{p}(\mathbb{H}^{n})$, endowed with the norm $\|u\|_{\dot{W}_{\varepsilon}^{1,p}}=\|\,|\nabla_{\varepsilon}u|_{\varepsilon}\|_{L^{p}}$, the case $\varepsilon=0$ corresponding to the horizontal Sobolev space $H\dot{W}^{1,p}(\mathbb{H}^{n})$ first introduced by Folland and Stein. Clearly such spaces are well defined because of the horizontal Sobolev inequality corresponding to the embedding $H\dot{W}^{1,p}(\mathbb{H}^{n})\subset L^{p^{*}}(\mathbb{H}^{n})$ and actually valid for $1\leq p<Q$ (see (2.4) below).
Note that such optimal functions solving (1.9a)-(1.9b) are nothing but the $p-$capacitary potentials of $\Omega^{c}$. For $1<p<Q$ and $\varepsilon\geq 0$ existence and uniqueness for this minimization problem is a standard consequence of the direct method in the Calculus of Variations for the coercive weakly l.s.c. functional (1.10) on the weakly closed convex subset (1.11) of the reflexive space $\dot{W}_{\varepsilon}^{1,p}(\mathbb{H}^{n})$. For $\varepsilon>0$ the $C^{1,\beta}$-regularity up to the boundary is nowadays classical in the regularity theory for $p$-harmonic functions on Riemannian manifolds (see, e.g., [17] and [31]), with exponent $\beta=\beta(p,\varepsilon)\in(0,1)$. Interior regularity is also known in the sub-Riemannian case $\varepsilon=0$, see, e.g., [37], so that $p-$capacitary potentials actually belong to the space $C(\overline{\Omega})$, where continuity up to the boundary is a consequence of the exterior gauge ball condition ($\mathbf{HP_{\Omega}}$), and with horizontal gradient in $C^{0,\beta^{\prime}}(\Omega)$ i for some $\beta^{\prime}\in(0,1)$. On the other hand, continuity up to the boundary of the horizontal gradient is at present not available for (1.9a) when $\varepsilon=0$ and this is indeed the major technical reason for us to address existence theory for (1.1) through the analysis of the two-parameter family of problems (1.6a)-(1.6b) together with their $p-$harmonic counterparts (1.9a)-(1.9b).
The first result of the present paper gives a uniform pointwise gradient bound for the $p-$capacitary potentials $v^{\varepsilon}_{p}$ in a form of a differential Harnack inequality (sometimes also known as Cheng-Yau’s inequality).
Theorem 1.2.
Let $\Omega\subset\mathbb{H}^{n}$ be an open set with $C^{2}$-smooth boundary and bounded complement satisfying the exterior uniform gauge-ball condition ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}$. For any $1<p<Q$ and $\varepsilon\in[0,1]$ let $v^{\varepsilon}_{p}\in\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ the unique finite energy solution to (1.9a) corresponding to the minimizer of (1.10) in the set (1.11). Then $v^{\varepsilon}_{p}\in C^{1}(\overline{\Omega})$ for any $\varepsilon\in(0,1]$, it is strictly positive in $\overline{\Omega}$ and there exists a constant $C>0$ depending only on $R_{0}$ and $Q$ such that in $\overline{\Omega}$ we have
(1.12) $|\nabla_{\varepsilon}v^{\varepsilon}_{p}|_{\varepsilon}\leq\frac{C}{p-1}v^{\varepsilon}_{p}\,.$
Moreover, for $1<p<Q$ and $\varepsilon=0$ the corresponding finite energy solution $v^{0}_{p}\in\dot{W}^{1,p}_{1,0}(\Omega)$ to (1.9a) satisfies $v^{0}_{p}\in C(\overline{\Omega})$, $\nabla_{0}v^{0}_{p}\in C(\Omega;\mathbb{R}^{2n})$ and (1.12) holds in $\Omega$ with the same $C>0$ above.
In view of (1.8), as a straightforward consequence of Theorem 1.2 one gets the following uniform gradient bound for solutions to (1.6a)-(1.6b).
Corollary 1.3.
Let $\Omega\subset\mathbb{H}^{n}$ be an open set with $C^{2}$-smooth boundary and bounded complement such that $\Omega$ satisfying assuption ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}$. For $1<p<Q$ and $\varepsilon\in[0,1]$ let $v^{\varepsilon}_{p}\in\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ the solution to (1.9a) corresponding to the unique minimizers of (1.10) in the set (1.11) and $u_{p}^{\varepsilon}$ given by (1.8). Then there exists $C>0$ depending only on $R_{0}$ and $Q$ such that
(1.13) $\|\,|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}\|_{L^{\infty}(\Omega)}\leq C\,.$
As already suggested above, the case $\varepsilon=0$ in the theorem is obtained by a limiting argument from the case $\varepsilon\in(0,1]$ for which boundary regularity is available, as the minimizers $\{v^{\varepsilon}_{p}\}$ are strongly convergent in $\dot{HW}^{1,p}(\mathbb{H}^{n})$ to $v^{0}_{p}$ as $\varepsilon\to 0$. Concerning the latter, here the strategy in the proof of Theorem 1.2 and in turn Corollary 1.3 is similar to the one adopted in [34], i.e., one constructs barriers to obtain the bound (1.12) for $v^{\varepsilon}_{p}$ at the boundary $\partial\Omega$ and then propagate this bound in the interior by a suitable use of the maximum principle and with some extra care since $\Omega$ is not bounded (see the remarks at the beginning of Section 6 for a more detailed discussion on this aspect).
Here, in constrast with [34] (as well as [30] and [15]), instead of working with $u^{\varepsilon}_{p}$ solving (1.6a) and applying the maximum principle on the equation satisfyied by $|\nabla_{\varepsilon}u^{\varepsilon}_{p}|^{2}$ we find more convenient to work directly with $v^{\varepsilon}_{p}$ and the corresponding quantity $|\nabla_{\varepsilon}v^{\varepsilon}_{p}|_{\varepsilon}^{2}-C^{2}(v^{\varepsilon}_{p})^{2}$ with the value of $C$ already obtained for the gradient estimate at the boundary. In addition, compared to [34] both the steps mentioned above turn out to be more difficult in the present case. Indeed, explicit solutions in terms of negative powers of the Korányi norm are available only for $\varepsilon=0$, whence, still under assumption ($\mathbf{HP_{\Omega}}$), a suitable perturbation of them must be used as barriers along $\partial\Omega$ in the full range $\varepsilon\in[0,1]$. Furthermore, the maximum principle argument alluded above actually does not seem to apply directly to $|\nabla_{\varepsilon}v^{\varepsilon}_{p}|_{\varepsilon}^{2}-C^{2}(v^{\varepsilon}_{p})^{2}$, so that we have to estimate separately first the vertical derivative $|T_{\varepsilon}v_{p}^{\varepsilon}|^{2}$ and then $|\nabla_{0}v^{\varepsilon}_{p}|_{0}^{2}$, by exploiting the explicit structure in $\mathbb{H}^{n}$ of the Ricci tensor $Ric_{\varepsilon}$ in connection with a key computation adapted from [30] involving the linearized operator (6.8) of (1.9a) at $v^{\varepsilon}_{p}$.
As for those in [34], both the steps our proof rely in an essential way on the $C^{1}$-boundary regularity of $p$-capacitary potentials on Riemannian manifolds, as they need to be differentiable at the boundary for the first step and must be of controlled size in tubular neighborhood of $\partial\Omega$ in order to obtain the second step (by running a contradiction argument well in the interior where even higher regularity is available and used). To our knowledge such boundary regularity property is not yet available in the sub-Riemannian case, therefore we have no direct proof of the theorem in the case $\varepsilon=0$ but we can only infer the same bound from (1.12) by letting $\varepsilon\to 0$, as already announced above. The same lack of boundary regularity affects the argument used for the recent existence result [15, Proposition 5.3] for the boundary value problem (1.1) in $\mathbb{H}^{1}$, so that their proof seems to contain a serious gap and our aim here is to provide an alternative approach to such existence property for IMCF on $\mathbb{H}^{n}$ through the Riemannian approximation.
In the Riemannian case, local gradient bounds for $p$-harmonic functions similar to (1.13) were first obtained in the literature in [30] on complete manifolds under lower bounds on the sectional curvature (see also [39] and [32] for similar results under the weaker assumption of lower bounds on the Ricci curvature). However, notice that along the vertical vector field $T_{\varepsilon}=\varepsilon\partial_{t}$ one has $|T_{\varepsilon}|_{\varepsilon}\equiv 1$ but $Ric_{\varepsilon}(T_{\varepsilon})\to-\infty$ as $\varepsilon\to 0$. Moreover, being uniform w.r.to $\varepsilon$, our global bounds here are clearly stronger and do not follow from previous results in the existing literature. On the other hand, it would be interesting to obtain directly purely sub-Riemannian local gradient estimates as done e.g. [3] in the linear case $p=2$.
As a consequence of the uniformity w.r.to $\varepsilon$, our estimate allows us to pass to the limit both as $p\to 1$ and as $\varepsilon\to 0$, so obtaining the following existence theorem which is the main result of our paper.
Theorem 1.4.
Let $\Omega\subset\mathbb{H}^{n}$ be an open set with $C^{2}$-smooth boundary and bounded complement such that $\Omega$ satisfies an exterior uniform gauge-ball condition ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}$. There exists $C>0$ depending only on $R_{0}$ and $Q$ and for each $0\leq\varepsilon\leq 1$ there exists a $d_{\varepsilon}$-Lipschitz function $u^{\varepsilon}:\mathbb{H}^{n}\to[0,\infty)$ such that $u^{\varepsilon}\equiv 0$ in $\mathbb{H}^{n}\setminus\Omega$, $Lip_{\varepsilon}(u^{\varepsilon})\leq C$ and $u^{\varepsilon}$ is a weak solution of (1.1) or (1.3).
Moreover, if $B_{R_{0}}(g_{0})\subset\Omega^{c}\subset B_{\bar{R}}(g_{0})$ for some $0<R_{0}<2<\bar{R}$ then there exist $\widehat{C}>1$ depending only on $R_{0}$ and $\bar{R}$ and there exists $C_{0}>1$ depending only on $Q$, $\partial\Omega$ and $\bar{R}$, such that we have
(1.14) $(Q-1)\log\frac{\|g_{0}^{-1}*g\|}{\bar{R}}-\log C_{0}\leq u^{\varepsilon}(g)\leq(Q-1)\log\frac{\|g_{0}^{-1}*g\|}{R_{0}}+\varepsilon^{4}\log\widehat{C}\,,$
for every $\varepsilon\in[0,1]$ and every $g\in\overline{\Omega}$. As a consequence, for any $\varepsilon\in[0,1]$ the solutions $u^{\varepsilon}$ are proper and their level sets $\Sigma_{s}:=\{u=s\}$ satisfty
(1.15) $\Sigma_{s}\cap B_{\varphi_{\varepsilon}(s)}(g_{0})=\emptyset\,\qquad\hbox{and}\qquad\Sigma_{s}\subset B_{\psi(s)}(g_{0})\,\quad\hbox{for any}\quad s\geq 0\,,$
where ${\varphi_{\varepsilon}(s)=\frac{R_{0}}{\widehat{C}^{\varepsilon^{4}/Q-1}}\exp{\left(\frac{s}{Q-1}\right)}}$ and ${\psi(s)=C_{0}^{\frac{1}{Q-1}}\bar{R}\exp{\left(\frac{s}{Q-1}\right)}}$.
The proof of the previous result follows from a standard compactness argument based on Corollary 1.3, passing to the limit the minimality property of the solutions $u^{\varepsilon}_{p}$ for the energy functionals $J^{p,\varepsilon}_{u^{\varepsilon}_{p}}(\cdot)$ in (1.7) as $p\to 1$ or $(\varepsilon,p)\to(0,1)$ respectively. Concerning the two-sided bound (1.14), the construction of subsolutions for Riemannian $p$-capacitary potential performed in Proposition 3.3 is precise enough to get the upper bound in (1.14) as $p\to 1$ in the whole range $\varepsilon\in[0,1]$. However, the lower bound in (1.14) and in turn properness of solutions turn out to be more subtle.
In the sub-Riemannian case $\varepsilon=0$ properness can be readily inferred from the construction of barriers in terms of explicit $p$-capacitary potentials for each $p\in(1,Q)$ which give exact solutions of (1.6a)-(1.6b) in the complement of a Korányi ball $B_{\bar{R}}(g_{0})$ and in turn an exact solution as in (1.2) as $p\to 1$. On the other hand, when $\varepsilon>0$ no such explicit solution is known for $p>1$, therefore in the whole range $\varepsilon\in[0,1]$ we prefer to follow a different strategy, exploiting decay properties of the $p$-capacitary potentials $v^{\varepsilon}_{p}$ at infinity. Thus, we obtain the lower bound in (1.14) by combining for $q=\frac{Q(p-1)}{Q-p/2}$ an $L^{\infty}-L^{q}$ bound on annuli together with a global weak-$L^{\sigma}$ bound, for $\sigma=\frac{Q(p-1)}{Q-p}>q$, with controlled dependence on $p>1$ and uniformly in $\varepsilon\in(0,1]$. As discussed in more details in Section 5, this approach allows to handle the whole range $\varepsilon\in[0,1]$, the inequalities (1.14) being eventually a consequence of the pointwise decay $v^{\varepsilon}_{p}(g)\simeq\|g\|^{\frac{p-Q}{p-1}}$ as $\|g\|\to\infty$ for the $p$-capacitary potentials which is the natural counterpart of the weak-$L^{\sigma}$ integrability.
The sharpness of the bounds in (1.14) and the related fact that the spheres in (1.15) do expand with the same speed, the latter depending on the homogeneous dimension $Q=2n+2$ of $\mathbb{H}^{n}$ in accordance with the sub-Riemannian nature of the space at infinity even for $\varepsilon\in(0,1]$, leave open the fundamental problem concerning the asymptotic behaviour of $u^{\varepsilon}$ as $\|g\|\to\infty$. In particular, it would be very interesting to know whether or not both the Riemannian and the sub-Riemannian IMCF evolution of the initial surface $\Sigma_{0}=\partial\Omega$ is asymptotically equivalent to a Korányi sphere with exponentially growing radius ${R_{s}\sim e^{\frac{s}{Q-1}}}$ as the time $s\to\infty$, in analogy with the results in [24] and [38] for the Euclidean case. The analysis of this problem seems to require a much better understanding of the fine asymptotic behaviour at infinity of the Riemannian $p$-capacitary potentials $v^{\varepsilon}_{p}$. In particular, both for a complement of a Korányi ball and for more general exterior domains, it would be interesting to study their decay properties as $\|g\|\to\infty$ in terms of an asymptotic expansion with an explicit dependence with respect to $p$ as $p\to 1$, in analogy with the analysis performed in [29] in the Euclidean spaces and in [4] on a class of manifolds with asymptotically nonnegative Ricci curvature.
As the fine behaviour of $p$-capacitary potential on Riemannian manifold is gaining considerable interest in recent years, see, e.g., [22], [1], [2], [23], we find this aspect worth of further investigation. At the same time in a subsequent paper for the solution given in Theorem 1.4 we aim to investigate curvature bounds and higher regularity of the level sets $\left\{\Sigma^{\varepsilon}_{s}\right\}_{s\geq 0}\subseteq\mathbb{H}^{n}$, where $\Sigma^{\varepsilon}_{s}=\{u^{\varepsilon}=s\}$, together with the validity of the Geroch monotonicity formulas from [36] both in the Riemannian and the sub-Riemannian setting.
The plan of the paper is as follows. In Section 2 we collect basic notations and facts about the Heisenberg group $\mathbb{H}^{n}$ and the relevant functional inequalities used here. In Section 3 we construct the relevant barriers for the boundary gradient estimate for $p$-capacitary potentials. In Section 4 we present two auxiliary results for $p$-harmonic functions, i.e., $L^{\infty}-L^{p}$ gradient estimates and Harnack inequality, which hold uniformly w.r.to the Riemannian approximation of the sub-Riemannian metric. In Section 5 we derive the weak-$L^{\sigma}$ bound and in turn the uniform pointwise two-sided bounds for the Riemannian $p$-capacitary potentials which hold globally in the domain. In Section 6 for the Riemannian $p$-capacitary potentials we derive gradient estimate at the boundary $\partial\Omega$ and we transfer it to the whole $\overline{\Omega}$ obtaining the Cheng-Yau’s inequality. In Section 7 we finally prove the main results of the paper, i.e., Theorem 1.2 and Theorem 1.4.
Acknowledgments: We thank Luca Capogna and Giovanna Citti for several interesting and useful discussions on the content of some results of Section 4 and the connections with their papers [9] and [10].
2. Basic notations and preliminary results
In this section we first collect the basic definitions and properties of the Heisenberg group $\mathbb{H}^{n}$ which will be needed throughout the paper, referring to [13] for a comprehensive introduction to the subject.
The Heisenberg group $\mathbb{H}^{n}$ is the non-abelian homogeneous Lie group $({\mathbb{R}}^{2n+1},\ast,\delta_{\lambda})$, where $\ast$ is the group operation and $\left(\delta_{\lambda}\right)_{\lambda>0}$ is the anisotropic family of dilations. More precisely, given two points $g=(x_{1},\ldots,x_{n},y_{1},\ldots,y_{n},t)$ and $g^{\prime}=(x^{\prime}_{1},\ldots,x^{\prime}_{n},y^{\prime}_{1},\ldots,y^{\prime}_{n},t^{\prime})$ in $\mathbb{H}^{n}$,
$g\ast g^{\prime}=\left(x_{1}+x^{\prime}_{1},\ldots,x_{n}+x^{\prime}_{n},y_{1}+y^{\prime}_{1},\ldots,y_{n}+y^{\prime}_{n},t+t^{\prime}+\tfrac{1}{2}\sum_{i=1}^{n}(x_{i}y^{\prime}_{i}-x^{\prime}_{i}y_{i})\right),$
and
$\delta_{\lambda}(g)=\left(\lambda x_{1},\ldots,\lambda x_{n},\lambda y_{1},\ldots,\lambda y_{n},\lambda^{2}t\right)\qquad\textrm{for every }\lambda>0.$
As already done in the Introduction, throughout the paper we denote by $Q=2n+2$ the homogeneous dimension of the group $\mathbb{H}^{n}$. Moreover, we will use the notation
$L_{g_{0}}(g):=g_{0}\ast g,\quad g\in\mathbb{H}^{n},$
to denote the left translation by any given point $g_{0}\in\mathbb{H}^{n}$, $g_{0}=0$ being the identity element.
The Heisenberg group is the simplest model of the so called Carnot or stratified Lie groups, indeed it presents a stratified structure at the level of its Lie algebra $\mathfrak{h}=\mathfrak{h}_{0}\oplus\mathfrak{h}_{1}\simeq T_{0}\mathbb{H}^{n}$. By left translation $\mathfrak{h_{0}}$ corresponds to the horizontal distribution, i.e., to the span at any $g=(x_{1},\ldots,x_{n},y_{1},\ldots,y_{n},t)$ of the horizontal vector fields $\{X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{n}\}$,
$X_{i}=\partial_{x_{i}}-\tfrac{y_{i}}{2}\partial_{t},\qquad Y_{i}=\partial_{y_{i}}+\tfrac{x_{i}}{2}\partial_{t}.$
On the other hand at the identity $\mathfrak{h}_{1}=\mathbb{R}T$, where $T=\partial_{t}=[X_{i},Y_{i}]$, for $i=1,\ldots,n$, is the vertical vector field and coincides with the commutator of horizontal vector fields. More generally, we have $[X_{i},Y_{j}]=\delta_{i,j}T$ and $[X_{i},X_{j}]=[Y_{i},Y_{j}]=[X_{i},T]=[Y_{i},T]=0$ for $i,j=1,\ldots,n$, which will be useful below.
For every $\varepsilon>0$ we introduce
$T_{\varepsilon}:=\varepsilon T,$
and we consider the family of Riemannian metrics $|\cdot|_{\varepsilon}$ such that $\{X_{i},Y_{i},T_{\varepsilon}\}$ form an orthonormal basis. Thus, given $V=\sum_{i=1}^{n}\left[a_{i}X_{i}+b_{i}Y_{i}\right]+cT_{\varepsilon}$ and $W=\sum_{i=1}^{n}\left[a^{\prime}_{i}X_{i}+b^{\prime}_{i}Y_{i}\right]+c^{\prime}T_{\varepsilon}$, we denote their scalar product and induced norms as
$\langle{V},{W}\rangle_{\varepsilon}=\sum_{i=1}^{n}\left[a_{i}a^{\prime}_{i}+b_{i}b^{\prime}_{i}\right]+cc^{\prime},\qquad|V|_{\varepsilon}^{2}=\langle{V},{V}\rangle_{\varepsilon}\,,$
with the same formula for $\varepsilon=0$ valid with $c=c^{\prime}\equiv 0$.
Note that for any $\varepsilon>0$ with respect to the corresponding Levi-Civita connection $\nabla^{\varepsilon}$ one has for arbitrary $U,V,W\in\mathfrak{h}$ the identity
$\langle{\nabla^{\varepsilon}_{U}V},{W}\rangle_{\varepsilon}=\frac{1}{2}\left(\langle{[U,V]},{W}\rangle_{\varepsilon}-\langle{U},{[V,W]}\rangle_{\varepsilon}-\langle{V},{[U,W]}\rangle_{\varepsilon}\right)\,,$
whence
$\nabla_{X_{i}}X_{j}$ $=\nabla^{\varepsilon}_{Y_{i}}Y_{j}=0,\quad\textrm{for all }i,j=1,\ldots,n,\quad\textrm{and}\quad\nabla^{\varepsilon}_{X_{i}}Y_{j}=0,\quad\textrm{for all }i\neq j,$
$\nabla^{\varepsilon}_{X_{i}}Y_{i}$ $=-\nabla^{\varepsilon}_{Y_{i}}X_{i}=\dfrac{T_{\varepsilon}}{2\varepsilon},\quad\nabla^{\varepsilon}_{X_{i}}T_{\varepsilon}=\nabla^{\varepsilon}_{T_{\varepsilon}}X_{i}=-\dfrac{Y_{i}}{2\varepsilon},\quad\textrm{and}\quad\nabla^{\varepsilon}_{Y_{i}}T_{\varepsilon}=\nabla^{\varepsilon}_{T_{\varepsilon}}Y_{i}=\dfrac{X_{i}}{2\varepsilon},\quad\textrm{for all }i=1,\ldots,n.$
For a given smooth function $u:\mathbb{H}^{n}\to{\mathbb{R}}$ its Riemannian gradient is the vector field given by
$\nabla_{\varepsilon}u=\sum_{i=1}^{n}\left[(X_{i}u)X_{i}+(Y_{i}u)Y_{i}\right]+(T_{\varepsilon}u)T_{\varepsilon}\,,$
whereas for $\varepsilon=0$ its horizontal gradient is the vector field given by $\nabla_{0}u=\sum_{i=1}^{n}\left[(X_{i}u)X_{i}+(Y_{i}u)Y_{i}\right]$.
On the other hand, for any smooth vector field $Z_{\varphi}=\sum_{i=1}^{n}\left[\varphi_{i}X_{i}+\varphi_{n+i}Y_{i}\right]+\varphi_{2n+1}T_{\varepsilon}$ associated to a smooth map $\varphi:\mathbb{R}^{2n+1}\to\mathbb{R}^{2n+1}$ its Riemannian divergence is given by
${\rm div}_{\varepsilon}Z_{\varphi}=\sum_{i=1}^{n}\left[(X_{i}\varphi_{i}+Y_{i}\varphi_{n+i}\right]+T_{\varepsilon}\varphi_{2n+1}\,,$
whereas for $\varepsilon=0$ its horizontal divergence is given by ${\rm div}_{0}Z_{\varphi}=\sum_{i=1}^{n}\left[(X_{i}\varphi_{i}+Y_{i}\varphi_{n+i}\right]$
In the region where $\nabla_{0}u\neq 0$ each level set $\{u=s\}$, $s\in\mathbb{R}$, defines a smooth surface with horizontal normal $\frac{\nabla_{0}u}{|\nabla_{0}u|_{0}}$ and horizontal mean curvature
(2.1) $\mathcal{H}_{H}:={\rm div}_{0}\frac{\nabla_{0}u}{|\nabla_{0}u|_{0}}=\sum_{i=1}^{n}\left[X_{i}\frac{X_{i}u}{|\nabla_{0}u|_{0}}+Y_{i}\frac{Y_{i}u}{|\nabla_{0}u|_{0}}\right]\,.$
Analogous formula holds for the mean curvature $\mathcal{H}_{\varepsilon}$ in the region where $\nabla_{\varepsilon}u\neq 0$, showing in particular that $\mathcal{H}_{\varepsilon}(g)\to\mathcal{H}_{H}(g)$ whenever $\nabla_{0}u(g)\neq 0$.
[…]
A major role in the paper is played by the Korányi norm. For any point $g=(x,y,t)\in\mathbb{H}^{n}$ with $x=(x_{1},\ldots,x_{n})$ and $y=(y_{1},\ldots,y_{n})$, its Korányi norm is given by
$\|g\|:=\left((|x|^{2}+|y|^{2})^{2}+16t^{2}\right)^{1/4}=\left(|z|^{4}+16t^{2}\right)^{1/4}\,,$
where ||$\cdot$|| ⋅ | is the Euclidean norm in $\mathbb{R}^{n}$ and to simplify the notation we denote from now on $(|x|^{2}+|y|^{2})$ simply by $|z|^{2}$, with $z=(x,y)$.
Given a point $g_{0}\in\mathbb{H}^{n}$ and number $r>0$, we denote by $B_{r}(g_{0})$ the open Korányi ball of center $g_{0}$ and radius $r$. More precisely, for any $g_{0}\in\mathbb{H}^{n}$
$B_{r}(g_{0}):=\left\{g\in\mathbb{H}^{n}:\|g_{0}^{-1}\ast g\|<r\right\}\,,$
so that $L_{g_{0}}(B_{r}(0))=B_{r}(g_{0})$ As it is well known, $d_{0}(\cdot,\cdot)$ and $\|(\,\cdot\,)^{-1}\cdot\|$ give equivalent distances and both induce the Euclidean topology on $\mathbb{H}^{n}$, so that only Korányi balls will be used for $\varepsilon=0$.
Here we recall that for an open set with smooth boundary $\Omega\subseteq\mathbb{H}^{n}$ a point $p\in\partial\Omega$ is characteristic if $T_{p}\Omega=span\{X_{1}(p),\ldots,X_{n}(p),Y_{1}(p),\ldots,Y_{n}(p)\}$. In case of a Korány ball radius $r>0$ centered at the origin the characteristic points are precisely $g=(0,0,\pm r^{2}/4)$.
Definition 2.1.
[…]
Clearly such an assumption amounts to require a quantitative one-sided flatness of $\partial\Omega$ at any of its characteristic point and it is clearly satisfied by the complement of any gauge-ball but also by gauge balls themselves (for the latter observe that for any $g\in\partial B_{r}(\bar{g})$ under the choice $g_{0}=g*{\bar{g}}^{-1}*g$ and $R_{0}=r$ it is straightforward to check that the ball $B_{R_{0}}(g_{0})$ has the desired properties because of Euclidean convexity of gauge balls).
Given the Riemannian metrics $|\cdot|_{\varepsilon}^{2}$, $\varepsilon>0$, the corresponding volume measures $dvol_{\varepsilon}(\cdot)$ are left-invariant and $dvol_{\varepsilon}(\cdot)=\varepsilon^{-1}\mathcal{L}^{2n+1}$, i.e., they coincide with the Lebesgue measure up to a constant factor. Note that, besides translation invariance, the Lebesgue measure under dilation satisfies the equality $\mathcal{L}^{2n+1}(\delta_{\lambda}(E))=\lambda^{Q}\mathcal{L}^{2n+1}(E)$ for any measurable subset $E\subseteq\mathbb{H}^{n}$ and for any $\lambda>0$. For bounded open set $\widetilde{\Omega}\subset\mathbb{H}^{n}$ with $C^{1}$-smooth boundary and $\varepsilon>0$ we will also consider the notion of Riemannian perimeter of $\widetilde{\Omega}$ according to the standard definition, by duality with vector fields $Z_{\varphi}=\sum_{i=1}^{n}\left[\varphi_{i}X_{i}+\varphi_{n+i}Y_{i}\right]+\varphi_{2n+1}T_{\varepsilon}$ associated to a smooth map $\varphi:\mathbb{H}^{n}\to\mathbb{R}^{2n+1}$, given by
$Per_{\varepsilon}(\widetilde{\Omega})=\sup\left\{\int_{\widetilde{\Omega}}{\rm div}_{\varepsilon}Z_{\varphi}\,dvol_{\varepsilon}\,\,;\quad\varphi\in C_{0}^{\infty}(\mathbb{H}^{n};\mathbb{R}^{2n+1})\,,\,|\varphi|\leq 1\right\}\,.$
As it is well known, due to the Riemannian divergence theorem we have $Per_{\varepsilon}(\widetilde{\Omega})=\mathcal{H}^{2n}_{\varepsilon}(\partial\widetilde{\Omega})$, where $\mathcal{H}^{2n}_{\varepsilon}(\cdot)$ is the Hausdorff measure associated to $d_{\varepsilon}$, which is also equal to the area computed with respect to the Riemannian volume form coming from the metric induced on $\partial\widetilde{\Omega}$. In the same way for $\varepsilon=0$ following, e.g., [12], [33] and [21], we will consider the orizontal perimeter $Per_{0}(\widetilde{\Omega})$ computed with a similar formula involving the horizontal divergence, horizontal vector fields (i.e., those with $\varphi_{2n+1}\equiv 0$) and where integration is with respect to $\mathcal{L}^{2n+1}$. The useful convergence property $\varepsilon Per_{\varepsilon}(\widetilde{\Omega})\to Per_{0}(\widetilde{\Omega})$ as $\varepsilon\to 0$ will be be discussed in Lemma 5.2 together with a representation formula for any $\varepsilon\geq 0$ in the spirit of [12, Section 3] and [33, Section 5].
Next, we recall three basic properties of the spaces $(\mathbb{H}^{n},d_{\varepsilon},dvol_{\varepsilon})$ which will be relevant in the sequel. Measures $dvol_{\varepsilon}(\cdot)$ are uniformly doubling, i.e., according to [19, Theorem 3.6], there exists constant $C_{D}\geq 1$ depending only on $Q$ such that for every $\varepsilon>0$ and $\bar{g}\in\mathbb{H}^{n}$ we have
(2.2) $|B^{\varepsilon}_{2r}(\bar{g})|\leq C_{D}|B^{\varepsilon}_{r}(\bar{g})|\quad\hbox{ for any }r>0\,.$
It follows from the doubling property that $dvol_{\varepsilon}(\cdot)$ support weak $(1,1)$-Poincarè inequality with uniform constants. More precisely, according to [19, Theorem 4.2], there exist constants $C_{P}\geq 1$ depending only on $Q$ such that for every $\varepsilon>0$, $\bar{g}\in\mathbb{H}^{n}$ and $r>0$ we have
(2.3) $\frac{1}{|B^{\varepsilon}_{r}(\bar{g})|}\int_{B^{\varepsilon}_{r}(\bar{g})}|U-U_{B^{\varepsilon}_{r}(\bar{g})}|\leq C_{P}r\frac{1}{|B^{\varepsilon}_{3r}(\bar{g})|}\int_{B^{\varepsilon}_{3r}(\bar{g})}|\nabla_{\varepsilon}U|_{\varepsilon}\,,$
where $U\in W^{1,1}_{\varepsilon}(B^{\varepsilon}_{3r}(\bar{g}))$ and $U_{B^{\varepsilon}_{r}(\bar{g})}$ is the average of $U$ over $B^{\varepsilon}_{r}(\bar{g})$.
Next, it is an immediate consequence of the horizontal Sobolev inequality for $p=1$ (i.e., the continuous embedding $H\dot{W}^{1,1}(\mathbb{H}^{n})\subset L^{1^{*}}(\mathbb{H}^{n})$ proved in [12, Theorem 1.1]) that for any $1\leq p<Q$ there exists a constant $C>0$ (depending only on $Q$) such that for $p^{*}=Qp/(Q-p)$ and for any $\varepsilon\geq 0$ we have
(2.4) $\left(\int_{\mathbb{H}^{n}}|U(g)|^{p^{*}}\right)^{1/p^{*}}\leq Cp^{*}\left(\int_{\mathbb{H}^{n}}|\nabla_{\varepsilon}U(g)|^{p}_{\varepsilon}\right)^{1/p}\,\quad\hbox{for any}\quad U\in\dot{W}^{1,p}_{\varepsilon}(\mathbb{H}^{n})\,.$
It is a remarkable consequence of (2.2) and (2.3) that Sobolev inequality also holds in a localized version uniformly on $\varepsilon>0$ (see [19, Theorem 4.6], see also [9, Lemma 7.4]). Thus, for any $1<p<Q$ there exists a constant $C_{S}$ (depending only on $p$ and $Q$) such that for every $\varepsilon>0$, $\bar{g}\in\mathbb{H}^{n}$ and $r>0$
(2.5) $\left(\frac{1}{|B^{\varepsilon}_{r}(\bar{g})|}\int_{B^{\varepsilon}_{r}(\bar{g})}|U(g)|^{p^{*}}\right)^{1/p^{*}}\leq C_{S}\,r\left(\frac{1}{|B^{\varepsilon}_{r}(\bar{g})|}\int_{B^{\varepsilon}_{r}(\bar{g})}|\nabla_{\varepsilon}U|^{p}_{\varepsilon}\right)^{1/p}\,\,,$
for any $U\in\dot{W}^{1,p}_{\varepsilon}(\mathbb{H}^{n})$ with support in some ball $B^{\varepsilon}_{r}(\bar{g})\subseteq\mathbb{H}^{n}$.
The proofs of (2.2), (2.3), and (2.5) in [19] rely on the equivalence (uniformly on $\varepsilon>0$, see [19, Corollary 3.9] ) of the distance $d_{\varepsilon}$ with the $\varepsilon$-gauge $\mathcal{N}_{\varepsilon}$, where $\mathcal{N}_{\varepsilon}(z,t)=|z|+{\rm min}\left\{\frac{|t|}{\varepsilon},\sqrt{|t|}\right\}$ (see also [9] for an alternative definition of equivalent $\varepsilon$-gauge). Such equivalence makes more transparent, at least on the subgroup $\mathbb{R}\simeq\{(0,0,t)\in\mathbb{H}^{n},\,\,t\in\mathbb{R}\}$, the Euclidean character of $d_{\varepsilon}$ at distances of order below $\varepsilon^{2}$ and its sub-Riemannian character otherwise. Here we adopt still another definition of a (uniformly) equivalent $\varepsilon$-gauge, namely $\|\cdot\|_{\varepsilon}$, where
(2.6) $\|g\|_{\varepsilon}:=\left(|z|^{4}+16\left({\rm min}\left\{\frac{|t|}{\varepsilon},\sqrt{|t|}\right\}\right)^{4}\right)^{1/4}={\rm min}\left\{\left(|z|^{4}+16\frac{t^{4}}{\varepsilon^{4}}\right)^{1/4},\|g\|\right\}\,,$
so that $\|\cdot\|_{\varepsilon}\to\|\cdot\|$ pointwise on $\mathbb{H}^{n}$ as $\varepsilon\to 0$. Then, straightforward computations in combination with [19, Corollary 3.7 and Corollary 3.9] lead to the following result that collects relevant properties the $\varepsilon$-gauge above and their consequences.
Lemma 2.2.
Let $\varepsilon>0$, $\|\cdot\|_{\varepsilon}$ as in (2.6) and $\mathcal{B}^{\varepsilon}_{r}(\bar{g})=\{g\in\mathbb{H}^{n}\hbox{ s.t. }\|\bar{g}^{-1}*g\|_{\varepsilon}<r\}$ the corresponding balls. Then there exists $C>1$ depending only on $Q$ such that the following five statements hold.
  1. (1)
    $\|g\|_{\varepsilon}\leq\|g\|$ for any $g\in\mathbb{H}^{n}$, whence $B_{r}(\bar{g})\subseteq\mathcal{B}^{\varepsilon}_{r}(\bar{g})$ for any $\bar{g}\in\mathbb{H}^{n}$ and for any $r>0$. Moreover, $\|g\|_{\varepsilon}\geq\frac{1}{2}\|g\|$ whenever $g\in\mathbb{H}^{n}$ satisfies $\|g\|_{\varepsilon}\geq 2\varepsilon$.
  2. (2)
    For any $\bar{g}\in\mathbb{H}^{n}$ we have $C^{-1}r^{Q}\leq|B_{r}(\bar{g})|\leq|B^{\varepsilon}_{r}(\bar{g})|$ for any $r>0$ and in turn $C^{-1}r^{Q}\leq|B^{\varepsilon}_{4r}(\bar{g})\setminus\overline{B_{2r}(\bar{g})}|$ for any $r>0$. Furthermore for any $\bar{g}\in\mathbb{H}^{n}$ we have $|B^{\varepsilon}_{r}(\bar{g})|\leq Cr^{Q}$ for any $r\geq\varepsilon$.
  3. (3)
    ${C^{-1}\|g\|_{\varepsilon}\leq d_{\varepsilon}(0,g)\leq C\|g\|_{\varepsilon}}$ for any $g\in\mathbb{H}^{n}$, i.e., $d_{\varepsilon}$ and $\|\cdot\|_{\varepsilon}$ are uniformly equivalent and in particular $\mathcal{B}^{\varepsilon}_{C^{-1}r}(\bar{g})\subseteq B^{\varepsilon}_{r}(\bar{g})\subseteq\mathcal{B}^{\varepsilon}_{Cr}(\bar{g})$ for any $\bar{g}\in\mathbb{H}^{n}$ and for any $r>0$.
  4. (4)
    $\|\cdot\|_{\varepsilon}$ are piecewise $C^{1}$ on $\mathbb{H}^{n}$ and ${|\nabla_{\varepsilon}\|\cdot\|_{\varepsilon}|_{\varepsilon}^{2}}+\left|T\|\cdot\|^{2}_{\varepsilon}\right|\leq C$ on $\mathbb{H}^{n}$; in particular each $\|\cdot\|_{\varepsilon}$ is $d_{\varepsilon}$-Lipschitz.
  5. (5)
    If $0<\rho_{1}<\rho_{2}\leq 2\rho_{1}$ and $h:[0,\infty)\to[0,1]$ is a Lipschitz and piecewhise linear function, with $\{h\equiv 1\}=[0,\rho_{1}]$, $\{h\equiv 0\}=[\frac{\rho_{1}+\rho_{2}}{2},\infty)$, and $h$ is linear otherwise, then for any $\bar{g}\in\mathbb{H}^{n}$ the function $\zeta(g)=h(\|\bar{g}^{-1}*g\|_{\varepsilon})$ is piecewise $C^{1}$ on $\mathbb{H}^{n}$ with $\zeta\equiv 1$ on $\mathcal{B}^{\varepsilon}_{\rho_{1}}(\bar{g})$, $spt\,\zeta\subset\mathcal{B}^{\varepsilon}_{\rho_{2}}(\bar{g})$ and
    (2.7) $|\nabla_{\varepsilon}\zeta|_{\varepsilon}^{2}+\left|T\zeta\right|\leq\frac{4C}{(\rho_{2}-\rho_{1})^{2}}\,.$
The last result of this section provides a uniform density lower bound for metric balls $B^{\varepsilon}_{r}(\bar{g})$, a fact which will be relevant in connection with the proof of the Harnack inequality.
Lemma 2.3.
Let $\varepsilon>0$ and $B^{\varepsilon}_{r}(\bar{g})\subset\mathbb{H}^{n}$ a fixed ball. There exists $C\geq 1$ depending only on $Q$ such that
(2.8) $|B^{\varepsilon}_{\rho}(\hat{g})|\leq C|B^{\varepsilon}_{\rho}(\hat{g})\cap B^{\varepsilon}_{r}(\bar{g})|\qquad\hbox{for any}\quad\hat{g}\in B^{\varepsilon}_{r}(\bar{g})\quad\hbox{and}\quad\rho\in(0,2r]\,.$
Proof.
Clearly we may assume $\hat{g}\neq\bar{g}$, otherwise the claim holds trivially for $\rho\leq r$ with $C=1$ and in view of (2.2) for $\rho\in(r,2r]$. For fixed $\varepsilon>0$ the space $(\mathbb{H}^{n},d_{\varepsilon})$ is a locally compact Riemannian manifold which is complete as a metric space. It follows from the Hopf-Rinow Theorem that for any $\hat{g}\in B^{\varepsilon}_{r}(\bar{g})$, $\hat{g}\neq\bar{g}$, there exists a (strictly positive) constant speed geodesic $\Gamma:[0,1]\to\mathbb{H}^{n}$ such that $\Gamma(0)=\bar{g}$, $\Gamma(1)=\hat{g}$ and $d_{\varepsilon}(\Gamma(t),\Gamma(t^{\prime}))=|t-t^{\prime}|d_{\varepsilon}(\bar{g},\hat{g})$ for any $t,t^{\prime}\in[0,1]$. Note that the last identity obviously yields $\Gamma(t)\in B^{\varepsilon}_{r}(\bar{g})$ for any $t\in[0,1]$.
Suppose first that $\rho\in[\frac{1}{4}r,2r]$, then we also have $\Gamma(t)\in B^{\varepsilon}_{\rho}(\hat{g})$ for any $t\in[\frac{3}{4},1]$. In addition, $B^{\varepsilon}_{\rho}(\hat{g})\subset B^{\varepsilon}_{3r}(\bar{g})$ by triangle inequality. Since $\tilde{g}:=\Gamma(7/8)\in B^{\varepsilon}_{\rho}(\hat{g})\cap B^{\varepsilon}_{r}(\bar{g})$ and $d_{\varepsilon}(\tilde{g},\bar{g})=7d_{\varepsilon}(\tilde{g},\hat{g})=\frac{7}{8}d_{\varepsilon}(\bar{g},\hat{g})$, then $B^{\varepsilon}_{r/{16}}(\tilde{g})\subset B^{\varepsilon}_{\rho}(\hat{g})\cap B^{\varepsilon}_{r}(\bar{g})$. Indeed for any $g\in B^{\varepsilon}_{r/{16}}(\tilde{g})$ we have $d_{\varepsilon}(g,\bar{g})\leq d_{\varepsilon}(g,\tilde{g})+d_{\varepsilon}(\tilde{g},\bar{g})<r/16+\frac{7}{8}r<r$ and, analogously, $d_{\varepsilon}(g,\hat{g})\leq d_{\varepsilon}(g,\tilde{g})+d_{\varepsilon}(\tilde{g},\hat{g})<r/16+r/8<\rho$.
Thus, using the left invariance of the measure together with the uniform doubling property (2.2) combined with the previous inclusions we have $|B^{\varepsilon}_{\rho}(\hat{g})|\leq|B^{\varepsilon}_{3r}(\bar{g})|=|B^{\varepsilon}_{3r}(\tilde{g})|\leq C|B^{\varepsilon}_{r/16}(\tilde{g})|\leq C|B^{\varepsilon}_{\rho}(\hat{g})\cap B^{\varepsilon}_{r}(\bar{g})|\,,$ which is the desired conclusion.
Finally, suppose instead that $40<\rho<r/$ 4. Clearly we may assume $d_{\varepsilon}(\bar{g},\hat{g})\geq\frac{3}{4}r$, because otherwise $B_{\rho}(\hat{g})\subset B_{r}(\bar{g})$ and (2.8) trivially holds with $C=1$. Then by continuity there exists $t_{1}\in(0,1)$ such that $d_{\varepsilon}(\Gamma(t_{1}),\hat{g})=\rho$ and $d_{\varepsilon}(\Gamma(t),\hat{g})<\rho$ for any $t\in(t_{1},1]$. Now we set $t_{2}=\frac{1}{2}(t_{1}+1)$ and $\tilde{g}=\Gamma(t_{2})\in B^{\varepsilon}_{\rho}(\hat{g})$, so that we clearly have $B^{\varepsilon}_{\rho}(\hat{g})\subset B^{\varepsilon}_{2\rho}(\tilde{g})$. Note that $d_{\varepsilon}(\tilde{g},\bar{g})=d_{\varepsilon}(\bar{g},\hat{g})-d_{\varepsilon}(\tilde{g},\hat{g})<r-\rho/2$ and therefore $B^{\varepsilon}_{\rho/{4}}(\tilde{g})\subset B^{\varepsilon}_{\rho}(\hat{g})\cap B^{\varepsilon}_{r}(\bar{g})$. Indeed for any $g\in B^{\varepsilon}_{\rho/4}(\tilde{g})$ we have $d_{\varepsilon}(g,\bar{g})\leq d_{\varepsilon}(g,\tilde{g})+d_{\varepsilon}(\tilde{g},\bar{g})<\rho/4+r-\rho/2<r$ and, analogously, $d_{\varepsilon}(g,\hat{g})\leq d_{\varepsilon}(g,\tilde{g})+d_{\varepsilon}(\tilde{g},\hat{g})<\rho/4+\rho/2<\rho$.
Hence, using as above the left invariance of the measure together with the uniform doubling property (2.2) combined with the previous inclusions we have $|B^{\varepsilon}_{\rho}(\hat{g})|\leq|B^{\varepsilon}_{2\rho}(\tilde{g})|\leq C|B^{\varepsilon}_{\rho/4}(\tilde{g})|\leq C|B^{\varepsilon}_{\rho}(\hat{g})\cap B^{\varepsilon}_{r}(\bar{g})|\,,$ which is the desired conclusion also in this second case. Note that in both cases we proved inequality (2.8) with a constant $C\geq 1$ depending only on the one in (2.2), hence only on $Q$ as claimed. ∎
3. Construction of barriers
In this section, in the range $1<p<Q$ we provide a family of barriers whch will be used for establishing the boundary gradient estimate for $p$-capacitary potentials. These functions are negative powers of the Korányi norm obtained as suitable deformation of those giving singular $p$-harmonic functions w.r.to the sub-Riemannian metric. As a result of a careful choice of the exponent, they will be singular subsolutions for the sub-Riemannian $p$-Laplace operator in the whole $\mathbb{H}^{n}$ and for the Riemannian $p$-Laplace operator in the complement of any fixed Korányi ball.
We first introduce a polynomial function that will be used frequently in the computation. For any $g=(x,y,t)\in\mathbb{H}^{n}$ we set
(3.1) $N(x,y,t):=\|g\|^{4}=\left(|x|^{2}+|y|^{2}\right)^{2}+16t^{2}=|z|^{4}+16t^{2}\,.$
We list below some useful and straightforward computations:
(3.2) $X_{i}N$ $=4x_{i}|z|^{2}-16y_{i}t,$
(3.3) $Y_{i}N$ $=4y_{i}|z|^{2}+16x_{i}t,$
(3.4) $TN$ $=32t.$
(3.5) $X_{j}X_{i}N$ $=4\delta_{ij}|z|^{2}+8x_{i}x_{j}+8y_{i}y_{j},$
(3.6) $Y_{j}Y_{i}N$ $=4\delta_{ij}|z|^{2}+8x_{i}x_{j}+8y_{i}y_{j},$
(3.7) $TTN$ $=32,$
(3.8) $Y_{j}X_{i}N$ $=8x_{i}y_{j}-16\delta_{ij}t-8x_{j}y_{i},$
(3.9) $X_{j}Y_{i}N$ $=8x_{j}y_{i}+16\delta_{ij}t-8x_{i}y_{j},$
(3.10) $X_{i}TN$ $=TX_{i}N=-16y_{i},$
(3.11) $Y_{i}TN$ $=TY_{i}N=16x_{i}.$
We note that, by (3.2) and (3.3), it follows that
(3.12) $|\nabla_{0}N|_{0}^{2}=\sum_{i=1}^{n}(X_{i}N)^{2}+(Y_{i}N)^{2}=16|z|^{2}N,$
and
(3.13) $\sum_{i=1}^{n}x_{i}X_{i}N+y_{i}Y_{i}N=4|z|^{4}.$
Consequently, for $\varepsilon\in(0,1]$ we also have
(3.14) $|\nabla_{\varepsilon}N|_{\varepsilon}^{2}=|\nabla_{0}N|_{0}^{2}+\varepsilon^{2}(TN)^{2}=16|z|^{2}N+64\varepsilon^{2}16t^{2}\leq 16N^{3/2}+64N\,.$
Finally, by (3.5) and (3.6), with $i=j$, we have
(3.15) $\Delta_{0}N={\rm div}_{0}\left(\nabla_{0}N\right)=8(2+n)|z|^{2}.$
For every $\alpha\in{\mathbb{R}}$, we define the function $\Phi_{\alpha}:\mathbb{H}^{n}\setminus\{0\}\to{\mathbb{R}}$ as
(3.16) $\Phi_{\alpha}(x,y,t):=N(x,y,t)^{\alpha}=\left((|x|^{2}+|y|^{2})^{2}+16t^{2}\right)^{\alpha},\qquad(x,y,t)\in\mathbb{H}^{n}\setminus\{0\}\,.$
Given two vector fields $Z_{1}$ and $Z_{2}$, we have
[…]
and
$Z_{j}Z_{i}\Phi_{\alpha}=\alpha(\alpha-1)N^{\alpha-2}(Z_{j}N)(Z_{i}N)+\alpha N^{\alpha-1}Z_{j}Z_{i}N.$
Therefore,
(3.17) $X_{i}\Phi_{\alpha}$ $=\alpha N^{\alpha-1}X_{i}N=\alpha N^{\alpha-1}(4x_{i}|z|^{2}-16y_{i}t),$
(3.18) $Y_{i}\Phi_{\alpha}$ $=\alpha N^{\alpha-1}Y_{i}N=\alpha N^{\alpha-1}(4y_{i}|z|^{2}+16x_{i}t),$
(3.19) $T_{\varepsilon}\Phi_{\alpha}$ $=\alpha N^{\alpha-1}T_{\varepsilon}N=\alpha N^{\alpha-1}32\varepsilon t,$
(3.20) $T_{\varepsilon}X_{i}\Phi_{\alpha}$ $=\varepsilon\left(\alpha(\alpha-1)N^{\alpha-2}(128t(x_{i}|z|^{2}-4y_{i}t))-16\alpha N^{\alpha-1}y_{i}\right),$
(3.21) $T_{\varepsilon}Y_{i}\Phi_{\alpha}$ $=\varepsilon\left(\alpha(\alpha-1)N^{\alpha-2}(128t(y_{i}|z|^{2}+4x_{i}t))+16\alpha N^{\alpha-1}x_{i}\right),$
(3.22) $T_{\varepsilon}T_{\varepsilon}\Phi_{\alpha}$ $=\varepsilon^{2}\left(\alpha(\alpha-1)N^{\alpha-2}(32t)^{2}+32\alpha N^{\alpha-1}\right).$
The following lemma provides smooth explicit sub-Riemannian $p$-subharmonic functions in $\mathbb{H}^{n}\setminus\{0\}$.
Lemma 3.1.
Let $1<p<Q$, $\alpha<0$ and $\Phi_{\alpha}$ as defined in (3.16). Then
(3.23) ${\rm div}_{0}\left(|\nabla_{0}\Phi_{\alpha}|_{0}^{p-2}\nabla_{0}\Phi_{\alpha}\right)=(4|\alpha|)^{p-2}|z|^{p}N^{\beta}4\alpha\left(4\alpha(p-1)-(p-Q)\right),\qquad\beta=\dfrac{2\alpha(p-1)-p}{2}<0\,.$
As a consequence, in $\mathbb{H}^{n}\setminus\{0\}$
${\rm div}_{0}\left(|\nabla_{0}\Phi_{\alpha}|_{0}^{p-2}\nabla_{0}\Phi_{\alpha}\right)\geq 0\quad\Longleftrightarrow\quad\alpha\leq\dfrac{p-Q}{4(p-1)}<0.$
Proof.
To simplify the notation, we will denote $\Phi_{\alpha}$ simply by $\Phi$. Firstly, notice that
(3.24) ${\rm div}_{0}\left(|\nabla_{0}\Phi|_{0}^{p-2}\nabla_{0}\Phi\right)$ $=\langle{\nabla_{0}\left(|\nabla_{0}\Phi|^{p-2}\right)},{\nabla_{0}\Phi}\rangle_{0}+|\nabla_{0}\Phi|^{p-2}\Delta_{0}\Phi$
$=|\nabla_{0}\Phi|^{p-4}\left(\frac{p-2}{2}\langle{\nabla_{0}\left(|\nabla_{0}\Phi|^{2}\right)},{\nabla_{0}\Phi}\rangle_{0}+|\nabla_{0}\Phi|^{2}\Delta_{0}\Phi\right)\,.$
It follows from (3.17) and (3.18) that
(3.25) $\nabla_{0}\Phi=\alpha N^{\alpha-1}\nabla_{0}N\quad\textrm{and }\quad|\nabla_{0}\Phi|_{0}^{2}=16\alpha^{2}|z|^{2}N^{2\alpha-1},$
hence
(3.26) $|\nabla_{0}\Phi|_{0}^{p-4}=(4|\alpha|)^{p-4}|z|^{p-4}N^{(2\alpha-1)(p-4)/2}\,.$
Since for $i=1,\ldots,n$,
$X_{i}\left(|\nabla_{0}\Phi|_{0}^{2}\right)=16\alpha^{2}\left(2x_{i}N^{2\alpha-1}+(2\alpha-1)|z|^{2}N^{2\alpha-2}X_{i}N\right)$
and
$Y_{i}\left(|\nabla_{0}\Phi|_{0}^{2}\right)=16\alpha^{2}\left(2y_{i}N^{2\alpha-1}+(2\alpha-1)|z|^{2}N^{2\alpha-2}Y_{i}N\right)\,,$
combining the previous identities with (3.2), (3.3) and (3.12) we obtain
(3.27) $\frac{p-2}{2}\langle{\nabla_{0}\left(|\nabla_{0}\Phi|_{0}^{2}\right)},{\nabla_{0}\Phi}\rangle_{0}=(4\alpha)^{3}(4\alpha-1)(p-2)|z|^{4}N^{3\alpha-2}\,.$
Let us now consider the second term in the r.h. side of (3.24). By (3.12) and (3.15) we have
(3.28) $\Delta_{0}\Phi$ $=\alpha(\alpha-1)N^{\alpha-2}|\nabla_{0}N|_{0}^{2}+\alpha N^{\alpha-1}\Delta_{0}N$
$=16\alpha(\alpha-1)|z|^{2}N^{\alpha-1}+8(2+n)\alpha|z|^{2}N^{\alpha-1}$
$=8\alpha(2\alpha+n)|z|^{2}\,N^{\alpha-1}\,,$
whence
(3.29) $|\nabla_{0}\Phi|_{0}^{2}\Delta_{0}\Phi=(4\alpha)^{3}2(2\alpha+n)|z|^{4}N^{3\alpha-2}.$
Finally, combining (3.27) and (3.29) with (3.24) and (3.26) we see that
${\rm div}_{0}\left(|\nabla_{0}\Phi|_{0}^{p-2}\nabla_{0}\Phi\right)=(4|\alpha|)^{p-2}|z|^{p}N^{\beta}\left(4\alpha(4\alpha-1)(p-2)+8\alpha(2\alpha+n)\right),$
where
$\beta:=\dfrac{(2\alpha-1)(p-4)}{2}+3\alpha-2=\dfrac{2\alpha(p-1)-p}{2}\,,$
whence (3.23) holds and the conclusion follows. ∎
We now consider the function $\Phi_{\alpha}$ as in (3.16) for $\alpha\leq\tfrac{p-Q}{4(p-1)}$ and for any given point $g_{0}\in\mathbb{H}^{n}$ and $R_{0}>0$ we consider the auxiliary function $\Phi^{(g_{0},R_{0})}_{\alpha}:\mathbb{H}^{n}\setminus\{g_{0}\}\to{\mathbb{R}}$ given by
(3.30) $\Phi^{(g_{0},R_{0})}_{\alpha}(x,y,t):=\dfrac{\Phi_{\alpha}\circ L_{g_{0}^{-1}}(x,y,t)}{R_{0}^{4\alpha}}\,.$
Some elementary properties of $\Phi^{(g_{0},R_{0})}_{\alpha}$ are given in the following auxiliary lemma.
Lemma 3.2.
For $p>1$ and $K>0$ let $\alpha=-\frac{K}{p-1}<0$ and $\Phi^{(g_{0},R_{0})}_{\alpha}$ as in (3.30). Then $\Phi^{(g_{0},R_{0})}_{\alpha}=1$ on $\partial B_{R_{0}}(g_{0})$ and there exist $C_{0}>0$ depending only on $R_{0}$ such that for any $\varepsilon\in(0,1]$ we have
(3.31) $\||\nabla_{\varepsilon}\Phi^{(g_{0},R_{0})}_{\alpha}|_{\varepsilon}\|_{L^{\infty}(\partial B_{R_{0}}(g_{0}))}\leq\frac{C_{0}K}{p-1}\,.$
Moreover, for the function $\tilde{\Phi}^{(g_{0},R_{0})}_{\alpha}=\min\{1,\Phi^{(g_{0},R_{0})}_{\alpha}\}$ we have $\tilde{\Phi}^{(g_{0},R_{0})}_{\alpha}\in\dot{W}_{\varepsilon}^{1,p}(\mathbb{H}^{n})$ whenever $K>\frac{(Q-p)(p-1)}{4p}$ and the norms are uniformly bounded for $\varepsilon\in(0,1]$.
Proof.
Up to translations we may assume $g_{0}=0$. Next for $\Phi=\Phi^{(0,R_{0})}_{\alpha}$ inequality (3.14) yields
(3.32) $|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}=\left|\nabla_{\varepsilon}\left(\frac{N}{R_{0}^{4}}\right)^{\alpha}\right|_{\varepsilon}^{2}=\alpha^{2}\Phi^{2}\frac{|\nabla_{\varepsilon}N|_{\varepsilon}^{2}}{N^{2}}\leq\alpha^{2}\Phi^{2}(64N^{-1}+16N^{-1/2})\,.$
Evaluating on $\partial B_{R_{0}}(0)$ gives the conclusion with $C_{0}=\frac{4}{R_{0}^{2}}\sqrt{4+R_{0}^{2}}$.
Concerning integrability, assuming up to translations $g_{0}=0$ and denoting $A_{j}=\{2^{j}R_{0}<\|\cdot\|\leq 2^{j+1}R_{0}\}$, for any $j\geq 0$, by change of variables we have $\mathcal{H}^{Q}(A_{j})\simeq(2^{Q})^{j}$ and in turn
$\int_{\|\cdot\|>R_{0}}\Phi_{\alpha}^{p^{*}}=R_{0}^{-4\alpha p^{*}}\sum_{j=0}^{\infty}\int_{A_{j}}\left(N^{\alpha}\right)^{p^{*}}\simeq\sum_{j=0}^{\infty}(2^{4\alpha p^{*}})^{j}\mathcal{H}^{Q}(A_{j})\simeq\sum_{j=0}^{\infty}(2^{4\alpha p^{*}+Q})^{j}\,.$
As a consequence $\tilde{\Phi}^{(g_{0},R_{0})}_{\alpha}\in L^{p^{*}}$ if and only if $04\alpha p^{*}+Q<$ 0, which holds for $K>0$ precisely as in the statement above.
On the other hand, applying again (3.32) we have $|\nabla_{\varepsilon}\Phi|_{\varepsilon}\lesssim\Phi N^{-1/4}$ in $\mathbb{H}^{n}\setminus B_{R_{0}}$ uniformly on $\varepsilon\in(0,1]$, whence
$\int_{\|\cdot\|>R_{0}}|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{p}\lesssim\sum_{j=0}^{\infty}\int_{A_{j}}\left(N^{\alpha}\right)^{p}N^{-p/4}\simeq\sum_{j=0}^{\infty}(2^{(4\alpha-1)p})^{j}\mathcal{H}^{Q}(A_{j})\simeq\sum_{j=0}^{\infty}(2^{(4\alpha-1)p+Q})^{j}\,,$
so that the integral is finite and uniformly bounded on $\varepsilon\in(0,1]$ whenever $(4\alpha-1)p+Q<0$, which is precisely the same choice of $K>0$ in the statement above. ∎
Aiming to use functions in (3.30) as barriers in the next section for $\varepsilon>0$, in the following preliminary result we show that they are subsolutions in the complement of Korànyi balls.
Proposition 3.3.
For any $R_{0}>0$, $1<p<Q$ and $\varepsilon\in(0,1]$ let $K\geq\frac{Q-p}{4}+\varepsilon^{4}\left(\frac{2}{R_{0}^{2}}+\frac{8Q}{R_{0}^{4}}\right)$. Then for any $g_{0}\in\mathbb{H}^{n}$ the function $\Phi^{(g_{0},R_{0})}_{\alpha}$ defined in (3.30) with $\alpha(p-1)=-K$ satisfies
(3.33) $\mathrm{div}_{\varepsilon}\left(|\nabla_{\varepsilon}\Phi^{(g_{0},R_{0})}_{\alpha}|_{\varepsilon}^{(p-2)}\nabla_{\varepsilon}\Phi^{(g_{0},R_{0})}_{\alpha}\right)\geq 0\quad\textrm{in }B^{c}_{R_{0}}(g_{0})\,.$
Proof.
Without loss of generality, we can assume up to a translation that $g_{0}=0$. To simplify further the readability, we set $\Phi=\Phi_{\alpha}$ and we define
(3.34) $\Phi^{\prime}:=\Phi^{(0,R_{0})}_{\alpha}=R_{0}^{-4\alpha}\Phi\,.$
We further observe that
(3.35) $\mathrm{div}_{\varepsilon}\left(|\nabla_{\varepsilon}\Phi^{\prime}|_{\varepsilon}^{(p-2)}\nabla_{\varepsilon}\Phi^{\prime}\right)$ $=R_{0}^{-4\alpha(p-1)}\mathrm{div}_{\varepsilon}\left(\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}\right)^{(p-2)/2}\nabla_{\varepsilon}\Phi\right)$
$=R_{0}^{-4\alpha(p-1)}\left[\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}\right)^{\tfrac{p-2}{2}}\Delta_{\varepsilon}\Phi+\langle{\nabla_{\varepsilon}\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}\right)^{\tfrac{p-2}{2}}},{\nabla_{\varepsilon}\Phi}\rangle_{\varepsilon}\right]$
$=:R_{0}^{-4\alpha(p-1)}\left((i)+(ii)\right).$
Now, for every smooth vector field $W$, it holds that
(3.36) $W$ $\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}\right)^{\tfrac{p-2}{2}}=\tfrac{p-2}{2}\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}\right)^{\tfrac{p-4}{2}}W(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2})$
$=\tfrac{p-2}{2}\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}\right)^{\tfrac{p-4}{2}}W\left(\sum_{i=1}^{n}\left[(X_{i}\Phi)^{2}+(Y_{i}\Phi)^{2}\right]+(T_{\varepsilon}\Phi)^{2}\right)$
$=(p-2)\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}\right)^{\tfrac{p-4}{2}}\left[\sum_{i=1}^{n}\left(X_{i}\Phi_{\alpha}WX_{i}\Phi+Y_{i}\Phi_{\alpha}WY_{i}\Phi\right)+T_{\varepsilon}\Phi_{\alpha}WT_{\varepsilon}\Phi\right].$
Applying (3.36) with $W=\nabla_{\varepsilon}\Phi$ in $B_{R_{0}}(0)^{c}$ yields
$(ii)$ $=(p-2)\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}\right)^{\tfrac{p-4}{2}}\sum_{j=1}^{n}\left[(X_{j}\Phi)\left(\sum_{i=1}^{n}(X_{i}\Phi X_{j}X_{i}\Phi+Y_{i}\Phi X_{j}Y_{i}\Phi)+T_{\varepsilon}\Phi X_{j}T_{\varepsilon}\Phi\right)\right]$
$+(p-2)\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}\right)^{\tfrac{p-4}{2}}\sum_{j=1}^{n}\left[(Y_{j}\Phi)\left(\sum_{i=1}^{n}(X_{i}\Phi Y_{j}X_{i}\Phi+Y_{i}\Phi Y_{j}Y_{i}\Phi)+T_{\varepsilon}\Phi Y_{j}T_{\varepsilon}\Phi\right)\right]$
$+(p-2)\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}\right)^{\tfrac{p-4}{2}}T_{\varepsilon}\Phi\left(\sum_{i=1}^{n}(X_{i}\Phi T_{\varepsilon}X_{i}\Phi+Y_{i}\Phi T_{\varepsilon}Y_{i}\Phi)+T_{\varepsilon}\Phi T_{\varepsilon}T_{\varepsilon}\Phi\right).$
Commuting $X_{i}$ and $Y_{i}$ with $T_{\varepsilon}=\varepsilon T$ and reordering with respect to the powers of $\varepsilon$ we obtain
(3.37) $(ii)$ $=(p-2)\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}\right)^{\tfrac{p-4}{2}}\sum_{j=1}^{n}\left[(X_{j}\Phi)\sum_{i=1}^{n}(X_{i}\Phi X_{j}X_{i}\Phi+Y_{i}\Phi X_{j}Y_{i}\Phi)\right]$
$+(p-2)\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}\right)^{\tfrac{p-4}{2}}\sum_{j=1}^{n}\left[(Y_{j}\Phi)\sum_{i=1}^{n}(X_{i}\Phi Y_{j}X_{i}\Phi+Y_{i}\Phi Y_{j}Y_{i}\Phi)\right]$
$+2\varepsilon^{2}(p-2)\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}\right)^{\tfrac{p-4}{2}}\sum_{j=1}^{n}\left(X_{j}\Phi T\Phi X_{j}T\Phi+Y_{j}\Phi T\Phi Y_{j}T\Phi\right)$
$+\varepsilon^{4}(p-2)\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}\right)^{\tfrac{p-4}{2}}(T\Phi)^{2}TT\Phi$
$=(p-2)\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}\right)^{\tfrac{p-4}{2}}\left[\frac{1}{2}\langle{\nabla_{0}|\nabla_{0}\Phi|^{2}_{0}},{\nabla_{0}\Phi}\rangle_{0}+\varepsilon^{2}T\Phi T\left(|\nabla_{0}\Phi|^{2}_{0}\right)+\varepsilon^{4}(T\Phi)^{2}TT\Phi\right]\,.$
On the other hand, using again the identity $T_{\varepsilon}=\varepsilon T$ we also have
(3.38) $(i)$ $=\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}\right)^{\tfrac{p-4}{2}}\left(|\nabla_{0}\Phi|_{0}^{2}+\varepsilon^{2}(T\Phi)^{2}\right)(\Delta_{0}\Phi+\varepsilon^{2}TT\Phi)$
$=\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}\right)^{\tfrac{p-4}{2}}\left[|\nabla_{0}\Phi|_{0}^{2}\Delta_{0}\Phi+\varepsilon^{2}\left(|\nabla_{0}\Phi|_{0}^{2}TT\Phi+(T\Phi)^{2}\Delta_{0}\Phi\right)+\varepsilon^{4}(T\Phi)^{2}TT\Phi\right]\,.$
Thus, (3.35), (3.38) and (3.37) yield
(3.39) $\mathrm{div}_{\varepsilon}\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{(p-2)}\nabla_{\varepsilon}\Phi\right)=(i)+(ii)=\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}\right)^{\tfrac{p-4}{2}}\left[(I)+(II)+(III)\right],$
where
(3.40) $(I):$ $=|\nabla_{0}\Phi|_{0}^{2}\Delta_{0}\Phi+\frac{(p-2)}{2}\langle{\nabla_{0}|\nabla_{0}\Phi|^{2}_{0}},{\nabla_{0}\Phi}\rangle_{0}\,,$
$(II):$ $=\varepsilon^{2}\left(|\nabla_{\varepsilon}\Phi|_{0}^{2}TT\Phi+(T\Phi)^{2}\Delta_{0}\Phi+(p-2)T\Phi T\left(|\nabla_{0}\Phi|^{2}_{0}\right)\right)\,,$
$(III):$ $=\varepsilon^{4}(p-1)(T\Phi)^{2}(TT\Phi)\,.$
Concerning the first term in (3.40) identities (3.27) and (3.29) give
(3.41) $(I)=(4\alpha)^{3}|z|^{4}N^{3\alpha-2}(4\alpha(p-1)+Q-p)>0\,,$
for any $K>\frac{Q-p}{4}$.
Next for $\alpha<0$ and restricting to $\mathbb{H}^{n}\setminus B_{R_{0}}$ we bound separately from below the terms $(II)$ and $(III)$ in (3.40). To this end, notice that on $\mathbb{H}^{n}\setminus B_{R_{0}}(0)$ we have $N\geq R_{0}^{4}$ an therefore
(3.42) $-N^{\gamma}\geq-R_{0}^{4\gamma},\quad\textrm{for every }\gamma<0.$
Concerning $(II)$, using (3.19), (3.22), (3.25) and (3.28) we compute all the terms and obtain
$(T\Phi)^{2}\Delta_{0}\Phi$ $=16(8\alpha)^{3}(2\alpha+n)N^{3\alpha-3}\,|z|^{2}\,t^{2},$
$|\nabla_{0}\Phi|^{2}TT\Phi$ $=(8\alpha)^{3}N^{3\alpha-3}\,|z|^{2}(|z|^{4}+(2\alpha-1)16t^{2})\,,$
$(p-2)T\Phi T\left(|\nabla_{0}\Phi|^{2}_{0}\right)$ $=32(p-2)(8\alpha)^{3}(2\alpha-1)N^{3\alpha-3}\,|z|^{2}\,t^{2}\,,$
whence
(3.43) $(II)$ $=\varepsilon^{2}(8\alpha)^{3}N^{3\alpha-3}|z|^{2}\left[16(2\alpha+n)t^{2}+|z|^{4}+16(2\alpha-1)t^{2}+32(p-2)(2\alpha-1)t^{2}\right]$
$=\varepsilon^{2}(8\alpha)^{3}N^{3\alpha-3}|z|^{2}\left[|z|^{4}+8t^{2}(Q+4(2\alpha-1)(p-1))\right]\,.$
Note that for $p>1$ we have $Q+4(2\alpha-1)(p-1)<0$ by our choice of $\alpha$, hence combining on $\mathbb{H}^{n}\setminus B_{R_{0}}$ the last identity with (3.1) and (3.42) we obtain uniformly w.r.to $\varepsilon\in(0,1]$
(3.44) $(II)\geq\varepsilon^{2}(8\alpha)^{3}|z|^{6}N^{3\alpha-3}\geq\varepsilon^{4}(4\alpha)^{3}|z|^{4}N^{3\alpha-2}8N^{-1/2}\geq\varepsilon^{4}(4\alpha)^{3}|z|^{4}N^{3\alpha-2}8R_{0}^{-2}\,.$
On the other hand, applying (3.19) and (3.22) we get
(3.45) $(III)$ $=\varepsilon^{4}(p-1)(T\Phi)^{2}(TT\Phi)=\varepsilon^{4}(p-1)(32\alpha N^{\alpha-1}t)^{2}\left[32\alpha N^{\alpha-2}(N+(\alpha-1)32t^{2})\right]$
$=\varepsilon^{4}(32\alpha)^{3}(p-1)N^{3\alpha-4}t^{2}\left[|z|^{4}+\underbrace{16(2\alpha-1)t^{2}}_{\leq 0}\right],$
therefore on $\mathbb{H}^{n}\setminus B_{R_{0}}$ for all $\varepsilon\in(0,1]$ by (3.1) and (3.42) we obtain
(3.46) $(III)\geq\varepsilon^{4}8^{3}(4\alpha)^{3}(p-1)N^{3\alpha-4}t^{2}|z|^{4}\geq\varepsilon^{4}32(4\alpha)^{3}(p-1)|z|^{4}N^{3\alpha-2}N^{-1}\geq\varepsilon^{4}(4\alpha)^{3}|z|^{4}N^{3\alpha-2}32QR_{0}^{-4}\,.$
Combining (3.39), (3.41), (3.44), and (3.46) in the whole $\mathbb{H}^{n}\setminus B_{R_{0}}$ we infer
$\mathrm{div}_{\varepsilon}\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{(p-2)}\nabla_{\varepsilon}\Phi\right)\geq(4\alpha^{3})|z|^{4}N^{3\alpha-2}\left(|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}\right)^{\tfrac{p-4}{2}}\left[-4K+Q-p+\varepsilon^{4}\left(\frac{8}{R_{0}^{2}}+\frac{32Q}{R_{0}^{4}}\right)\right]\,,$
whence the conclusion follows from the choice of $K$ as $\alpha<0$. ∎
Remark 3.4.
Note that in order to construct subsolutions in a Riemannian case one has to adjust the exponent comparing to the one in the sub-Riemannian case treated in Lemma 3.1. However, the difference of the exponents corresponding to the two choices can be made smaller and smaller as $\varepsilon$ gets smaller or $R_{0}>0$ gets larger. We will exploit this fact in the proof of Proposition 5.1. On the other hand, the construction of Riemannian supersolutions via neagative powers of the Korányi norm seems to be impossible. Indeed, considering (3.41), (3.43) and (3.45) restricted to the vertical axis (i.e. for $|z|=0$) clearly gives the opposite of the desired sign. A similar behaviour will be faced even considering the $\varepsilon$-gauge defined in (2.6), as they agree with the Korányi norm for $|t|\geq\varepsilon$.
4. Two uniform estimates in mustiscale geometries.
In this section we present two results concerning $p$-harmonic functions on the Riemannian Heisenberg group $(\mathbb{H}^{n},d_{\varepsilon})$ which for fixed $p\in(1,Q)$ hold uniformly w.r.to the parameter $\varepsilon\in(0,1]$, namely, an $L^{p}-L^{\infty}$ estimates for their gradient in Proposition 4.1 and a local Harnack inequality in Proposition 4.2. As it will be clear in the next sections, the these two results will be essential in Proposition 6.1 in order to establish decay properties at infinity for Riemannian $p$-capacitary potentials in exterior domains and in turn to prove the gradient bounds announced in Theorem 1.2 and Corollary 1.3.
Note that the $L^{p}-L^{\infty}$ estimates for their gradient discussed here is the counterpart of the one obtained in [37] for the horizontal gradient of $p$-harmonic functions on the sub-Riemannian Heisenberg group. The result presented here is essentially contained in [10], although not explicitely stated there. Indeed, in [10] the key estimate concerning a weighted Caccioppoli inequality for the gradient is obtained and here we fill the missing details to derive the claim through uniform local Sobolev inequality from [19] and Moser iteration. Similarly, the uniform local Harnack inequality discussed here is the Riemannian counterpart of the one proved in [11] for the sub-Riemannian case (see also [9] for analogous results concerning the Riemannian approximation for parabolic equations modeled on the $p$-Laplacian for $p\geq 2$). Here, following closely the argument in [11] for the sub-Riemannian case, we detail how to obtain the Harnack inequality for $p$-harmonic functions in the Riemannian case uniformly on $\varepsilon\in(0,1]$, relying on the main results in [19] (see also [9]), i.e., the uniform doubling property and the uniform Poincaré inequalities reviewed in Section 2 in combination with Moser iteration and John-Nirenberg inequality.
Since we were not able to find a reference in the literature and although the arguments are essentially standard and well-known to the experts, for all the three results in the sequel we sketch the proofs mainly for the reader’s convenience, so that uniformity w.r.to $\varepsilon\in(0,1]$ will be transparent and the dependence on $p$ will be made explicit when needed.
Proposition 4.1.
Let $1<p<Q$, $\varepsilon\in(0,1]$, and let $\Omega\subset\mathbb{H}^{n}$ be an open set. There exist constants $C>0$ and $\theta>1$ depending only on $p$ and $Q$ such that for every $\varepsilon\in(0,1]$ and every $v=v^{\varepsilon}_{p}\in\dot{W}^{1,p}_{\varepsilon,loc}(\Omega)$ which is weakly $p$-harmonic in $\Omega$ we have
(4.1) $\||\nabla_{\varepsilon}v|_{\varepsilon}\|_{L^{\infty}(B_{\theta^{-1}r}(\bar{g}))}\leq C\left(\frac{1}{|B_{r}^{\varepsilon}(\bar{g})|}\int_{B_{r}^{\varepsilon}(\bar{g})}|\nabla_{\varepsilon}v|_{\varepsilon}^{p}\right)^{1/p}$
whenever $B_{2r}^{\varepsilon}(\bar{g})\subseteq\Omega$, for some $\bar{g}\in\Omega$ and $r>0$.
Proof.
The proof relies essentially on the uniform Caccioppoli inequality for the gradient of $p$-harmonic functions proved in [10, Theorem 5.3] combined with the with the Moser iteration technique. However, instead of working with the metric balls $B^{\varepsilon}_{r}$, here we prefer to perform the iteration using the balls $\mathcal{B}^{\varepsilon}_{r}$ associated to the $\varepsilon$-gauge $\|\cdot\|_{\varepsilon}$, so that well-behaved cut-off functions are available in view of Lemma 2.2-(5).
For fixed $\varepsilon\in(0,1]$ and $B^{\varepsilon}_{2r}(\bar{g})\subseteq\Omega$ we consider the auxiliary domain $\Omega^{\prime}:=\mathcal{B}^{\varepsilon}_{2\bar{C}^{-1}r}(\bar{g})\subseteq B^{\varepsilon}_{2r}(\bar{g})$, i.e., a ball associated to the $\varepsilon$-gauge $\|\cdot\|_{\varepsilon}$ discussed in Lemma 2.2 for a constant $\bar{C}>1$ depending only on $Q$ obtained there. Thus, chosing $\theta=2\bar{C}^{2}>1$ and setting for brevity $\bar{r}=(2\bar{C})^{-1}r$, Lemma 2.2-(3) yields $B^{\varepsilon}_{\theta^{-1}r}(\bar{g})\subseteq\mathcal{B}^{\varepsilon}_{\bar{r}}(\bar{g})\subseteq\mathcal{B}^{\varepsilon}_{2\bar{r}}(\bar{g})\subseteq B^{\varepsilon}_{r}(\bar{g})$ and the lemma is proved once we show that
(4.2) $\||\nabla_{\varepsilon}v|_{\varepsilon}\|_{L^{\infty}(\mathcal{B}^{\varepsilon}_{\bar{r}}(\bar{g}))}\leq C\left(\frac{1}{|\mathcal{B}_{2\bar{r}}^{\varepsilon}(\bar{g})|}\int_{\mathcal{B}_{2\bar{r}}^{\varepsilon}(\bar{g})}|\nabla_{\varepsilon}v|_{\varepsilon}^{p}\right)^{1/p}\,,$
for some constant $C>0$ depending only on $p$ and $Q$ (note that the volumes of the balls above are all comparable uniformly on $\varepsilon\in(0,1]$ and $r>0$ because of (2.2)).
In order to prove (4.2) we rely on the approach in [10] and proceed by a further regularization of (1.9a) and (1.10), thus for any $\sigma>0$ we consider the Euler-Lagrange equations
(4.3) $\left\{\begin{array}[]{ll}{\rm div}_{\varepsilon}\left(\left(\sigma+|\nabla_{\varepsilon}w|^{2}_{\varepsilon}\right)^{(p-2)/2}\nabla_{\varepsilon}w\right)=0&\textrm{in }\Omega^{\prime},\\ w=v&\textrm{on }\partial\Omega^{\prime}\,,\end{array}\right.$
corresponding to the minimization of the energy functional
(4.4) $E^{\varepsilon,\sigma}_{p}(w)=\frac{1}{p}\int_{\Omega^{\prime}}\left(\sigma+|\nabla_{\varepsilon}w|^{2}_{\varepsilon}\right)^{p/2}\,,$
among Sobolev functions $W^{1,p}_{v,\varepsilon}(\Omega^{\prime})=\left\{w\in W^{1,p}_{\varepsilon}(\Omega^{\prime})\quad\hbox{s.t.}\,\,w=v\,\,\hbox{a.e. on}\,\,\partial\Omega^{\prime}\right\}$. Note that by definition of $\|\cdot\|_{\varepsilon}$ the domain $\Omega^{\prime}$ has piecewise smooth boundary, hence $W^{1,p}_{v,\varepsilon}(\Omega^{\prime})$ is a well defined closed convex subset of $W^{1,p}_{\varepsilon}(\Omega^{\prime})$ in view of standard trace theory for Sobolev functions.
By the direct method in the Calculus of Variations equations (4.3) do have unique solutions $w^{\sigma}=w^{\varepsilon,\sigma}_{p}$ which are minimizers of (4.4) in the class above (here uniqueness follows from the strict convexity of the energy functionals). Thus, $E^{\varepsilon,\sigma}_{p}(w^{\sigma})\leq E^{\varepsilon,\sigma}_{p}(v^{\varepsilon}_{p})\leq C$ uniformly on $\sigma$, so that $w^{\sigma}\to w^{*}$ as $\sigma\to 0$ (possibly up to subsequences) weakly in $W^{1,p}_{\varepsilon}$ and strongly in $L^{p}$ , in particular $w^{*}\in W^{1,p}_{v,\varepsilon}$ by weak continuity of the trace operator. Since $E^{\varepsilon,\sigma}_{p}(v^{\varepsilon}_{p})\to E^{\varepsilon}_{p}(v^{\varepsilon}_{p})$ as $\sigma\to 0$, by weak lower semicontinuity of $E_{p}^{\varepsilon}(\cdot)$ and monotonicity of the map $\sigma\to E^{\varepsilon,\sigma}_{p}(w)$ for any fixed $w$ we conclude that $E^{\varepsilon}_{p}(w^{*})\leq E^{\varepsilon}_{p}(v^{\varepsilon}_{p})$, so that $w^{*}=v^{\varepsilon}_{p}$ because the latter is the unique minimizer of $E^{\varepsilon}$ in the class above again in view of the strict convexity of $E_{p}^{\varepsilon}(\cdot)$. Thus, $E^{\varepsilon}_{p}(w^{\sigma})\to E^{\varepsilon}_{p}(v^{\varepsilon}_{p})$ as $\sigma\to 0$, so that $w^{\sigma}\to v^{\varepsilon}_{p}$ as $\sigma\to 0$ strongly in $W^{1,p}_{\varepsilon}$ and (possibly up to subsequences) a.e. convergence of the gradients holds. Taking into account the previous regularization scheme and its convergence properties as $\sigma\to 0$, inequality (4.2) follows once we prove that
(4.5) $\||\nabla_{\varepsilon}w^{\sigma}|_{\varepsilon}\|_{L^{\infty}(\mathcal{B}^{\varepsilon}_{\bar{r}}(\bar{g}))}\leq C\left(\frac{1}{|\mathcal{B}_{2\bar{r}}^{\varepsilon}(\bar{g})|}\int_{\mathcal{B}_{2\bar{r}}^{\varepsilon}(\bar{g})}\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)^{p/2}\right)^{1/p}\,,$
for some constant $C>0$ depending only on $p$ and $Q$ (in particular, independent of $\sigma$).
In order to obtain (4.5) we rely on [10, Theorem 5.3], so that for any $\beta\geq 2$ and any $\eta\in C^{\infty}_{0}(\Omega^{\prime})$ we have
(4.6) $\int_{\Omega^{\prime}}\eta^{2}\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)^{\frac{p-2+\beta}{2}}|\nabla^{2}_{\varepsilon}w^{\sigma}|^{2}\leq C\beta^{10}\left(\||\nabla_{\varepsilon}\eta|_{\varepsilon}\|_{L^{\infty}}^{2}+\|\eta T\eta\|_{L^{\infty}}\right)\int_{spt\,\eta}\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)^{\frac{p+\beta}{2}}\,,$
where $C>0$ depends only on $p$ and $Q$ and $|\nabla^{2}_{\varepsilon}w^{\sigma}|^{2}$ is the sum of the squared second derivatives w.r. to the orthonormal frame $\{X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{n},T_{\varepsilon}\}$.
Now notice that $\Omega^{\prime}=\mathcal{B}_{4\bar{r}}^{\varepsilon}(\bar{g})$ and chosing $\bar{r}\leq\rho_{1}<\rho_{2}\leq 2\bar{r}\leq 2\rho_{1}\leq 4\bar{r}$ we have $\mathcal{B}_{\rho_{1}}^{\varepsilon}(\bar{g})\subset\mathcal{B}_{\rho_{2}}^{\varepsilon}(\bar{g})\subseteq\Omega^{\prime}$, so that by a simple limiting argument inequality (4.6) still holds for $\eta=\zeta$, the Lipschitz cut-off function given in Lemma 2.2-(5). Thus, combining (4.6) with the elementary inequality $|\nabla_{\varepsilon}\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)|^{2}_{\varepsilon}\leq 4\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)|\nabla^{2}_{\varepsilon}w^{\sigma}|^{2}$ we obtain for any $\beta\geq 2$
$\int_{\mathcal{B}_{\rho_{2}}^{\varepsilon}(\bar{g})}\zeta^{2}\left|\nabla_{\varepsilon}\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)^{\frac{p+\beta}{4}}\right|_{\varepsilon}^{2}\leq\frac{C(1+\beta)^{12}}{(\rho_{2}-\rho_{1})^{2}}\int_{\mathcal{B}_{\rho_{2}}^{\varepsilon}(\bar{g})}\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)^{\frac{p+\beta}{2}}\,.$
On the other hand still by Lemma 2.2-(5) we have for any $\beta\geq 2$
$\int_{\mathcal{B}_{\rho_{2}}^{\varepsilon}(\bar{g})}|\nabla_{\varepsilon}\zeta|_{\varepsilon}^{2}\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)^{\frac{p+\beta}{2}}\leq\frac{C}{(\rho_{2}-\rho_{1})^{2}}\int_{\mathcal{B}_{\rho_{2}}^{\varepsilon}(\bar{g})}\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)^{\frac{p+\beta}{2}}\,,$
which combined with the previous inequality and Leibnitz’s rule easily yields for any $\beta\geq 2$
(4.7) $\int_{\mathcal{B}_{\rho_{2}}^{\varepsilon}(\bar{g})}\left|\nabla_{\varepsilon}\left\{\zeta\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)^{\frac{p+\beta}{4}}\right\}\right|_{\varepsilon}^{2}\leq\frac{C(1+\beta)^{12}}{(\rho_{2}-\rho_{1})^{2}}\int_{\mathcal{B}_{\rho_{2}}^{\varepsilon}(\bar{g})}\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)^{\frac{p+\beta}{2}}\,.$
From now on we set for brevity $\Psi_{\sigma}=\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)^{\frac{1}{2}}$. Since $spt\,\zeta\subset\mathcal{B}_{\rho_{2}}^{\varepsilon}(\bar{g})\subseteq\Omega^{\prime}\subseteq B^{\varepsilon}_{2r}(\bar{g})$, from the localized Sobolev inequality (2.5) (with $p=2$ and on the ball $B^{\varepsilon}_{2r}(\bar{g})$) we infer for any $\beta\geq 2$
$\left(\frac{1}{|B^{\varepsilon}_{2r}(\bar{g})|}\int_{B^{\varepsilon}_{2r}(\bar{g})}\left|\zeta\Psi_{\sigma}^{\frac{p+\beta}{2}}\right|^{\frac{2Q}{Q-2}}\right)^{\frac{Q-2}{2Q}}\leq 2rC\,\left(\frac{1}{|B^{\varepsilon}_{2r}(\bar{g})|}\int_{B^{\varepsilon}_{2r}(\bar{g})}\left|\nabla_{\varepsilon}\left\{\zeta\Psi_{\sigma}^{\frac{p+\beta}{2}}\right\}\right|_{\varepsilon}^{2}\right)^{1/2}\,\,,$
where the constant $C>0$ depends only on $p$ and $Q$.
Now we observe that $\mathcal{B}_{\bar{r}}^{\varepsilon}(\bar{g})\subset\mathcal{B}^{\varepsilon}_{\rho_{1}}(\bar{g})\subset spt\,\zeta\subset\mathcal{B}_{\rho_{2}}^{\varepsilon}(\bar{g})\subseteq\Omega^{\prime}\subseteq B^{\varepsilon}_{2r}(\bar{g})=B^{\varepsilon}_{4\bar{C}\bar{r}}(\bar{g})$, hence all the balls have uniformly equivalent volumes in view of (2.2). Thus, combining the last inequality with (4.7) and taking into account the properties of $\zeta$, for any $\beta\geq 2$ and $\bar{r}\leq\rho_{1}<\rho_{2}\leq 2\bar{r}\leq 2\rho_{1}\leq 4\bar{r}$ we finally get
(4.8) $\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{\rho_{1}}(\bar{g})|}\int_{\mathcal{B}^{\varepsilon}_{\rho_{1}}(\bar{g})}\left(\Psi_{\sigma}^{p+\beta}\right)^{\frac{Q}{Q-2}}\right)^{\frac{Q-2}{Q}}\leq\frac{C\rho_{2}^{2}(1+\beta)^{12}}{(\rho_{2}-\rho_{1})^{2}}\frac{1}{|\mathcal{B}^{\varepsilon}_{\rho_{2}}(\bar{g})|}\int_{\mathcal{B}_{\rho_{2}}^{\varepsilon}(\bar{g})}\Psi_{\sigma}^{p+\beta}\,\,,$
where the constant $C>0$ depends only on $p$ and $Q$.
Next we perform Moser iteration, i.e., for $i\in\mathbb{N}$ we repeatedly use (4.8) with the choice $(\rho_{1},\rho_{2},\beta)=(\tilde{\rho}_{i+1},\tilde{\rho}_{i},\beta_{i})$, where $\tilde{\rho}_{i}=\bar{r}(1+2^{-i-1})$ for $i\geq 0$ and $\beta_{i+1}=\frac{Q}{Q-2}\beta_{i}+\frac{2p}{Q-2}$, $\beta_{0}=2$. Note that $\{\beta_{i}\}$ is positive, increasing and $\beta_{i}\to+\infty$ as $i\to\infty$, moreover $p+\beta_{i+1}=(p+\beta_{i})\frac{Q}{Q-2}$ and $\rho_{2}-\rho_{1}=\tilde{\rho}_{i}-\tilde{\rho}_{i+1}=\bar{r}2^{-i-2}$ for any $i\geq 0$. Thus, under the previous choice (4.8) leads to
(4.9) $\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{\tilde{\rho}_{i+1}}(\bar{g})|}\int_{\mathcal{B}^{\varepsilon}_{\tilde{\rho}_{i+1}}(\bar{g})}\Psi_{\sigma}^{p+\beta_{i+1}}\right)^{\frac{1}{p+\beta_{i+1}}}\leq\left(4^{i}C(p+\beta_{i})^{12}\right)^{\frac{1}{p+\beta_{i}}}\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{\tilde{\rho}_{i}}(\bar{g})|}\int_{\mathcal{B}_{\tilde{\rho}_{i}}^{\varepsilon}(\bar{g})}\Psi_{\sigma}^{p+\beta_{i}}\right)^{\frac{1}{p+\beta_{i}}}\,\,,$
whence the pointwise bound $|\nabla_{\varepsilon}w^{\sigma}|_{\varepsilon}\leq\Psi_{\sigma}$, the inclusion $\mathcal{B}^{\varepsilon}_{\bar{r}}(\bar{g})\subset\mathcal{B}^{\varepsilon}_{\tilde{\rho}_{i+1}}(\bar{g})\subset\mathcal{B}^{\varepsilon}_{2\bar{r}}(\bar{g})$ with uniform equivalence of volumes, and a simple iteration of (4.9) give
(4.10) $\||\nabla_{\varepsilon}w^{\sigma}|_{\varepsilon}\|_{L^{\infty}(\mathcal{B}^{\varepsilon}_{\bar{r}}(\bar{g}))}$ $=\lim_{i\to\infty}\left(\frac{1}{\mathcal{B}^{\varepsilon}_{\bar{r}}(\bar{g})}\int_{\mathcal{B}^{\varepsilon}_{\bar{r}}(\bar{g})}|\nabla_{\varepsilon}w^{\sigma}|_{\varepsilon}^{p+\beta_{i+1}}\right)^{\frac{1}{p+\beta_{i+1}}}$
$\leq\hat{C}\left(\frac{1}{|\mathcal{B}_{\frac{3}{2}\bar{r}}^{\varepsilon}(\bar{g})|}\int_{\mathcal{B}_{\frac{3}{2}\bar{r}}^{\varepsilon}(\bar{g})}\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)^{\frac{p+2}{2}}\right)^{\frac{1}{p+2}}\,,$
where $\hat{C}=\Pi_{i=0}^{\infty}\left(4^{i}C(p+\beta_{i})^{12}\right)^{\frac{1}{p+\beta_{i}}}<\infty$ depends only on $p$ and $Q$ and in particular it is independent of $\sigma$ (notice that by its very definition $\beta_{i}\geq\left(\frac{Q}{Q-2}\right)^{i}$ for any $i\geq 0$, whence $\sum_{i=0}^{\infty}\frac{\log(p+\beta_{i})}{p+\beta_{i}}<\infty$ and in turn the convergence of the infinite product follows easily).
In order to conclude the proof it remains to show how to infer (4.5) from (4.10), i.e., how to decrease the exponent from $p+2$ to $p$ on the r.h.s. of (4.10) (while inflating slightly the corresponding ball), and to do this we adapt the interpolation argument in [27, Lemma 3.38]. First, using (4.8) with $\beta=2$ for any $\bar{r}\leq\rho_{1}<\rho_{2}\leq 2\bar{r}\leq 2\rho_{1}\leq 4\bar{r}$, so that $\lambda=\rho_{1}/\rho_{2}\in[\frac{\bar{r}}{\rho_{2}},1)\subseteq[\frac{1}{2},1)$, we deduce that the following inequality holds
(4.11) $\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{\lambda\rho_{2}}(\bar{g})|}\int_{\mathcal{B}^{\varepsilon}_{\lambda\rho_{2}}(\bar{g})}\left(\Psi_{\sigma}^{p+2}\right)^{\frac{Q}{Q-2}}\right)^{\frac{Q-2}{Q(p+2)}}\leq\frac{C}{(1-\lambda)^{\frac{2}{p+2}}}\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{\rho_{2}}(\bar{g})|}\int_{\mathcal{B}_{\rho_{2}}^{\varepsilon}(\bar{g})}\Psi_{\sigma}^{p+2}\,\right)^{\frac{1}{p+2}}\,,$
where the constant $C>0$ depends only on $p$ and $Q$.
Next we optimize (4.11) with respect to $\lambda\in[\frac{\bar{r}}{\rho_{2}},1)$ while keeping $\rho_{2}\in[\bar{r},2\bar{r}]$ fixed. Since $p<p+2<(p+2)\frac{Q}{Q-2}$, if we choose $\mu\in(0,1)$ such that
(4.12) $\frac{1}{\mu}-1=\frac{Q}{p}\,,\quad\hbox{i.e.}\,,\qquad\frac{1}{p+2}=\frac{\mu}{p}+\frac{1-\mu}{(p+2)\frac{Q}{Q-2}}\,,$
then in view of Hölder’s inequalty, (4.11) and $1<p<Q$ we have for fixed $\rho_{2}\in[\bar{r},2\bar{r}]$
(4.13) $\Phi(\rho_{2};p):=\sup_{\lambda\in[\frac{\bar{r}}{\rho_{2}},1)}(1-\lambda)^{\frac{2(1-\mu)}{\mu(p+2)}}\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{\lambda\rho_{2}}(\bar{g})|}\int_{\mathcal{B}_{\lambda\rho_{2}}^{\varepsilon}(\bar{g})}\Psi_{\sigma}^{p+2}\right)^{\frac{1}{p+2}}\leq C\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{\rho_{2}}(\bar{g})|}\int_{\mathcal{B}_{\rho_{2}}^{\varepsilon}(\bar{g})}\Psi_{\sigma}^{p+2}\,\right)^{\frac{1}{p+2}}\,,$
where the constant $C>0$ depends only on $p$ and $Q$.
For given $\lambda\in[\frac{\bar{r}}{\rho_{2}},1)\subseteq[\frac{1}{2},1)$ we set $\lambda^{\prime}=\frac{1}{2}(1+\lambda)\in(\lambda,1)\subset[\frac{1}{2},1)$, so that $\lambda^{\prime}$ is admissible in (4.13) and $\bar{r}\leq\rho_{1}=\lambda\rho_{2}<\lambda^{\prime}\rho_{2}<2\bar{r}\leq 2\rho_{1}<4\bar{r}$, so that (4.11) also holds with radii $(\lambda\rho_{2},\lambda^{\prime}\rho_{2})$ instead of $(\lambda\rho_{2},\rho_{2})$. Note that $1-\lambda/\lambda^{\prime}=(1-\lambda^{\prime})/\lambda^{\prime}=\frac{1-\lambda}{2\lambda^{\prime}}$, so that $1-\lambda/\lambda^{\prime}$ and $1-\lambda$ are uniformly comparable to $1-\lambda^{\prime}$ for $\lambda\in[\frac{1}{2},1)$, hence combining the previous two inequalities with (4.12) we have the estimate
(4.14) $(1-\lambda)^{\frac{2}{\mu(p+2)}}\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{\lambda\rho_{2}}(\bar{g})|}\int_{\mathcal{B}^{\varepsilon}_{\lambda\rho_{2}}(\bar{g})}\left(\Psi_{\sigma}^{p+2}\right)^{\frac{Q}{Q-2}}\right)^{\frac{Q-2}{Q(p+2)}}\leq C\frac{(1-\lambda)^{\frac{2}{\mu(p+2)}}}{(1-\lambda/\lambda^{\prime})^{\frac{2}{p+2}}}\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{\lambda^{\prime}\rho_{2}}(\bar{g})|}\int_{\mathcal{B}_{\lambda^{\prime}\rho_{2}}^{\varepsilon}(\bar{g})}\Psi_{\sigma}^{p+2}\right)^{\frac{1}{p+2}}\leq\widetilde{C}\Phi(\rho_{2};p)\,,$
where $\widetilde{C}>0$ depends only on $p$ and $Q$ and $\lambda\in[\frac{\bar{r}}{\rho_{2}},1)$ is arbitrary.
Now for any $\delta>0$ we chose $\tilde{\lambda}\in[\frac{\bar{r}}{\rho_{2}},1)$ in (4.13) so that in view also of (4.12) and Hölder inequality we have
$\Phi(\rho_{2};p)<(1-\tilde{\lambda})^{\frac{2(1-\mu)}{\mu(p+2)}}\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{\tilde{\lambda}\rho_{2}}(\bar{g})|}\int_{\mathcal{B}_{\tilde{\lambda}\rho_{2}}^{\varepsilon}(\bar{g})}\Psi_{\sigma}^{\mu(p+2)}\Psi_{\sigma}^{(1-\mu)(p+2)}\right)^{\frac{1}{p+2}}+\delta\,,$
$\leq(1-\tilde{\lambda})^{\frac{2(1-\mu)}{\mu(p+2)}}\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{\tilde{\lambda}\rho_{2}}(\bar{g})|}\int_{\mathcal{B}_{\tilde{\lambda}\rho_{2}}^{\varepsilon}(\bar{g})}\Psi_{\sigma}^{p}\right)^{\frac{\mu}{p}}\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{\tilde{\lambda}\rho_{2}}(\bar{g})|}\int_{\mathcal{B}^{\varepsilon}_{\tilde{\lambda}\rho_{2}}(\bar{g})}\left(\Psi_{\sigma}^{p+2}\right)^{\frac{Q}{Q-2}}\right)^{\frac{Q-2}{Q(p+2)}(1-\mu)}+\delta\,.$
Next, for $\nu>0$ to be chosen later and $a,b\geq 0$ we use Young’s inequality $ab\leq c(\nu,\mu)a^{1/\mu}+\nu b^{1/(1-\mu)}$ to infer from the last inequality that
$\Phi(\rho_{2};p)<c(\nu,\mu)\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{\tilde{\lambda}\rho_{2}}(\bar{g})|}\int_{\mathcal{B}_{\tilde{\lambda}\rho_{2}}^{\varepsilon}(\bar{g})}\Psi_{\sigma}^{p}\right)^{\frac{1}{p}}+\nu(1-\tilde{\lambda})^{\frac{2}{\mu(p+2)}}\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{\tilde{\lambda}\rho_{2}}(\bar{g})|}\int_{\mathcal{B}^{\varepsilon}_{\tilde{\lambda}\rho_{2}}(\bar{g})}\left(\Psi_{\sigma}^{p+2}\right)^{\frac{Q}{Q-2}}\right)^{\frac{Q-2}{Q(p+2)}}+\delta\,,$
whence equivalence of volumes between $\mathcal{B}^{\varepsilon}_{\tilde{\lambda}\rho_{2}}(\bar{g})$ and $\mathcal{B}^{\varepsilon}_{\rho_{2}}(\bar{g})$ for the first term together with (4.14) for the second one yield
(4.15) $\Phi(\rho_{2};p)<c(\nu,\mu,p,Q)\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{\rho_{2}}(\bar{g})|}\int_{\mathcal{B}_{\rho_{2}}^{\varepsilon}(\bar{g})}\Psi_{\sigma}^{p}\right)^{\frac{1}{p}}+\nu\,\widetilde{C}\,\Phi(\rho_{2};p)+\delta\,.$
Finally, chosing $\nu=\frac{1}{2\widetilde{C}}$ and letting $\delta\to 0$ in (4.15) we get $\Phi(\rho_{2};p)\leq 2c(\nu,\mu,p,Q)\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{\rho_{2}}(\bar{g})|}\int_{\mathcal{B}_{\rho_{2}}^{\varepsilon}(\bar{g})}\Psi_{\sigma}^{p}\right)^{\frac{1}{p}}$, which together with (4.13) for $\rho_{1}=\frac{3}{2}\bar{r}$, $\rho_{2}=2\bar{r}$ and $\lambda=\frac{\rho_{1}}{\rho_{2}}=\frac{3}{4}>\frac{1}{2}$ give
$\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{\frac{3}{2}\bar{r}}(\bar{g})|}\int_{\mathcal{B}_{\frac{3}{2}\bar{r}}^{\varepsilon}(\bar{g})}\Psi_{\sigma}^{p+2}\right)^{\frac{1}{p+2}}\leq C\Phi(p)\leq C\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{2\bar{r}}(\bar{g})|}\int_{\mathcal{B}_{2\bar{r}}^{\varepsilon}(\bar{g})}\Psi_{\sigma}^{p}\right)^{\frac{1}{p}}\,,$
for constants $C>0$ depending only on $p$ and $Q$. Combining the last inequality with (4.10) the estimate (4.5) follows and the proof is complete.
The second tool needed in the next sections is the following uniform Harnack inequality for $p$-harmonic functions.
Proposition 4.2.
Let $1<p<Q$ and $\bar{r}>0$ be fixed and for any $\varepsilon>0$ and $\bar{g}\in\mathbb{H}^{n}$ let $v=v^{\varepsilon}_{p}\in C^{1}(B^{\varepsilon}_{9\bar{r}}(\bar{g}))$ be a nonnegative $p$-harmonic function w.r. to the metric $|\cdot|^{2}_{\varepsilon}$. There exists a constant $C_{H}\geq 1$ depending only on $p,Q$ and $\bar{r}$ such that
(4.16) $\underset{B^{\varepsilon}_{r}(\bar{g})}{\rm{ess}\sup}\,\,v\leq C_{H}\,\underset{B^{\varepsilon}_{r}(\bar{g})}{\rm{ess}\inf}\,\,v\qquad\hbox{for any }0<r\leq\frac{1}{2}\bar{r}\,.$
Proof.
As already announced, we follow closely the strategy in [11], although with some simplifications in treating the two involved $L^{\infty}-L^{p}$ bounds through Moser iteration in a unified way and without relying on the Dahlberg and Kenig trick (compare [11, Lemma 3.29]). Moreover, relying on Lemma 2.3, we can apply the abstract John-Nirenberg inequality for BMO functions on homogeneous spaces from [5] and infer (4.16) from the $L^{\infty}-L^{p}$ bounds.
We first notice that $v>0$ in $B^{\varepsilon}_{9\bar{r}}(\bar{g})$ by the strong maximum principle, unless $v\equiv 0$ and (4.16) holds trivially. Now we derive a Caccioppoli-type inequality for arbitrary powers of $v$. Thus, for fixed $\varepsilon>0$, $\beta\in\mathbb{R}\setminus\{-1/p\}$ and for nonnegative $d_{\varepsilon}$-Lipschitz function $\zeta$ compactly supported in $B^{\varepsilon}_{9\bar{r}}(\bar{g})$ to be specified later, we test the equation for $v$ with $\zeta^{p}v^{1+\beta p}$, hence integration by parts, Cauchy-Schwarz and Young inequalities yield
$\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}\zeta^{p}|\nabla_{\varepsilon}v|_{\varepsilon}^{p}v^{\beta p}=-\frac{1}{1+\beta p}\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}|\nabla_{\varepsilon}v|_{\varepsilon}^{p-2}v^{1+\beta p}\langle\nabla_{\varepsilon}v,\nabla_{\varepsilon}\zeta^{p}\rangle_{\varepsilon}$
$\leq\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}\left(\zeta^{p-1}|\nabla_{\varepsilon}v|_{\varepsilon}^{p-1}v^{\beta(p-1)}\right)\left(\frac{p}{|1+\beta p|}v^{1+\beta}|\nabla_{\varepsilon}\zeta|_{\varepsilon}\right)$
$\leq\left(1-\frac{1}{p}\right)\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}\zeta^{p}|\nabla_{\varepsilon}v|_{\varepsilon}^{p}v^{\beta p}+\frac{1}{p}\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}\left(\frac{p}{|1+\beta p|}\right)^{p}v^{(1+\beta)p}|\nabla_{\varepsilon}\zeta|^{p}_{\varepsilon}\,.$
Thus, for $\beta=-1$ the previous inequality in lerms of $\log v$ becomes
(4.17) $\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}\zeta^{p}|\nabla_{\varepsilon}\log v|_{\varepsilon}^{p}\leq\left(\frac{p}{p-1}\right)^{p}\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}|\nabla_{\varepsilon}\zeta|^{p}_{\varepsilon}\,,$
while for $\beta\neq-1$ in terms of $v^{1+\beta}$ we have
$\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}\zeta^{p}|\nabla_{\varepsilon}v^{1+\beta}|_{\varepsilon}^{p}\leq\left(\frac{|1+\beta|p}{|1+\beta p|}\right)^{p}\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}\left(v^{(1+\beta)}|\nabla_{\varepsilon}\zeta|_{\varepsilon}\right)^{p}\,,$
which combined with Leibniz rule and a convexity inequality gives
(4.18) $\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}|\nabla_{\varepsilon}\left(\zeta v^{1+\beta}\right)|_{\varepsilon}^{p}\leq 2^{p-1}\left[\left(\frac{|1+\beta|p}{|1+\beta p|}\right)^{p}+1\right]\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}\left(v^{(1+\beta)}|\nabla_{\varepsilon}\zeta|_{\varepsilon}\right)^{p}\,.$
Next, we specialize inequalities (4.17) and (4.18) by chosing test functions adapted to the metric $d_{\varepsilon}$. Instead of resorting to the general theory of spaces of homogenous type as in, e.g., [9, Section 7], here we argue as in Lemma 2.2. Notice that $|\nabla_{\varepsilon}d_{\varepsilon}(\cdot,\bar{g})|_{\varepsilon}\leq 1$ on $\mathbb{H}^{n}$ for any $\bar{g}\in\mathbb{H}^{n}$ and for each $\varepsilon>0$ because such function is $1$-Lipschitz w.r.to $d_{\varepsilon}$ just by triangle inequality. Thus, if $0<\rho_{1}<\rho_{2}\leq 2\rho_{1}$ and $h:[0,\infty)\to[0,1]$ is a Lipschitz and piecewhise linear function, with $\{h\equiv 1\}=[0,\rho_{1}]$, $\{h\equiv 0\}=[\frac{\rho_{1}+\rho_{2}}{2},\infty)$, and $h$ is linear otherwise, then for any $\bar{g}\in\mathbb{H}^{n}$ the function $\zeta(g)=h(d_{\varepsilon}(g,\bar{g}))$ is Lipschitz on $\mathbb{H}^{n}$, with $\zeta\equiv 1$ on $B^{\varepsilon}_{\rho_{1}}(\bar{g})$, $spt\,\zeta\subset B^{\varepsilon}_{\rho_{2}}(\bar{g})$ and $|\nabla_{\varepsilon}\zeta|_{\varepsilon}\leq\frac{2}{\rho_{2}-\rho_{1}}\,$.
Thus, for $\hat{g}\in B^{\varepsilon}_{\bar{r}}(\bar{g})$ and $0<\rho\leq 2\bar{r}$ we have $B^{\varepsilon}_{3\rho}(\hat{g})\subset B^{\varepsilon}_{7\bar{r}}(\bar{g})$ and $B^{\varepsilon}_{4\rho}(\hat{g})\subset B^{\varepsilon}_{9\bar{r}}(\bar{g})$, hence for $(\rho_{1},\rho_{2})=(3\rho,4\rho)$ in view of the properties of $\zeta$ inequality (4.17) together with the doubling property (2.2) and the uniform Poincaré inequality (2.3) yield for $w=\log v$
(4.19) $\frac{1}{|B^{\varepsilon}_{\rho}(\hat{g})|}\int_{B^{\varepsilon}_{\rho}(\hat{g})}|w-w_{B^{\varepsilon}_{\rho}(\hat{g})}|\leq C\frac{\rho}{|B^{\varepsilon}_{3\rho}(\hat{g})|}\int_{B^{\varepsilon}_{3\rho}(\hat{g})}|\nabla_{\varepsilon}w|_{\varepsilon}\leq C\rho\left(\frac{1}{|B^{\varepsilon}_{4\rho}(\hat{g})|}\int_{B^{\varepsilon}_{4\rho}(\hat{g})}\zeta^{p}|\nabla_{\varepsilon}w|^{p}_{\varepsilon}\right)^{1/p}\leq C\,,$
where $C>0$ depends only on $p$ and $Q$. Then, setting $\widetilde{B}^{\varepsilon}_{\rho}(\hat{g})=B^{\varepsilon}_{\rho}(\hat{g})\cap B^{\varepsilon}_{\bar{r}}(\bar{g})$ and applying inequality (2.8) we easily infer from (4.19) and the triangle inequality that for $\hat{g}\in B^{\varepsilon}_{\bar{r}}(\bar{g})$ and $0<\rho\leq 2\bar{r}$ we have
(4.20) $\frac{1}{|\widetilde{B}^{\varepsilon}_{\rho}(\hat{g})|}\int_{\widetilde{B}^{\varepsilon}_{\rho}(\hat{g})}|w-w_{\widetilde{B}^{\varepsilon}_{\rho}(\hat{g})}|$ $\leq C\frac{1}{|B^{\varepsilon}_{\rho}(\hat{g})|^{2}}\int_{B^{\varepsilon}_{\rho}(\hat{g})}\int_{B^{\varepsilon}_{\rho}(\hat{g})}|w(g)-w(g^{\prime})|$
$\leq C\frac{1}{|B^{\varepsilon}_{\rho}(\hat{g})|}\int_{B^{\varepsilon}_{\rho}(\hat{g})}|w-w_{B^{\varepsilon}_{\rho}(\hat{g})}|\leq C,$
where $C>0$ depends only on $p$ and $Q$. In particular, the last estimate shows that $w\in BMO(B^{\varepsilon}_{\bar{r}}(\bar{g}),d_{\varepsilon},|\cdot|)$ as the ball $\widetilde{B}^{\varepsilon}_{\rho}(\hat{g})\subset B^{\varepsilon}_{\bar{r}}(\bar{g})$ is an arbitrary ball in the metric space $(B^{\varepsilon}_{\bar{r}}(\bar{g}),d_{\varepsilon})$.
Note that for all $\varepsilon>0$ the spaces $(B^{\varepsilon}_{\bar{r}}(\bar{g}),d_{\varepsilon},|\cdot|)$ are of homogeneous type, the doubling property for them holds uniformly w.r.to $\varepsilon>0$ due to (2.2) together with (2.8). Then, it follows from the bound of the BMO norms (4.20) and the abstract John-Nirenberg inequality (see, e.g, [5, Theorem 1]) that there exists $\hat{p}>0$ and for each $\bar{p}\in(0,\hat{p})$ there exists a corresponding $C=C(\bar{p})>0$ both dependent on $p$ and $Q$, possibly depending on $\bar{r}$ but independent of $\varepsilon$, such that $\int_{B^{\varepsilon}_{r}(\bar{g})}exp\left(\bar{p}|w-w_{B^{\varepsilon}_{r}(\bar{g})}|\right)\leq C|B^{\varepsilon}_{r}(\bar{g})|$ for any $0<r\leq\bar{r}$. Then, the identity $w=\log v$ and $e^{\pm\bar{p}\log v}=v^{\pm\bar{p}}$ together with exponential integrabiliy readily yield for any $0<r\leq\bar{r}$ the estimate
(4.21) $\left(\frac{1}{|B^{\varepsilon}_{r}(\bar{g})|}\int_{B^{\varepsilon}_{r}(\bar{g})}v^{\bar{p}}\right)^{1/\bar{p}}\leq C\left(\frac{1}{|B^{\varepsilon}_{r}(\bar{g})|}\int_{B^{\varepsilon}_{r}(\bar{g})}v^{-\bar{p}}\right)^{-1/\bar{p}}\,,$
where $C>0$ depends only on $\bar{p}$, $p$, $Q$ and $\bar{r}$.
In order to proceed in the proof we go back to (4.18) and use the localized Sobolev inequality for fixed $0<r\leq\bar{r}$. Taking the cut-off function $\zeta$ as above, since $spt\,\zeta\subset B_{\rho_{2}}^{\varepsilon}(\bar{g})\subseteq B^{\varepsilon}_{r}(\bar{g})$, from inequality (2.5) and (2.2) we infer for any $\beta\in\mathbb{R}\setminus\{-1,-1/p\}$ and $r/4\leq\rho_{1}<\rho_{2}\leq r$ that
(4.22) $\left(\frac{1}{|B^{\varepsilon}_{\rho_{1}}(\bar{g})|}\int_{B^{\varepsilon}_{\rho_{1}}(\bar{g})}\left(v^{(1+\beta)}\right)^{\frac{pQ}{(Q-p)}}\right)^{\frac{Q-p}{pQ|1+\beta|}}$
$\leq\left(\frac{C\rho_{2}}{\rho_{2}-\rho_{1}}\right)^{\frac{1}{|1+\beta|}}\left[\left(\frac{|1+\beta|p}{|1+\beta p|}\right)^{p}+1\right]^{\frac{1}{p|\beta+1|}}\left(\frac{1}{|B^{\varepsilon}_{\rho_{2}(\bar{g})}|}\int_{B^{\varepsilon}_{\rho_{2}}(\bar{g})}\left(v^{(1+\beta)}\right)^{p}\right)^{\frac{1}{p|\beta+1|}},$
where the constant $C>0$ depends only on $p$ and $Q$.
Next we perform Moser iteration with exponents $\{\beta_{i}\}_{i\geq 0}$ where $\beta_{0}\in(-1,-1/p)\cup(-1/p,\infty)$ is to be chosen later and $\beta_{i+1}=p/(Q-p)+\beta_{i}Q/(Q-p)$, so that $\{\beta_{i}\}\subset(-1,\infty)$ and it is increasing. Note that if $s=p(\beta+1)$ then $\beta>-1$ iff $s>0$ and $\beta\neq-1/p$ iff $s\neq p-1$. Hence, setting for brevity $s_{i}=p(\beta_{i}+1)$, we have that $s_{i+1}=s_{i}Q/(Q-p)=s_{0}\frac{Q^{i+1}}{(Q-p)^{i+1}}\to+\infty$ as $i\to\infty$, so that the sequence $\{\beta_{i}\}$ is eventually nonnegative and increasing to +$\infty$+ ∞ as $\beta_{i+1}\geq C\frac{Q^{i}}{(Q-1)^{i}}$ for any $i\geq 0$ large enough. Moreover, for such positive $\beta=\beta_{i}$ one has $\left(\frac{|1+\beta|p}{|1+\beta p|}\right)^{p}+1\leq 2p^{p}$ Thus, we can select $s_{0}=\bar{p}\in(0,\hat{p})$ such that $s_{i}\neq p-1$ for all $i\geq 0$ (i.e., $\beta_{i}\in(-1,-1/p)\cup(-1/p,\infty)$ for all $i\geq 0$).
Thus, for fixed $0<r\leq\bar{r}$ we repeatedly use (4.22) with the choice $(\rho_{1},\rho_{2},\beta)=(\tilde{\rho}_{i+1},\tilde{\rho}_{i},\beta_{i})$, where $\tilde{\rho}_{i}=r(\frac{1}{2}+2^{-i-1})$ for $i\geq 0$, so that inequality(4.8) rewrites for all $i\geq 0$ as
(4.23) $\left(\frac{1}{|B^{\varepsilon}_{\tilde{\rho}_{i+1}}(\bar{g})|}\int_{B^{\varepsilon}_{\tilde{\rho}_{i+1}}(\bar{g})}v^{s_{i+1}}\right)^{\frac{1}{|s_{i+1}|}}\leq C^{\frac{i+1}{|s_{i}|}}\left(\frac{1}{|B^{\varepsilon}_{\tilde{\rho}_{i}}(\bar{g})|}\int_{B_{\tilde{\rho}_{i}}^{\varepsilon}(\bar{g})}v^{s_{i}}\right)^{\frac{1}{|s_{i}|}}\,\,,$
for a constant $C>0$ depending only on $p$ and $Q$.
Iterating (4.23) as in the proof of the previous lemma we easily obtain the inequality
(4.24) $\|v\|_{L^{\infty}(B^{\varepsilon}_{r/2}(\bar{g}))}\leq C^{\prime}\left(\frac{1}{|B_{r}^{\varepsilon}(\bar{g})|}\int_{B_{r}^{\varepsilon}(\bar{g})}v^{\bar{p}}\right)^{\frac{1}{\bar{p}}}\,,$
where $C^{\prime}=\Pi_{i=0}^{\infty}C^{\frac{i+1}{|s_{i}|}}<\infty$ depends only on $p$ and $Q$ (notice that $s_{i}=s_{0}\frac{Q^{i}}{(Q-p)^{i}}$ for any $i\geq 0$, whence $\sum_{i=0}^{\infty}\frac{i+1}{s_{i}}<\infty$ and in turn the convergence of the infinite product follows).
On the other hand, for fixed $0<r\leq\bar{r}$ we perform again Moser iteration to show that
(4.25) $\left(\underset{B^{\varepsilon}_{r/2}(\bar{g})}{\rm{ess}\inf}\,\,v\right)^{-1}=\left\|v^{-1}\right\|_{L^{\infty}(B^{\varepsilon}_{r/2})}\leq\widetilde{C}\left(\frac{1}{|B^{\varepsilon}_{r}(\bar{g})|}\int_{B^{\varepsilon}_{r}(\bar{g})}v^{-\bar{p}}\right)^{1/\bar{p}}\,,$
where $\widetilde{C}>0$ depends only on $\bar{p}$, $Q$ and $\bar{r}$. Indeed, note that for $\{\beta_{i}\}_{i\geq 0}$ and $\{s_{i}\}_{i\geq 0}$ as above and $\beta_{0}<-1$ one has $\beta_{i}<-1$ and in turn $\beta_{i+1}<\beta_{i}$ for all $i\geq 0$. On the other hand for $s_{0}<0$ one has both $s_{i}=s_{0}\left(\frac{Q}{Q-p}\right)^{i}\to-\infty$ and in turn $\beta_{i}\to-\infty$ as $i\to\infty$. Then, we can chose $\beta_{0}<-1$ such that $s_{0}=p(1+\beta_{0})=-\bar{p}$ and as above we can repeatedly use (4.22) with the same choice $(\rho_{1},\rho_{2},\beta)=(\tilde{\rho}_{i+1},\tilde{\rho}_{i},\beta_{i})$ and where still $\tilde{\rho}_{i}=r(\frac{1}{2}+2^{-i-1})$ for $i\geq 0$ to get once more (4.23) for all $i\geq 0$. As in the proof of (4.24), iterating (4.23) for these new set of exponents inequality (4.25) follows easily as $i\to\infty$
Finally, combining inequalities (4.24), (4.21), and (4.25) the desired inequality (4.16) follows with $C_{H}>0$ depending only on $p$, $Q$ and $\bar{r}$. ∎
5. Two-sided control at infinity for $p$-capacitary potentials.
In this section, we consider $p$-capacitary potentials associated to exterior domains, the model case being the complement of a Korányi ball, and we discuss their two-sided behaviour at infinity. On the one hand, we are going to establish a sharp universal pointwise lower bound through an iterative use of the barriers constructed in Section 3. This lower bound will be relevant to obtain for the limit $u$ of solutions to (1.6a)-(1.6b)-(1.8)-(1.9a)-(1.9b) as $p\to 1$ a lower bound for the size of their level sets of the form $diam_{\varepsilon}(\{u=s\})\gtrsim e^{s/(Q-1)}$.
Concerning the upper bound, we first obtain it in the form of integrability at infinity below the Sobolev exponent $p^{*}$ in the space $L^{\sigma,\infty}$, where $\sigma=\frac{Q(p-1)}{Q-p}$ is the natural exponent dictated by the explicit singular solutions provided in Lemma 3.1 , from which higher integrability follows by interpolation. Then, as a byproduct of our analysis performed in Proposition 4.2, in Proposition 5.4 we derive a Harnack inequality for $p$-harmonic functions on annular regions which holds uniformly on $\varepsilon\in(0,1]$ and with constants of controlled growth as $p\to 1$. As a consequence, for $p>1$ we provide a power-type decay of $p$-capacitary potentials of the form $v^{\varepsilon}_{p}(g)\lesssim\|g\|^{\frac{p-Q}{p-1}}$ as $\|g\|\to\infty$ and with constant of controlled growth as $p\to 1$. In turn, this pointwise behaviour of the potentials guarantee that for solutions $u^{\varepsilon}=\lim_{p\to 1}u^{\varepsilon}_{p}$ one has an upper bound of the form $diam_{\varepsilon}(\{u=s\})\lesssim e^{s/(Q-1)}$.
The pointwise lower bound is given by the following result.
Proposition 5.1.
Let $1<p\leq 2$, $\varepsilon\in[0,1]$, and let $\Omega\subset\mathbb{H}^{n}$ be an open set with $C^{2}$-smooth boundary and bounded complement satisfying assumption assuption ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}$ such that $B_{R_{0}}(g_{0})\subset\Omega^{c}$ and $\mathbb{H}^{n}\setminus\Omega\subset B_{\bar{R}}(g_{0})$ for some $R_{0}<2<\bar{R}$. There exist $C>0$ depending only on $R_{0}$ and $Q$ and there exists $\widehat{C}>1$ depending only on $R_{0}$, $\bar{R}$ and $Q$ such that for $v=v^{\varepsilon}_{p}$ the minimizer of (1.10) in the set $\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ we have
(5.1) $v(g)\geq\widehat{C}^{-\frac{\varepsilon^{4}}{p-1}}\left(\frac{\|g_{0}^{-1}*g\|}{R_{0}}\right)^{-\gamma}\qquad\hbox{on }\quad\overline{\Omega}\,,\qquad\hbox{with}\quad\gamma=\frac{Q-p}{p-1}\,,\quad\widehat{C}=\left(4\frac{\bar{R}}{R_{0}}\right)^{C}\,.$
Proof.
First notice that the claim is true for $\varepsilon=0$ by chosing any $\widehat{C}>0$. Indeed in that case both sides in (5.1) are continuous and finite energy $p$-harmonic functions in $\Omega$ with respect to the sub-Riemannian metric $|\cdot|_{0}$, moreover they are clearly well ordered on $\partial\Omega$, so that inequality (5.1) holds in the whole $\Omega$ as it follows from the comparison principle for finite energy $p$-harmonic functions.
Assuming $\varepsilon\in(0,1]$ in the sequel, up to traslations we may also assume $g_{0}=0$. Notice that the right hand side in (5.1) is no longer $p$-harmonic in $\Omega$ but a slight tilt in the exponent still gives a subsolution according to Proposition 3.3. Then we iterate this property at different scales and with different tilts in the exponent in order to prove that (5.1) actually holds for suitably chosen $\widehat{C}>0$.
In what follows for $\varepsilon\in(0,1]$ we set $\mu(\varepsilon)=\frac{\varepsilon^{4}}{p-1}\left(\frac{8}{R_{0}^{2}}+\frac{32Q}{R_{0}^{4}}\right)$, so that in view of Lemma 3.2 and Proposition 3.3 the function $\|\cdot\|^{-(\gamma+\mu(\varepsilon))}$ has finite energy and it satisfies (3.33), i.e., it is weakly $p$-subharmonic w.r.to the metric $|\cdot|_{\varepsilon}$ outside the ball $B_{R_{0}}(0)$, whence comparison principle yields $R_{0}^{-\gamma-\mu(\varepsilon)}v\geq\|\cdot\|^{-(\gamma+\mu(\varepsilon))}$ in $\overline{\Omega}$. Now for $\bar{R}>2>R_{0}$ such that $\mathbb{H}^{n}\setminus\Omega\subset B_{\bar{R}}(0)$, if $\widetilde{\Omega}=\mathbb{H}^{n}\setminus\overline{B_{\bar{R}}(0)}$ then we have $v\in\dot{W}^{1,p}_{\varepsilon}(\widetilde{\Omega})\cap C(\overline{\widetilde{\Omega}})$. Next, using homogeneous dilations we define by induction
$v_{0}(g)=\bar{R}^{\gamma}v(\delta_{\bar{R}}(g))\,,\qquad v_{j+1}(g)=2^{\gamma}v_{j}(\delta_{2}(g))=\left(2^{\gamma}\right)^{j+1}v_{0}(\delta_{2^{j+1}}(g))\,,\quad j\in\mathbb{N}\,.$
Clearly if we set $\bar{R}_{j}=\bar{R}2^{j}$ and $\varepsilon_{j}=\varepsilon/\bar{R}_{j}$ then $\varepsilon_{j}\in(0,1]$ for all $j$, $\varepsilon_{j}\to 0$ as $j\to\infty$, moreover $v_{j}(g)=\bar{R}^{\gamma}_{j}v(\delta_{\bar{R}_{j}}(g))$ for any $j\in\mathbb{N}$ and each $v_{j}$ is $p$-harmonic in $B_{1}(0)^{c}$ w.r. to the metric $|\cdot|_{\varepsilon_{j}}$.
Note that, as above, if $w\in\dot{W}^{1,p}_{\tilde{\varepsilon}}(\overline{B_{1}(0)}^{c})\cap C(B_{1}(0)^{c})$ is $p$-harmonic in $\overline{B_{1}(0)}^{c}$ w.r. to the metric $|\cdot|_{\tilde{\varepsilon}}$ then $w\geq\left(\min_{\partial B_{1}(0)}\,w\right)\|\cdot\|^{-(\gamma+\mu(\tilde{\varepsilon}))}$ on $B_{1}(0)^{c}$, hence the same holds on $B_{1}(0)^{c}\cap\overline{B_{2}(0)}$ and in particular we have the inequality $2^{\gamma}\left(\min_{\partial B_{2}(0)}\,w\right)\geq 2^{-\mu(\tilde{\varepsilon})}\left(\min_{\partial B_{1}(0)}\,w\right)$. Applying the latter inequalities to $w=v_{j}$ and $\tilde{\varepsilon}=\varepsilon_{j}$ as defined above and setting $m_{j}=\min_{\partial B_{1}(0)}\,v_{j}$ we obtain for each $j\geq 0$
(5.2) $v_{j}\geq m_{j}\|\cdot\|^{-(\gamma+\mu(\varepsilon_{j}))}\quad\hbox{on}\,\,\,B_{1}(0)^{c}\cap\overline{B_{2}(0)}\,,\qquad m_{j+1}\geq 2^{-\mu(\varepsilon_{j})}m_{j}\geq m_{0}2^{-\sum_{k=0}^{j}\mu(\varepsilon_{k})}\,.$
Clearly $\mu(\varepsilon_{j})\leq\frac{C\varepsilon^{4}}{p-1}2^{-j-1}$ for all $j\geq 0$, where $C>0$ depends only on $R_{0}$ and $Q$, whence $\sum_{j}\mu(\varepsilon_{j})\leq\frac{C\varepsilon^{4}}{p-1}$, $m_{j}\geq m_{0}\left(2^{C}\right)^{\varepsilon^{4}/(1-p)}$ and $v_{j}\geq m_{j}\|\cdot\|^{-(\gamma+\mu(\varepsilon_{j}))}\geq m_{0}\left(4^{C}\right)^{\varepsilon^{4}/(1-p)}\|\cdot\|^{-\gamma}$ on $B_{1}(0)^{c}\cap\overline{B_{2}(0)}$.
Thus, by construction
$m_{0}=\bar{R}^{\gamma}\,\min_{\partial B_{\bar{R}}(0)}\,v\geq\bar{R}^{\gamma}\left(\frac{\bar{R}}{R_{0}}\right)^{-\gamma-\mu(\varepsilon)}\geq R_{0}^{\gamma}\left(\frac{\bar{R}}{R_{0}}\right)^{-\frac{C\varepsilon^{4}}{p-1}}$
and in turn for each $j\geq 0$
$v_{j}(g)\geq R_{0}^{\gamma}\left(\frac{\bar{R}}{R_{0}}\right)^{-\frac{C\varepsilon^{4}}{p-1}}\left(4^{C}\right)^{\varepsilon^{4}/(1-p)}\|g\|^{-\gamma}=\left(4\frac{\bar{R}}{R_{0}}\right)^{-\frac{C\varepsilon^{4}}{p-1}}\left(\frac{\|g\|}{R_{0}}\right)^{-\gamma}\quad\hbox{for each}\quad g\in B_{1}(0)^{c}\cap\overline{B_{2}(0)}\,,$
therefore the identity $v_{j}(g)=\bar{R}^{\gamma}_{j}v(\delta_{\bar{R}_{j}}(g))$ for any $j\in\mathbb{N}$ yield inequality (5.1) on $\mathbb{H}^{n}\setminus B_{\bar{R}}(0)$ by chosing $\widehat{C}=\left(4\frac{\bar{R}}{R_{0}}\right)^{C}>1$. Finally, since $R_{0}^{-\gamma-\mu(\varepsilon)}v\geq\|\cdot\|^{-(\gamma+\mu(\varepsilon))}$ in $\overline{\Omega}$ yields $R_{0}^{-\gamma}v\geq(4\bar{R}/R_{0})^{-\mu(\varepsilon)}\|\cdot\|^{-\gamma}$ in $\overline{\Omega}\cap B_{\bar{R}}(0)$, then from the choice of $C$ above and the corresponding definition of $\widehat{C}$ inequality (5.1) actually holds in the whole $\overline{\Omega}$. This closes the proof. ∎
In order to obtain a pointwise upper bound for the $p$-capacitary potentials we start with the following auxiliary result announced in Section 2 about the Riemannian perimeters $Per_{\varepsilon}(\cdot)$ and their relation with the horizontal perimeter $Per_{0}(\cdot)$. The reader should refer to [33, Section 5] for a similar statement, although for a slightly different definition of $Per_{\varepsilon}(\cdot)$ corresponding to a family of vector fields different from $\{X_{i},Y_{i},T_{\varepsilon}\}$.
Lemma 5.2.
Let $\widetilde{\Omega}\subset\mathbb{H}^{n}$ be a bounded open set with $C^{1}$-smooth boundary, $\mathcal{H}^{2n}$ be the Eucledean $2n$-dimensional Hausdorff measure and let $\hat{n}$ be the Euclidean outer normal vector field.
For each $\varepsilon>0$ we have the equality
[…]
As a consequence $\varepsilon\to\varepsilon Per_{\varepsilon}(\widetilde{\Omega})$ is non decreasing, $\varepsilon Per_{\varepsilon}(\widetilde{\Omega})\leq Per_{1}(\widetilde{\Omega})$ for any $\varepsilon\in(0,1]$, and $\varepsilon Per_{\varepsilon}(\widetilde{\Omega})\to Per_{0}(\widetilde{\Omega})$ as $\varepsilon\to 0$.
Proof.
It is clearly enough to establish the validity of (5.3), whence the the other properties follow easily taking into account the representation formula for $P_{0}(\widetilde{\Omega})$ already proved in [12, Section 3] (see also [33] and [21, Section 7]).
The proof of (5.3) is very similar to the one in [12, Section 3] for $Per_{0}(\widetilde{\Omega})$. Indeed, recall that for any smooth vector field $Z_{\varphi}=\sum_{i=1}^{n}\left[\varphi_{i}X_{i}+\varphi_{n+i}Y_{i}\right]+\varphi_{2n+1}T_{\varepsilon}$ associated to a smooth map $\varphi:\mathbb{H}^{n}\to\mathbb{R}^{2n+1}$ its Riemannian divergence is given by ${\rm div}_{\varepsilon}Z_{\varphi}=\sum_{i=1}^{n}\left[(X_{i}\varphi_{i}+Y_{i}\varphi_{n+i}\right]+T_{\varepsilon}\varphi_{2n+1}\,,$ and moreover that we have the identities $X_{i}\varphi_{i}=\partial_{x_{i}}\varphi_{i}-\partial_{t}(\frac{y_{i}}{2}\varphi_{i})$, $Y_{i}\varphi_{i}=\partial_{y_{i}}\varphi_{n+i}+\partial_{t}(\frac{x_{i}}{2}\varphi_{n+i})$ and $T_{\varepsilon}\varphi_{2n+1}=\partial_{t}(\varepsilon\varphi_{2n+1})$.
Then using the identity $\varepsilon\,dvol_{\varepsilon}(\cdot)=\mathcal{L}^{2n+1}$ and the Euclidean divergence theorem we obtain
[…]
whence (5.3) follows from the definition of $Per_{\varepsilon}(\cdot)$ taking the supremum over maps $\varphi\in C_{0}^{\infty}(\mathbb{H}^{n};\mathbb{R}^{2n+1})$ satisfying $|\varphi|\leq 1$ everywhere in $\mathbb{H}^{n}$.
Next, relying on the previous lemma we prove a global weak-$L^{\sigma}$ integrability of $p$-capacitary potentials for $\sigma=(p-1)\frac{Q}{Q-p}$. Recall that for any $\sigma>0$ the weak-$L^{\sigma}$ space $L^{\sigma,\infty}(\Omega)$ is the set of measurable functions $w:\Omega\to\mathbb{R}$ (up to a.e. equivalence) having finite (quasi)norm (see, e.g., [25]),
(5.4) $\|w\|_{L^{\sigma,\infty}(\Omega)}=\sup_{\rho>0}\rho^{\sigma}|\{|w|>\rho\}|\,.$
The following result gives control of $\|v^{\varepsilon}_{p}\|_{L^{\sigma,\infty}(\Omega)}$ uniformly in $\varepsilon\in[0,1]$ and quantified w.r.to $p\in(1,2]$.
Lemma 5.3.
Let $1<p\leq 2$, $\varepsilon\in(0,1]$ and $\sigma=\frac{Q(p-1)}{Q-p}$. Let $\Omega\subset\mathbb{H}^{n}$ be an open set with $C^{2}$-smooth boundary and bounded complement, $\widetilde{\Omega}=\mathbb{H}^{n}\setminus\overline{\Omega}$ and let $v=v^{\varepsilon}_{p}$ be the minimizer of (1.10) in the set $\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$.
Then there exists $C>0$ depending only on $Q$ such that
(5.5) $\|v\|_{L^{\sigma,\infty}(\Omega)}\leq C\|\nabla_{\varepsilon}v\|^{\sigma}_{L^{\infty}(\partial\Omega)}Per_{1}(\widetilde{\Omega})^{Q/(Q-p)}\,.$
As a consequence, if $\Omega$ satisfies also assumption $(\bf{HP_{\Omega}})$ with parameter $R_{0}>0$ then there exist $C^{\prime}>0$ depending only on $Q$ and $R_{0}$ such that $\|v\|_{L^{\sigma,\infty}(\Omega)}\leq C^{\prime}$.
Proof.
For fixed $0<\tau<\rho\leq\frac{1}{2}$ we consider auxiliary piecewise affine functions $g_{\tau}:[0,1]\to[0,\infty)$ given by $g_{\tau}(r)=\min\{r,\rho,\frac{\rho}{\tau}(1-r)\}$, so that in particular $g^{\prime}_{\tau}(r)=\chi_{(0,\rho)}(r)-\frac{\rho}{\tau}\chi_{(1-\tau,1)}(r)$ a.e. in $(0,1)$ and $g_{\tau}(1)=0$.
Since $0\leq v\leq 1$ and it is continuous in $\mathbb{H}^{n}$, by the chain rule we have $\varphi=g_{\tau}(v)\in\dot{W}^{1,p}_{\varepsilon,0}(\Omega)$, $\varphi\equiv 0$ in $\widetilde{\Omega}$ and $\nabla_{\varepsilon}\varphi=(\chi_{\{v<\rho\}}-\frac{\rho}{\tau}\chi_{\{1-\tau<v\}})\nabla_{\varepsilon}v$ a.e. in $\Omega$. Using $\varphi$ as an admissible test function for $v$ we obtain
$\int_{\{v<\rho\}}|\nabla_{\varepsilon}v|_{\varepsilon}^{p}=\frac{\rho}{\tau}\int_{\{1-\tau<v\}}|\nabla_{\varepsilon}v|_{\varepsilon}^{p}\leq\||\nabla_{\varepsilon}v|_{\varepsilon}\|_{L^{\infty}(\{1-\tau\leq v\})}^{p-1}\frac{\rho}{\tau}\int_{\{1-\tau<v\}}|\nabla_{\varepsilon}v|_{\varepsilon}\,.$
In order to estimate the right hand side as $\tau\to 0$, for fixed $\varepsilon$ we rely on the $C^{1}$-boundary regularity of $v$ to obtain control of $|\nabla_{\varepsilon}v|_{\varepsilon}$. Note that $v<1$ in $\Omega$ by the strong maximum principle, moreover $\nabla_{\varepsilon}v\neq 0$ on $\partial\Omega$ because of the Hopf lemma applied at any (maximum) point of $\partial\Omega$. As a consequence $\cap_{\tau>0}\{v\geq 1-\tau\}=\widetilde{\Omega}$ and $\||\nabla_{\varepsilon}v|_{\varepsilon}\|_{L^{\infty}(\{1-\tau\leq v\})}=(1+o(1))\||\nabla_{\varepsilon}v|_{\varepsilon}\|_{L^{\infty}(\partial\Omega)}$ as $\tau\to 0$.
On the other hand, since $\widetilde{\Omega}$ has smooth boundary $\partial\widetilde{\Omega}=\partial\Omega$, the same holds for $\widetilde{\Omega}_{r,\varepsilon}=\{v>1-r\}$ for any $0<r<\tau$ and for $\tau$ small enough because $\nabla_{\varepsilon}v\neq 0$ on $\partial\Omega$. As a consequence $Per_{\varepsilon}(\widetilde{\Omega}_{r,\varepsilon})=(1+o(1))Per_{\varepsilon}(\widetilde{\Omega})$ as $r\to 0$. Since $\mathcal{L}^{n+1}=\varepsilon\,dvol_{\varepsilon}(\cdot)$, then the Riemannian coarea formula together with the previous expansions yield
[…]
hence, from Lemma 5.2 as $\tau\to 0$, we conclude that
(5.6) $\int|\nabla_{\varepsilon}\min\{v,\rho\}|_{\varepsilon}^{p}=\int_{\{v<\rho\}}|\nabla_{\varepsilon}v|_{\varepsilon}^{p}\leq\rho\||\nabla_{\varepsilon}v|_{\varepsilon}\|_{L^{\infty}(\partial\Omega)}^{p-1}Per_{1}(\widetilde{\Omega})\,,$
for any $\varepsilon\in(0,1]$, $1<p\leq 2$ and $\rho\in(0,\frac{1}{2}]$.
Next, applying Chebyshev’s inequality and the Sobolev inequality (2.4) for $\rho\in(0,\frac{1}{2}]$ from (5.6) we infer that
$\rho^{p}|\{v>\rho\}|^{p/p^{*}}\leq\left(\int(\min\{v,\rho\})^{p^{*}}\right)^{p/p^{*}}\leq(Cp^{*})^{p}\rho\||\nabla_{\varepsilon}v|_{\varepsilon}\|_{L^{\infty}(\partial\Omega)}^{p-1}Per_{1}(\widetilde{\Omega})\,,$
for $C>0$ depending only on $Q$.
Finally, for $1<p\leq 2$ we have $p^{*}\leq 2Q/(Q-2)$ and $\sigma\leq Q/(Q-2)$, therefore raising both sides of the previous inequality to $p^{*}/p$ we get
$\|v\|_{L^{\sigma,\infty}(\Omega)}=\sup_{0<\rho<1}\rho^{\sigma}|\{v>\rho\}\cap\Omega|\leq\sup_{0<\rho\leq 1/2}2^{\sigma}\rho^{\sigma}|\{v>\rho\}\cap\Omega|\leq C\||\nabla_{\varepsilon}v|_{\varepsilon}\|_{L^{\infty}(\partial\Omega)}^{\sigma}Per_{1}(\widetilde{\Omega})^{Q/(Q-p)}\,,$
for $C>0$ depending only on $Q$, which completes the proof of (5.5).
Finally, in order to finish the proof we rely on Proposition 6.2, so that inequality (6.4) for some $\overline{C}>0$ depending only on $R_{0}$ and $Q$ together with (5.5) yield
$\|v\|_{L^{\sigma,\infty}(\Omega)}\leq C\left(\frac{\overline{C}^{p-1}}{(p-1)^{p-1}}Per_{1}(\widetilde{\Omega})\right)^{\frac{Q}{Q-p}}\,$
and the conclusion follows as the right hand side is uniformly bounded for $p\in(1,2]$. ∎
The following result provides uniform Harnack inequalities on annular regions.
Proposition 5.4.
Let $1<p\leq 2$, $p_{*}=\frac{Q(p-1)}{Q-p/2}$ and for $r>\varepsilon>0$ let us set $\Omega_{r}=B^{\varepsilon}_{5r}(\bar{g})\setminus\overline{B^{\varepsilon}_{r}(\bar{g})}$ and $\widetilde{\Omega}_{r}=B^{\varepsilon}_{4r}(\bar{g})\setminus\overline{B^{\varepsilon}_{2r}(\bar{g})}\subset\Omega_{r}$ There exists $C_{*}>1$ depending only on $Q$ such that if $v=v^{\varepsilon}_{p}\in W^{1,p}_{\varepsilon}(\Omega_{r})$ is nonegative and $p$-harmonic in $\Omega_{r}$ w.r. to the metric $|\cdot|^{2}_{\varepsilon}$ then for any $q\geq p_{*}$ we have
(5.7) $\|v\|_{L^{\infty}(\widetilde{\Omega}_{r})}\leq C_{*}^{\frac{1}{p-1}}\left(\frac{1}{|\Omega_{r}|}\int_{\Omega_{r}}v^{q}\right)^{1/q}\,.$
Proof.
First notice that $v\in C^{1}(\Omega_{r})$ by the regularity theory for $p$-harmonic functions on Riemannian manifolds and, as in Proposition 4.2, $v>0$ by the strong maximum principle unless $v\equiv 0$ so that (5.7) is trivially satisfied. In addition, in view of Hölder inequality it is clearly enough to prove (5.7) for $q=p_{*}$, moreover, since both the $p$-harmonic equation and the inequality (5.7) are translation invariant, from now on we may assume $\bar{g}=0$ and drop the dependence on $\bar{g}$ in the sequel.
As in the proof of Proposition 4.2 we first derive a Caccioppoli-type inequality for suitable powers of $v$. Thus, for $\beta>-1/p$ and for nonnegative $d_{\varepsilon}$-Lipschitz function $\zeta$ compactly supported in $\Omega_{r}$ to be specified later, we test the equation for $v$ with $\zeta^{p}v^{1+\beta p}$, hence integration by parts, Hölder and Young inequalities combined as in the proof of (4.18) yield
(5.8) $\int_{\Omega_{r}}|\nabla_{\varepsilon}\left(\zeta v^{1+\beta}\right)|_{\varepsilon}^{p}\leq\left(\frac{2(1+\beta)p}{1+p\beta}\right)^{p}\int_{\Omega_{r}}\left(v^{(1+\beta)}|\nabla_{\varepsilon}\zeta|_{\varepsilon}\right)^{p}\,,$
where we have used that $\frac{p(1+\beta)}{1+p\beta}>1$ for $p>1$ and $\beta>-1/p$.
Next, in analogy with the previous proof we specialize inequality (5.8) by chosing suitable test functions defined in terms of the metric $d_{\varepsilon}$ in order to obtain (5.7) from (5.8) through Moser iteration. For $\frac{r}{2}<\frac{1}{2}\rho_{2}\leq\rho_{1}<\rho_{2}\leq 2r$ we consider $h:[0,\infty)\to[0,1]$ a Lipschitz and piecewhise linear function, with $\{h\equiv 1\}=[3r-\rho_{1},3r+\rho_{1}]$, $\{h\equiv 0\}=[0,3r-\frac{\rho_{1}+\rho_{2}}{2}]\cup[3r+\frac{\rho_{1}+\rho_{2}}{2},\infty)$, and $h$ is linear otherwise. Hence, in view of the inequality $|\nabla_{\varepsilon}d_{\varepsilon}(\cdot,\bar{g})|_{\varepsilon}\leq 1$ on $\mathbb{H}^{n}$ the function $\zeta(g)=h(d_{\varepsilon}(g,0))$ by construction is $d_{\varepsilon}$-Lipschitz on $\mathbb{H}^{n}$ with $\zeta\equiv 1$ on $\overline{B^{\varepsilon}_{3r+\rho_{1}}}\setminus B^{\varepsilon}_{3r-\rho_{1}}$, $spt\,\zeta\subset\subset B^{\varepsilon}_{3r+\rho_{2}}\setminus\overline{B^{\varepsilon}_{3r-\rho_{2}}}$, and $|\nabla_{\varepsilon}\zeta|_{\varepsilon}\leq\frac{2}{\rho_{2}-\rho_{1}}\,$ a.e. in $\mathbb{H}^{n}$.
Now note that for $\rho_{1}$ and $\rho_{2}$ as above we have the following chain of inclusions, namely $B^{\varepsilon}_{\frac{7}{2}r}\setminus\overline{B^{\varepsilon}_{\frac{5}{2}r}}\subset B^{\varepsilon}_{3r+\rho_{1}}\setminus\overline{B^{\varepsilon}_{3r-\rho_{1}}}\subset B^{\varepsilon}_{3r+\rho_{2}}\setminus\overline{B^{\varepsilon}_{3r-\rho_{2}}}\subset B^{\varepsilon}_{3r+\rho_{2}}\subseteq B^{\varepsilon}_{5r}$, hence all the volumes are uniformly equivalent for $r>\varepsilon>0$ because in such a range of parameters there exists absolute constant $C_{0}>1$ depending only on $Q$ such that $C_{0}^{-1}|B^{\varepsilon}_{5r}|\leq r^{Q}\leq C_{0}|B^{\varepsilon}_{\frac{7}{2}r}\setminus\overline{B^{\varepsilon}_{\frac{5}{2}r}}|$ as follows from (2.2) and Lemma 2.6.
Going back to (5.8) we apply the uniform Sobolev inequality (2.4) for functions supported on $B^{\varepsilon}_{3r+\rho_{2}}$ with uniform constant $Cp^{*}\leq\overline{C}$ for a constant $\overline{C}>0$ independent of $p\in(1,2]$. Thus, taking the aforementioned equivalence of volumes into account, the previous choice of $\zeta$ and Lemma 2.2-(5) for $\varepsilon<3r+\rho_{2}<5\rho_{2}$ yield
(5.9) $\left(\frac{1}{|B^{\varepsilon}_{3r+\rho_{1}}\setminus\overline{B^{\varepsilon}_{3r-\rho_{1}}}|}\int_{B^{\varepsilon}_{3r+\rho_{1}}\setminus\overline{B^{\varepsilon}_{3r-\rho_{1}}}}v^{(1+\beta)p^{*}}\right)^{\frac{1}{(1+\beta)p^{*}}}\leq\left(\frac{C_{0}^{2}}{|B^{\varepsilon}_{3r+\rho_{2}}|}\int_{B^{\varepsilon}_{3r+\rho_{2}}}\left(\zeta v^{(1+\beta)}\right)^{p^{*}}\right)^{\frac{1}{(1+\beta)p^{*}}}$
$\leq{C_{0}}^{\frac{2}{(1+\beta)p^{*}}}\,\overline{C}^{1/(1+\beta)}|B^{\varepsilon}_{3r+\rho_{2}}|^{\frac{1}{Q(1+\beta)}}\left(\frac{1}{|B^{\varepsilon}_{3r+\rho_{2}}|}\int_{B^{\varepsilon}_{3r+\rho_{2}}}|\nabla_{\varepsilon}\left(\zeta v^{1+\beta}\right)|_{\varepsilon}^{p}\right)^{\frac{1}{(1+\beta)p}}$
$\leq\left[C\cdot\frac{\rho_{2}}{\rho_{2}-\rho_{1}}\cdot\frac{2(1+\beta)p}{1+p\beta}\right]^{\frac{1}{1+\beta}}\left(\frac{1}{|B^{\varepsilon}_{3r+\rho_{2}}\setminus\overline{B^{\varepsilon}_{3r-\rho_{2}}}|}\int_{B^{\varepsilon}_{\rho_{2}}(\bar{g})}v^{(1+\beta)p}\right)^{\frac{1}{(1+\beta)p}}\,,$
where we have also used that $C_{0}^{1/p^{*}}<C_{0}$ as $C_{0}>1$ and $p^{*}=\frac{Qp}{Q-p}>p>1$, and (5.8) to obtain the last inequality. Note that for $p\in(1,2]$ the constant $C>0$ in the square brackets depends only on $Q$.
Next, using (5.9) we perform Moser iteration with exponents $\{\beta_{i}\}_{i\geq 0}$ where $\beta_{0}>-1/p$ is chosen such that $(1+\beta_{0})p=p_{*}$ and $\beta_{i+1}=p/(Q-p)+\beta_{i}Q/(Q-p)$, so that $\{\beta_{i}\}\subset(-1/p,\infty)$ and it is increasing to infinity. Indeed, note that if $s=p(\beta+1)$ then $\beta>-1/p$ iff $s>p-1$ so that for $s_{i}=p(\beta_{i}+1)$ we have $s_{i+1}=s_{i}Q/(Q-p)=p_{*}\frac{Q^{i+1}}{(Q-p)^{i+1}}$ is increasing and $s_{i}>(p-1)\frac{Q^{i+1}}{(Q-1/2)^{i+1}}\uparrow+\infty$ as $i\to\infty$. Moreover, since $\beta\to\frac{(1+\beta)p}{1+\beta p}$ is decreasing for $\beta>-1/p$ one has $\frac{(1+\beta_{i})p}{1+\beta_{i}p}\leq\frac{(1+\beta_{0})p}{1+\beta_{0}p}\leq 2Q$ for any $i\geq 0$ and $p>1$.
For fixed $r>\varepsilon>0$ and $\frac{r}{2}<\frac{1}{2}\rho_{2}\leq\rho_{1}<\rho_{2}\leq 2r$ we repeatedly use (5.9) with the choice $(\rho_{1},\rho_{2},\beta)=(\tilde{\rho}_{i+1},\tilde{\rho}_{i},\beta_{i})$, where $\tilde{\rho}_{i}=r(1+2^{-i})$ for $i\geq 0$, so that $\frac{\rho_{2}}{\rho_{2}-\rho_{1}}=\frac{\tilde{\rho}_{i}}{\tilde{\rho}_{i}-\tilde{\rho}_{i+1}}\leq 2^{i+2}$ for any $i\geq 0$.
Thus, inequality(4.8) rewrites for all $i\geq 0$ as
(5.10) $\left(\frac{1}{|B^{\varepsilon}_{3r+\tilde{\rho}_{i+1}}\setminus\overline{B^{\varepsilon}_{3r-\tilde{\rho}_{i+1}}}|}\int_{B^{\varepsilon}_{3r+\tilde{\rho}_{i+1}}\setminus\overline{B^{\varepsilon}_{3r-\tilde{\rho}_{i+1}}}}v^{s_{i+1}}\right)^{\frac{1}{s_{i+1}}}$
$\quad\quad\leq\left(C2^{i+2}\right)^{\frac{1}{1+\beta_{i}}}\left(\frac{1}{|B^{\varepsilon}_{3r+\tilde{\rho}_{i}}\setminus\overline{B^{\varepsilon}_{3r-\tilde{\rho}_{i}}}|}\int_{B^{\varepsilon}_{3r+\tilde{\rho}_{i}}\setminus\overline{B^{\varepsilon}_{3r-\tilde{\rho}_{i}}}}v^{s_{i}}\right)^{\frac{1}{s_{i}}},$
for a constant $C>1$ depending only on $Q$.
Since $s_{0}=p_{*}$ and $\tilde{\rho}_{0}=2r$, iterating (5.10) as in the proof of the previous lemma we easily obtain (5.7), with a constant factor $C_{*}>1$ depending only on $Q$ ensuring the bound
$\Pi_{i=0}^{\infty}\left(C2^{i+2}\right)^{\frac{p}{s_{i}}}\leq\Pi_{i=0}^{\infty}\left(C2^{i+2}\right)^{\frac{2(Q-1/2)^{i+1}}{(p-1)Q^{i+1}}}=\left(\Pi_{i=0}^{\infty}\left(C^{2}4^{i+2}\right)^{\frac{(Q-1/2)^{i+1}}{Q^{i+1}}}\right)^{\frac{1}{p-1}}=:C_{*}^{\frac{1}{p-1}}<\infty\,,$
where the convergence of the infinite product is a simple consequence of $\sum_{i=0}^{\infty}\frac{i(Q-1/2)^{i}}{Q^{i}}<\infty$. This closes the proof. ∎
The pointwise upper bound for the $p$-capacitary potentials is given by the following result.
Proposition 5.5.
Let $1<p\leq 2$, $\varepsilon\in[0,1]$, and let $\Omega\subset\mathbb{H}^{n}$ be an open set with $C^{2}$-smooth boundary and bounded complement satisfying assumption assuption ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}$ such that $B_{R_{0}}(g_{0})\subset\Omega^{c}$ and $\mathbb{H}^{n}\setminus\Omega\subset B_{\bar{R}}(g_{0})$ for some $R_{0}<2<\bar{R}$. There exist $C_{0}>1$ depending only on $R_{0}$, $\bar{R}$, $\partial\Omega$ and $Q$ such that for $v=v^{\varepsilon}_{p}$ the minimizer of (1.10) in the set $\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ we have
(5.11) $v(g)\leq{C_{0}}^{\frac{1}{p-1}}\left(\frac{\|g_{0}^{-1}*g\|}{\bar{R}}\right)^{-\gamma}\qquad\hbox{on }\quad\overline{\Omega}\,,\qquad\hbox{with}\quad\gamma=\frac{Q-p}{p-1}\,.$
Proof.
First we notice that inequality (5.11) trivially holds in $\overline{\Omega}\cap B_{\bar{R}}(g_{0})$ for any $C_{0}>1$ because $0<v\leq 1$ in $\overline{\Omega}$. Then the claim is still true for $\varepsilon=0$ in the whole $\overline{\Omega}$ by chosing any $C_{0}>1$. Indeed in that case both sides in (5.1) are continuous and finite energy $p$-harmonic functions in $\mathbb{H}^{n}\setminus B_{\bar{R}}(g_{0})$ with respect to the sub-Riemannian metric $|\cdot|_{0}$, moreover they are clearly well ordered on $B_{\bar{R}}(g_{0})$, so that inequality (5.1) holds in the whole $\mathbb{H}^{n}\setminus B_{\bar{R}}(g_{0})$ and in turn on $\Omega$, as it follows from the comparison principle for finite energy $p$-harmonic functions.
When dealing with (5.11) for $\varepsilon\in(0,1]$ we rely instead on Proposition 5.4 and Lemma 5.3. More precisely, for $1<p\leq 2$ we derive an upper bound for $v=v^{\varepsilon}_{p}$ from the intrinsic Harnack inequality (5.7) with exponent $q=p_{*}=\frac{Q(p-1)}{Q-p/2}$ applied to the corresponding potential $v$ on dyadic annuli and we combine it with the global weak-$L^{\sigma}$ integrability, with $\sigma=\frac{Q(p-1)}{Q-p}>q$, given in (5.5) and further localized on the annuli.
Indeed, note that by assumption $\Omega^{c}\subset B_{\bar{R}}(g_{0})$ and $B_{\bar{R}}(g_{0})\subset B^{\varepsilon}_{C\bar{R}}(g_{0})$ for $C>1$ depending only on $Q$ given by Lemma 2.2, so that $\mathbb{H}^{n}\setminus B^{\varepsilon}_{C\bar{R}}(g_{0})\subset\Omega$. For $g\not\in B^{\varepsilon}_{C2\bar{R}}(g_{0})$ we select $i\in\mathbb{N}$ such that $C2^{i+1}\bar{R}\leq d_{\varepsilon}(g_{0},g)<C2^{i+2}\bar{R}$ and we apply Proposition 5.4 with $r=r_{i}=C2^{i}\bar{R}$, because $v^{\varepsilon}_{p}$ is $p$-harmonic in $\Omega_{r}=\Omega_{r_{i}}:=\{C2^{i}\bar{R}<d_{\varepsilon}(g_{0},\cdot)<5C2^{i}\bar{R}\}\subset\Omega$ and $g\in\overline{\widetilde{\Omega}_{r}}=\{C2^{i+1}\bar{R}\leq d_{\varepsilon}(g_{0},g)\leq C2^{i+2}\bar{R}\}$ in view of the choice of $i\in\mathbb{N}$. Thus inequality (5.7) yields
(5.12) $v(g)\leq C_{*}^{\frac{1}{p-1}}\left(\frac{1}{|\Omega_{r_{i}}|}\int_{\Omega_{r_{i}}}v^{p_{*}}\right)^{1/p_{*}}$
for a constant $C_{*}>1$ depending only on $Q$.
On the other hand, notice that still Lemma 2.2 gives $2C\varepsilon<C2^{i}\bar{R}\leq d_{\varepsilon}(g_{0},g)\leq C\|g_{0}^{-1}*g\|_{\varepsilon}$, whence $\|g_{0}^{-1}*g\|\leq 2\|g_{0}^{-1}*g\|_{\varepsilon}\leq 2Cd_{\varepsilon}(g_{0},g)\leq 8C^{2}2^{i}\bar{R}=R_{0}\tilde{r}_{i}$ and in turn $v(g)\geq\widehat{C}^{1/(1-p)}\tilde{r}_{i}^{-\gamma}$ on $\Omega_{r_{i}}$, where $\tilde{r}_{i}=8C^{2}2^{i}\bar{R}/R_{0}=8Cr_{i}/R_{0}$, because of Proposition 5.1.
Thus, from Lemma 5.3 and $\sigma/p_{*}=1+\frac{p/2}{Q-p}>1$ we infer that
[…]
$=|\Omega_{r_{i}}|\left(\widehat{C}^{1/(1-p)}\tilde{r}_{i}^{-\gamma}\right)^{p_{*}}+\frac{1}{\sigma/p_{*}-1}\left(\widehat{C}^{1/(1-p)}\tilde{r}_{i}^{-\gamma}\right)^{p_{*}-\sigma}\|v\|_{L^{\sigma,\infty}(\Omega)}$
$=|\Omega_{r_{i}}|\left(\widehat{C}^{1/(1-p)}\tilde{r}_{i}^{-\gamma}\right)^{p_{*}}\left(1+\frac{Q-p}{p/2}\widehat{C}^{Q/(Q-p)}\frac{\tilde{r}_{i}^{Q}}{|\Omega_{r_{i}}|}\|v\|_{L^{\sigma,\infty}(\Omega)}\right)\,.$
Now, observe that $|\Omega_{r_{i}}|\geq(c_{*}r_{i})^{Q}$ for a constant $c_{*}\in(0,1)$ depending only on $Q$ because of (2.2) and Lemma 2.2-(2), hence the choice of $r_{i}=C2^{i}\bar{R}=\tilde{r}_{i}R_{0}/(8C)$ combined with the previous inequalities for $1<p\leq 2$ give
$\frac{1}{|\Omega_{r_{i}}|}\int_{\Omega_{r_{i}}}v^{p_{*}}\leq\left(\widehat{C}^{1/(1-p)}\tilde{r}_{i}^{-\gamma}\right)^{p_{*}}\left(1+2Q\widehat{C}^{Q/(Q-2)}\left(\frac{8C}{c_{*}R_{0}}\right)^{Q}\|v\|_{L^{\sigma,\infty}(\Omega)}\right)\,,$
so that from the last inequality, (5.12) and $\tilde{r}_{i}\geq\frac{2C}{R_{0}}d_{\varepsilon}(g_{0},g)\geq\frac{1}{2C\bar{R}}d_{\varepsilon}(g_{0},g)$ we infer that
(5.13) $v(g)\leq\left(\frac{C_{*}}{\widehat{C}}\right)^{\frac{1}{p-1}}\left(1+2Q\widehat{C}^{Q/(Q-2)}\left(\frac{8C}{c_{*}R_{0}}\right)^{Q}\|v\|_{L^{\sigma,\infty}(\Omega)}\right)^{\frac{1}{p-1}}\left(\frac{d_{\varepsilon}(g_{0},g)}{2C\bar{R}}\right)^{-\gamma}\,,$
whenever $g\not\in B^{\varepsilon}_{C2\bar{R}}(g_{0})$ and in turn for any $g\in\overline{\Omega}$, because for $g\in B^{\varepsilon}_{C2\bar{R}}(g_{0})$ we have $v(g)\leq 1$ but each factor on the right hand side is greater than one.
As already recalled above, by Lemma 2.2-(3) when $g\not\in B^{\varepsilon}_{C2\bar{R}}(g_{0})$ we have $\|g_{0}^{-1}*g\|_{\varepsilon}\geq 2\bar{R}\geq 2\varepsilon\bar{R}>2\varepsilon>0$, so that $d_{\varepsilon}(g,g_{0})\geq C^{-1}\|g_{0}^{-1}*g\|_{\varepsilon}\geq(2C)^{-1}\|g_{0}^{-1}*g\|$. On the other hand, observing that the same lemma yields $B_{R_{0}}(g_{0})\subset B_{\bar{R}}(g_{0})\subset\mathcal{B}^{\varepsilon}_{\bar{R}}(g_{0})\subset B_{C2\bar{R}}^{\varepsilon}(g_{0})\subset B^{1}_{C2\bar{R}}(g_{0})$, if we set $M(\bar{R})=\max\{\|g_{0}^{-1}*g\|\,,\,\,g\in\overline{B^{1}_{C2\bar{R}}(g_{0})}\}$ and $m(\bar{R})=\min\{d_{1}(g_{0},g)\,,\,\,g\in\overline{B^{1}_{C2\bar{R}}(g_{0})}\cap\overline{\Omega}\}>0$, then for any $\varepsilon\in(0,1]$ we get
(5.14) $\|g_{0}^{-1}*g\|\leq\left(2C+\frac{M(\bar{R})}{m(\bar{R})}\right)d_{\varepsilon}(g_{0},g)\qquad\hbox{for any}\quad g\in\overline{\Omega}\,.$
Combining (5.13) with (5.14) and observing that $12Q\widehat{C}^{Q/(Q-2)}\left(\frac{8C}{c_{*}R_{0}}\right)^{Q}>$ 1 we eventually obtain
(5.15) $v(g)\leq\overline{C}^{\frac{1}{p-1}}\left(1+\|v\|_{L^{\sigma,\infty}(\Omega)}\right)^{\frac{1}{p-1}}\left(\frac{\|g_{0}^{-1}*g\|}{\bar{R}}\right)^{-\gamma}\,,$
for a constant $\overline{C}>0$ depending only on $Q$, $R_{0}$ and $\bar{R}$. Finally, the uniform bound for $\|v\|_{L^{\sigma,\infty}(\Omega)}$ contained in Lemma 5.3 combined with (5.15) yield the desired conclusion.
6. Uniform gradient estimates for $p$-capacitary potentials.
In this section we lay the technical ground for the proof of Theorem 1.2 and in turn Corollary 1.3, i.e., to obtain the poinwise gradient estimate (1.12) for the $p-$capacitary potentials $v_{p}^{\varepsilon}$ associated to $\overline{\Omega}^{c}$ and in turn the corresponding bound (1.13) for solutions $u_{p}^{\varepsilon}$ of (1.6a). For the sake of clarity, here only the preliminary results are discussed concerning the boundary gradient estimate and the way to propagate in the interior (see Proposition 6.2, Lemma 6.4 and Proposition 6.3), while the proofs of the main results above are postponed to the next section.
The strategy for those proofs is essentially the one outlined in the Introduction. However, when trying to apply the maximum principle in the exterior domain $\Omega$ an extra ingredient which is needed is the condition $|\nabla_{\varepsilon}v^{\varepsilon}_{p}|_{\varepsilon}=o(v_{p}^{\varepsilon})$ as $\|g\|\to\infty$ and this is the main content of Proposition 6.1 below. Indeed, both in the setting of the Euclidean space treated in [34] and in the case of sub-Riemannian Heisenberg group $\mathbb{H}^{1}$ as attempted in [15], such property for the $p$-capacitary potentials follows from a scale-invariant Harnack inequality for $p-$harmonic functions and $L^{p}-L^{\infty}$ estimates for their gradient obtained in [37]. On the other hand in the Riemannian Heisenberg group no such scale invariance is available, therefore in the present case we rely on the two uniform properties for $p$-harmonic functions established in the previous section.
The next result gives the key decay property at infinity for $p$-harmonic functions in exterior domains.
Proposition 6.1.
Let $1<p<Q$, $\varepsilon\in(0,1]$, and let $\Omega\subset\mathbb{H}^{n}$ be an open set with bounded complement and smooth boundary $\partial\Omega$. Let $v=v^{\varepsilon}_{p}$ be the minimizer of (1.10) in the set $\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$. Then $v\in C^{1}(\overline{\Omega})$ and there exists a constant $C>0$ depending only on $p$ and $Q$ such that for every $\varepsilon\in(0,1]$ the gradient bound (4.1) holds. Moreover, $|\nabla_{\varepsilon}v_{p}^{\varepsilon}|_{\varepsilon}=o(v_{p}^{\varepsilon})$ as $\|g\|\to\infty$.
Proof.
As already recalled, the $C^{1}$-regularity property up to the boundary is a consequence of the interior and boundary regularity theory for $p$-harmonic functions from [17] and [31]. Then, the global $L^{\infty}$-bound for $\nabla_{\varepsilon}v$ follows directly from Proposition 4.1 applied to arbitrary balls $B_{1}^{\varepsilon}(\bar{g})\subseteq\Omega$.
In order to justify the asymptotic decay at infinity we adapt the original proof from [34], relying on Proposition 4.1, Proposition 4.2 and a simple scaling argument. Thus, setting fo brevity $v=v^{\varepsilon}_{p}$ we have $v\in C^{1}(\Omega)$ by elliptic regularity and for any sequence $\{g_{n}\}\subset\Omega$ such that $\|g_{n}\|\to\infty$ we are going to show that $|\nabla_{\varepsilon}v(g_{n})|_{\varepsilon}=o(v(g_{n}))$ as $n\to\infty$. To do this, we first chose a sequence $\{r_{n}\}\subset(1,\infty)$ such that $r_{n}\to+\infty$ and $r_{n}=o(\|g_{n}\|)$ as $n\to\infty$ and then we consider rescaled functions $w_{n}(g)=v(L_{g_{n}}\circ\delta_{r_{n}}(g))$. Setting $\tilde{\varepsilon}_{n}=\varepsilon r_{n}^{-1}$, by construction for $n$ large enough we have $w_{n}\in C^{1}(B^{\tilde{\varepsilon}_{n}}_{36})$, where we have also the identity
(6.1) $|\nabla_{\tilde{\varepsilon}_{n}}w_{n}|^{2}_{\tilde{\varepsilon}_{n}}=r^{2}_{n}|\nabla_{\varepsilon}v|^{2}_{\varepsilon}\circ(L_{g_{n}}\circ\delta_{r_{n}})\,,$
Note that in particular the function $w_{n}$ is weakly $p$-harmonic in $B^{\tilde{\varepsilon}_{n}}_{36}$ w.r. to the rescaled metric $|\cdot|^{2}_{\tilde{\varepsilon}_{n}}$ because $v$ is weakly $p$-harmonic in $B^{\varepsilon}_{36r_{n}}(g_{n})$ w.r.to the metric $|\cdot|^{2}_{\varepsilon}$ for $n$ large enough.
Now, applying Proposition 4.2 to $\{w^{\tilde{\varepsilon}_{n}}\}\subseteq C^{1}(B_{2}^{\tilde{\varepsilon}_{n}})$ it follows from the gradient bound (4.1) that
(6.2) $|\nabla_{\tilde{\varepsilon}_{n}}w_{n}(0)|^{p}_{\tilde{\varepsilon}_{n}}\leq C\int_{B_{1}^{\tilde{\varepsilon}_{n}}(0)}|\nabla_{\tilde{\varepsilon}_{n}}w_{n}|_{\tilde{\varepsilon}_{n}}^{p}\,,$
for a constant $C>0$ independent of $n$.
Next, we let $\psi_{\tilde{\varepsilon}}(g)=\min\{1,\max\{-1+d_{\tilde{\varepsilon}}(g,0),0\}\}$, so that $0\leq\psi_{\tilde{\varepsilon}}\leq 1$ on $\mathbb{H}^{n}$, $\psi_{\tilde{\varepsilon}}\equiv 1$ on $B_{1}^{\tilde{\varepsilon}}$, $spt\,\psi_{\tilde{\varepsilon}}\subset\overline{B_{2}^{\tilde{\varepsilon}}}$ and, being 1-Lipschitz w.r. to $d_{\tilde{\varepsilon}}$, we have $|\nabla_{\tilde{\varepsilon}}\psi_{\tilde{\varepsilon}}|_{\tilde{\varepsilon}}\leq 1$. Testing the equation for $w_{n}$ with $\psi_{n}^{p}w_{n}$, where $\psi_{n}=\psi_{\tilde{\varepsilon}_{n}}$, and using Young’s inequality we have
$\int_{\mathbb{H}^{n}}|\nabla_{\tilde{\varepsilon}_{n}}w_{n}|^{p}_{\tilde{\varepsilon}_{n}}\psi_{n}^{p}\leq p\int_{\mathbb{H}^{n}}|\nabla_{\tilde{\varepsilon}_{n}}w_{n}|^{p-1}_{\tilde{\varepsilon}_{n}}\psi_{n}^{p-1}w_{n}|\nabla_{\tilde{\varepsilon}_{n}}\psi_{\tilde{\varepsilon}_{n}}|_{\tilde{\varepsilon}_{n}}\leq\frac{p-1}{p}\int_{\mathbb{H}^{n}}|\nabla_{\tilde{\varepsilon}_{n}}w_{n}|^{p}_{\tilde{\varepsilon}_{n}}\psi_{n}^{p}+p^{p-1}\int_{\mathbb{H}^{n}}|\nabla_{\tilde{\varepsilon}_{n}}\psi_{n}|^{p}_{\tilde{\varepsilon}_{n}}w_{n}^{p}\,,$
whence we have the following Caccioppoli inequality with uniform constant,
(6.3) $\int_{B_{1}^{\tilde{\varepsilon}_{n}}}|\nabla_{\tilde{\varepsilon}_{n}}w_{n}|^{p}_{\tilde{\varepsilon}_{n}}\leq p^{p}\int_{B_{2}^{\tilde{\varepsilon}_{n}}}w_{n}^{p}\,.$
Finally, combining (6.1), (6.2), and (6.3) with the Harnack inequality (4.16) in Proposition 4.2 (with the choice $2r=4=\bar{r}$) we have
$r^{p}_{n}|\nabla_{\varepsilon}v(g_{n})|^{p}_{\varepsilon}\leq C\int_{B_{1}^{\tilde{\varepsilon}_{n}}(0)}|\nabla_{\tilde{\varepsilon}_{n}}w_{n}|_{\tilde{\varepsilon}_{n}}^{p}\leq C\int_{B_{2}^{\tilde{\varepsilon}_{n}}}w_{n}^{p}\leq C\,\,|B_{2}^{\tilde{\varepsilon}_{n}}|\,\,\underset{B_{2}^{\tilde{\varepsilon}_{n}}}{\rm{ess}\inf}\,\,w^{p}_{n}\leq C(w_{n}(0))^{p}=Cv(g_{n})^{p}\,,$
for constant $C>0$ depending on $p$ but independent of $n$ (compare also Lemma 2.2-$(2)$). Since $r_{n}\to\infty$ as $n\to\infty$ then the conclusion follows. ∎
The next result provides uniform control for the gradient of each Riemannian $p-$capacitary potential at the boundary as a consequence of the exterior gauge-ball condition.
Proposition 6.2.
Let $\Omega\subset\mathbb{H}^{n}$ be an open set with $C^{2}$-smooth boundary and bounded complement satisfying assuption ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}$. For $1<p<Q$ and $\varepsilon\in(0,1]$ let $v=v^{\varepsilon}_{p}\in\dot{W}^{1,p}_{1,\varepsilon}(\Omega)\cap C^{1,\beta}(\overline{\Omega})$ be the solution to (1.9a) corresponding to the unique minimizers of (1.10) in the set $\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$.
Then, there exists $\overline{C}>0$ depending only on $R_{0}$ and $Q$ such that the following inequality on $\partial\Omega$ holds
(6.4) $|\nabla_{\varepsilon}v|_{\varepsilon}\leq\frac{\overline{C}}{p-1}v\,.$
Proof.
The claim will follow from the comparison principle between $v$ and suitable subsolutions constructed in terms of the functions given by (3.30). More precisely, inequality (6.4) will follow with the choice $\overline{C}=C_{0}K$, with $C_{0}=C_{0}(R_{0},Q)$ as in Lemma 3.2 and $K=K(R_{0},Q)$ as in Proposition 3.3.
Let $R_{0}$ be as in assumption ($\mathbf{HP_{\Omega}}$) and for any $g\in\partial\Omega$ let us choose $g_{0}\in\mathbb{H}^{n}$ such that $B_{R_{0}}(g_{0})\subset\Omega^{c}$ and $\partial\Omega\cap\partial B_{R_{0}}(g_{0})=\{g\}$. Now consider $\Phi_{\alpha}=\Phi_{\alpha}^{(g_{0},R_{0})}$ as in (3.30) and $v\in\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ the Riemannian $p$-harmonic function associated to $\Omega$, so that $v\leq 1$ in $\overline{\Omega}$ in view of the the weak maximum principle (equivalently, $v=v\wedge 1$ since truncation decreases the energy). Here $\alpha=-K/(p-1)$ and $K>\frac{Q-p}{4}$, $K=K(R_{0},Q)$, is chosen independent of $p\in(1,Q)$ and such that inequality (3.33) holds for any $g_{0}$ as above and any $\varepsilon\in(0,1]$. Note that by construction $B_{R_{0}}(g_{0})\subset\Omega^{c}$, hence $\Phi_{\alpha}\leq 1$ on $\overline{\Omega}$. On the other hand, by the previous choice of $K$ the function $\Phi_{\alpha}$ is $p$-subharmonic in $\mathbb{H}^{n}\setminus B_{R_{0}}(g_{0})$, hence in $\Omega$, as proved in Proposition 3.3. Thus, by the weak comparison principle for (sub/super) $p$-harmonic functions in $\dot{W}^{1,p}_{\varepsilon}(\Omega)$ we have $\Phi_{\alpha}\leq v$ a.e. in $\Omega$ and in turn everywhere in $\overline{\Omega}$ because such functions are both continuous in $\overline{\Omega}$.
Next, note that in view of ($\mathbf{HP_{\Omega}}$) both $B_{R_{0}}(g_{0})^{c}=\{\Phi_{\alpha}\leq 1\}$ and $\Omega$ have smooth boundaries which are tangent at $g\in\partial\Omega\cap\partial\left(B_{R_{0}}(g_{0})^{c}\right)$, therefore they have the same outer normal vector there. Now, $\Phi_{\alpha}$ (resp. $v$) has zero tangential grandient at $g$ because it is identically one on $\partial\left(B_{R_{0}}(g_{0})^{c}\right)$ (resp. on $\partial\Omega$) by construction. On the other hand since $0<\Phi_{\alpha}\leq v\leq 1$ on $\overline{\Omega}$ then their normal derivatives are ordered as well and in turn $|\nabla_{\varepsilon}v(g)|_{\varepsilon}\leq|\nabla_{\varepsilon}\Phi_{\alpha}|_{\varepsilon}(g)$. Hence, inequality (6.4) on $\partial\Omega$ follows from inequality (3.31) in Lemma 3.2 since $v(g)=1$ and $g\in\partial\Omega$ is arbitrary. ∎
The next result propagates the boundary gradient estimates for the $p$-capacitary potentials $v^{\varepsilon}_{p}$ to the whole exterior domain.
Proposition 6.3.
Let $\Omega\subset\mathbb{H}^{n}$ be an open set with $C^{2}$-smooth boundary and bounded complement such that $\Omega$ satisfies an exterior uniform gauge-ball condition ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}$. For $1<p<Q$ and $0<\varepsilon\leq 1$ let $v=v_{p}^{\varepsilon}\in\dot{W}^{1,p}_{1,\varepsilon}(\Omega)\cap C^{1,\beta}(\overline{\Omega})$ be a weak solution to (1.9a) such that $|\nabla_{\varepsilon}v|_{\varepsilon}=o(v)$ as $\|g\|\to\infty$.
Then
(6.5) $\left\|\frac{|\nabla_{\varepsilon}v|_{\varepsilon}}{v}\right\|_{L^{\infty}(\Omega)}\leq\sqrt{1+n}\,\left\|\frac{|\nabla_{\varepsilon}v|_{\varepsilon}}{v}\right\|_{L^{\infty}(\partial\Omega)}\,.$
The previous result is a consequence of the following stronger claim for the vertical derivative $T_{\varepsilon}v_{p}^{\varepsilon}$.
Lemma 6.4.
Under the assumptions of Proposition 6.3 for $v=v^{\varepsilon}_{p}$ we have the estimate
(6.6) $\left\|\frac{\left(T_{\varepsilon}v_{p}^{\varepsilon}\right)^{2}}{\left(v_{p}^{\varepsilon}\right)^{2}}\right\|_{L^{\infty}(\Omega)}\leq\left\|\frac{\left(T_{\varepsilon}v_{p}^{\varepsilon}\right)^{2}}{\left(v_{p}^{\varepsilon}\right)^{2}}\right\|_{L^{\infty}(\partial\Omega)}\,,$
Proof.
First notice that for $v=v^{\varepsilon}_{p}$ the r.h.side of (6.6) is finite in view of Proposition 6.2. We argue by contradiction and suppose (6.6) is false. Note that in view of the $C^{1,\beta}$-regularity of $v$ in the whole $\overline{\Omega}$ and the assumption $|\nabla_{\varepsilon}v|_{\varepsilon}=o(v)$ as $\|g\|\to\infty$ the left hand side of (6.6) is by continuity a maximum denoted by $M$ in what follows. Being by contradiction an interior maximum, there exists $\bar{g}\in\Omega$ such that
(6.7) $M=\frac{\left(T_{\varepsilon}v\right)^{2}}{v^{2}}(\bar{g})>\left\|\frac{\left(T_{\varepsilon}v\right)^{2}}{v^{2}}\right\|_{L^{\infty}(\partial\Omega)}\,.$
Next, in view of the assumption $|\nabla_{\varepsilon}v|_{\varepsilon}=o(v)$ as $\|g\|\to\infty$ we fix can fix $R>0$ so large that $\bar{g}\in\Omega_{R}:=\Omega\cap B_{R}(0)$ for any maximum point $\bar{g}$ as in (6.7), $\partial B_{R}(0)\cap\partial\Omega=\emptyset$ and $|\nabla_{\varepsilon}v|_{\varepsilon}\leq\frac{M}{2}v$ on $\partial B_{R}(0)$.
Now let us define the auxiliary function
$w=\left(T_{\varepsilon}v\right)^{2}-Mv^{2}\,$
and notice that $w\leq 0$ in $\overline{\Omega_{R}}$, $w(\bar{g})=0$ and $w<0$ on $\partial\Omega_{R}$ because $v>0$ on $\overline{\Omega_{R}}$ and (6.7) holds. Thus, any $\bar{g}$ as in (6.7) is an interior local maximum for $w$ on $\overline{\Omega_{R}}$ and $w(\bar{g})=0$ yields $(T_{\varepsilon}v)^{2}(\bar{g})=M(v(\bar{g}))^{2}\neq 0$, hence $\nabla_{\varepsilon}v(\bar{g})\neq 0$; in particular, $v$ being $C^{1}$ it satisfies $T_{\varepsilon}v\neq 0$ and $\nabla_{\varepsilon}v\neq 0$ near $\bar{g}$ and therefore it is smooth near $\bar{g}$ by elliptic regularity.
Similarly, for any $\sigma>0$ define the auxiliary functions
$w_{\sigma}=\left(T_{\varepsilon}v\right)^{2}-(M+\sigma)v^{2}\,,$
and notice that $w_{\sigma}<0$ in $\overline{\Omega_{R}}$, because $w\leq 0$ and $\min_{\overline{\Omega_{R}}}v>0$. Clearly $w_{\sigma}\to w$ uniformly on $\overline{\Omega_{R}}$, hence for $\sigma$ small enough the set of maximum points for $w_{\sigma}$ is well inside $\Omega_{R}$, as the same property for $w$ has been shown above; furthermore for $\sigma$ small enough there are maxima $\bar{g}_{\sigma}\in\Omega_{R}$ for $w_{\sigma}$ such that up to subsequences $\bar{g}_{\sigma}\to\bar{g}\in\Omega_{R}$ for some $\bar{g}$ as above.
Sufficiently close to $\bar{g}$ let us now consider the following linear (uniformly) elliptic operator
(6.8) $L_{v}(\psi):=\mathrm{div}_{\varepsilon}(A_{v}\nabla_{\varepsilon}\psi)\,,$
where
$A_{v}:=|\nabla_{\varepsilon}v|_{\varepsilon}^{p-2}\left(I_{2n+1}+(p-2)\dfrac{\nabla_{\varepsilon}v}{|\nabla_{\varepsilon}v|_{\varepsilon}}\otimes\dfrac{\nabla_{\varepsilon}v}{|\nabla_{\varepsilon}v|_{\varepsilon}}\right).$
Note that $L_{v}(v)=0$ near $\bar{g}$ because of (1.9a). Moreover, since $v$ is smooth near $\bar{g}$ we can differentiate (1.9a) and since $[X_{i},T]=[Y_{i},T]=0$ for all $i=1,\ldots,n$ we also obtain $L_{v}(T_{\varepsilon}v)=0$ near $\bar{g}$.
Since $L_{v}$ is an elliptic operator and $\bar{g}_{\sigma}$ is an interior local maximum point for $w_{\sigma}$, from the smoothness of $v$ and in turn of $w_{\sigma}$ near $\bar{g}$ (i.e., near $\bar{g}_{\sigma}$ for $\sigma$ small enough) we have
(6.9) $L_{v}(w_{\sigma})(\bar{g}_{\sigma})\leq 0.$
On the other hand, the identities $L_{v}(v)=L_{v}(T_{\varepsilon}v)=0$ and $L_{v}(f^{2})=2fL_{v}(f)+2\langle{A_{v}\nabla_{\varepsilon}f},{\nabla_{\varepsilon}f}\rangle_{\varepsilon}$ for a sufficiently smooth function $f$, yield
(6.10) $L_{v}(w_{\sigma})=L_{v}\left((T_{\varepsilon}v)^{2}-(M+\sigma)v^{2}\right)=2\langle{A_{v}\nabla_{\varepsilon}T_{\varepsilon}v},{\nabla_{\varepsilon}T_{\varepsilon}v}\rangle_{\varepsilon}-2(M+\sigma)(p-1)|\nabla_{\varepsilon}v|_{\varepsilon}^{p}\,.$
Since $\bar{g}_{\sigma}$ is an interior local maximum point for $w_{\sigma}$, at that point $\nabla_{\varepsilon}w_{\sigma}=0$ yields $\nabla_{\varepsilon}T_{\varepsilon}v=(M+\sigma)\dfrac{v\nabla_{\varepsilon}v}{T_{\varepsilon}v}\,,$ therefore, again in $\bar{g}_{\sigma}$, we get
$2\langle{A_{v}\nabla_{\varepsilon}Tv},{\nabla_{\varepsilon}Tv}\rangle_{\varepsilon}=2(p-1)(M+\sigma)^{2}\dfrac{v^{2}}{(Tv)^{2}}|\nabla_{\varepsilon}v|_{\varepsilon}^{p}\,.$
Combining the last identity with (6.9) and (6.10) together with $w_{\sigma}(\bar{g}_{\sigma})<0$ we obtain
$0\geq L_{v}(w_{\sigma})(\bar{g}_{\sigma})=2(p-1)(M+\sigma)|\nabla_{\varepsilon}v(\bar{g}_{\sigma})|_{\varepsilon}^{p}\left((M+\sigma)\dfrac{v^{2}}{(T_{\varepsilon}v)^{2}}(\bar{g}_{\sigma})-1\right)>0,$
because $(T_{\varepsilon}v)^{2}(\bar{g}_{\sigma})\leq Mv^{2}(\bar{g}_{\sigma})$. Thus, the contradiction proves that $\dfrac{(T_{\varepsilon}v)^{2}}{v^{2}}$ attains its maximum at the boundary $\partial\Omega$ as desired. ∎
Before giving the proof of Proposition 6.3 we need the following auxiliary identity for solutions to (1.9a). The proof here is given just for the sake of completeness, as it is a mere adaptation of the one in [30, Lemma 2.1] where exactly the same identity is established but for solutions to (1.6a) instead of (1.9a).
Lemma 6.5.
Let $1<p<Q$, $\varepsilon>0$, $\widehat{\Omega}\subset\mathbb{H}^{n}$ an open set and $v\in C^{3}(\widehat{\Omega})$ such that $\nabla_{\varepsilon}v\neq 0$ in $\widehat{\Omega}$. If $v$ is $p$-harmonic in $\widehat{\Omega}$ and $L_{v}$ is the linear operator defined as in (6.8) then
(6.11) $L_{v}(|\nabla_{\varepsilon}v|_{\varepsilon}^{2})=|\nabla_{\varepsilon}v|_{\varepsilon}^{p-2}\left(2|D^{2}_{\varepsilon}v|_{\varepsilon}^{2}+2Ric_{\varepsilon}(\nabla_{\varepsilon}v)\right)+\dfrac{p-2}{2}|\nabla_{\varepsilon}v|_{\varepsilon}^{p-4}|\nabla_{\varepsilon}|\nabla_{\varepsilon}v|_{\varepsilon}^{2}|_{\varepsilon}^{2}\,,$
where $D^{2}_{\varepsilon}v$ is the covariant Hessian and $Ric_{\varepsilon}$ is the Ricci (2,0)-tensor corresponding to the Riemannian metric $|\cdot|^{2}_{\varepsilon}$.
Proof.
First notice that $p$-harmonicity of $v$ together with the assumption $\nabla_{\varepsilon}v\neq 0$ in $\widehat{\Omega}$ give $v\in C^{\infty}(\widehat{\Omega})$ by elliptic regularity. As $\varepsilon>0$ is fixed, for simplicity we drop the dependence on $\varepsilon$ in the sequel and, following the notation in [30], we let $f:=|\nabla v|^{2}$, $v_{ij}$ the Hessian of $v$ and $Ric_{ij}$ the Ricci curvature w.r. to any fixed local orthonormal frame. Then the equation (1.9a) satisfied by $v$ yields the identity
$f^{p/2-1}\Delta v+\left(\dfrac{p}{2}-1\right)f^{p/2-2}\langle\nabla f,\nabla v\rangle=0\,.$
Taking the gradient on both sides of the latter, and then computing its scalar product with $\nabla v$ we obtain
(6.12) $\left(\dfrac{p}{2}-1\right)f^{p/2-2}\Delta v\langle\nabla f,\nabla v\rangle+f^{p/2-1}\langle\nabla\Delta v,\nabla v\rangle$
$+\left(\dfrac{p}{2}-1\right)\left(\dfrac{p}{2}-2\right)f^{p/2-3}\langle\nabla f,\nabla v\rangle^{2}+\left(\dfrac{p}{2}-1\right)f^{p/2-2}(f_{ij}v_{i}v_{j}+v_{ij}f_{i}v_{j})=0.$
Now, following almost verbatim the computations made in [30, Lemma 2.1], we get
(6.13) $L_{v}(f)$ $=\left(\dfrac{p}{2}-1\right)f^{p/2-2}|\nabla f|^{2}+f^{p/2-1}\Delta f+(p-2)\Delta v\langle\nabla v,\nabla f\rangle f^{p/2-2}$
$+(p-2)\left(\dfrac{p}{2}-1\right)f^{p/2-3}\langle\nabla v,\nabla f\rangle^{2}$
$+(p-2)(v_{ij}f_{i}v_{j}f^{p/2-2}+f_{ij}v_{i}v_{j}f^{p/2-2}-\langle\nabla v,\nabla f\rangle^{2}f^{p/2-3})\,.$
On the other hand, recalling that the Bochner’s identity gives
$\Delta f=2|D^{2}v|^{2}+2\langle\nabla\Delta v,\nabla v\rangle+2Ric(\nabla v)\,,$
inserting this identity in (6.13) we get
$L_{v}(f)$ $=f^{p/2-1}\left(2|D^{2}v|^{2}+2Ric(\nabla v)\right)+\dfrac{p-2}{2}f^{p/2-2}|\nabla f|^{2}$
$+(p-2)f^{p/2-2}\Delta v\langle\nabla v,\nabla f\rangle+2f^{p/2-1}\langle\nabla\Delta v,\nabla v\rangle$
$+(p-2)\left(\dfrac{p}{2}-2\right)f^{p/2-3}\langle\nabla v,\nabla f\rangle^{2}+(p-2)f^{p/2-2}(v_{ij}f_{i}v_{j}+f_{ij}v_{i}v_{j})\,.$
Finally, taking (6.12) into account identity (6.11) follows from the previous formula as the last two lines vanish identically. ∎
Before presenting the proof of Proposition 6.3, we state a technical result which provides an explicit formula for the $Ric_{\varepsilon}$ tensor.
Lemma 6.6.
Let $U,V\in\mathfrak{h}$ be two vector fields. Then the Ricci tensor is given by
(6.14) $Ric_{\varepsilon}(U,V)=\dfrac{1}{2\varepsilon^{2}}\langle U,V\rangle_{\varepsilon}-\dfrac{n+1}{2\varepsilon^{2}}\langle U,T_{\varepsilon}\rangle_{\varepsilon}\,\langle V,T_{\varepsilon}\rangle_{\varepsilon}.$
In particular, ${Ric_{\varepsilon}(U):=Ric_{\varepsilon}(U,U)=\frac{1}{2\varepsilon^{2}}}|U|_{\varepsilon}^{2}-\frac{n+1}{2\varepsilon^{2}}|<U,T_{\varepsilon}>_{\varepsilon}|^{2}$ for any $U\in\mathfrak{h}$.
Proof.
We write $U,V\in\mathfrak{h}$ as linear combinations with constant coefficients,
$U=\sum_{i=1}^{n}\left(u_{i}X_{i}+u_{i+n}Y_{i}\right)+u_{2n+1}T_{\varepsilon}\quad\textrm{and}\quad V=\sum_{i=1}^{n}\left(v_{i}X_{i}+v_{i+n}Y_{i}\right)+v_{2n+1}T_{\varepsilon}.$
Taking into account the commutation rules recalled in Section 2, it follows that
$[U,X_{i}]=-\dfrac{u_{i+n}}{\varepsilon}T_{\varepsilon},\quad[U,Y_{i}]=\dfrac{u_{i}}{\varepsilon}T_{\varepsilon}\quad\textrm{and}\quad[U,T_{\varepsilon}]=0.$
Moreover, the action of $\nabla^{\varepsilon}$ on the orthonormal base recalled in Section 2 yields
$\nabla^{\varepsilon}_{U}V=\sum_{i=1}^{n}\left(\dfrac{u_{i+n}v_{2n+1}+u_{2n+1}v_{i+n}}{2\varepsilon}\right)X_{i}-\sum_{i=1}^{n}\left(\dfrac{u_{i}v_{2n+1}+u_{2n+1}v_{i}}{2\varepsilon}\right)Y_{i}+\sum_{i=1}^{n}\left(\dfrac{v_{i+n}u_{i}-u_{i+n}v_{i}}{2\varepsilon}\right)T_{\varepsilon}.$
Now recall that for $U,V\in\mathfrak{h}$ the corresponding curvature operator $R^{\varepsilon}(U,V):\mathfrak{h}\to\mathfrak{h}$ is given by
(6.15) $R^{\varepsilon}(U,V)W=\nabla^{\varepsilon}_{U}\nabla^{\varepsilon}_{V}W-\nabla^{\varepsilon}_{V}\nabla^{\varepsilon}_{U}W-\nabla^{\varepsilon}_{[U,V]}W\,,\quad W\in\mathfrak{h},$
so that we have to compute
$Ric_{\varepsilon}(U,V)=\sum_{i=1}^{n}\left(\langle R^{\varepsilon}(U,X_{i})V,X_{i}\rangle_{\varepsilon}+\langle R^{\varepsilon}(U,Y_{i})V,Y_{i}\rangle_{\varepsilon}\right)+\langle R^{\varepsilon}(U,T_{\varepsilon})V,T_{\varepsilon}\rangle_{\varepsilon}.$
To this aim, recalling (6.15) and the explicit computations made for the Levi-Civita connection $\nabla^{\varepsilon}$, it follows that
$\nabla^{\varepsilon}_{X_{i}}V$ $=-\dfrac{v_{2n+1}}{2\varepsilon}Y_{i}+\dfrac{v_{i+n}}{2\varepsilon}T_{\varepsilon},\quad\textrm{for every }i=1,\ldots,n,$
$\nabla^{\varepsilon}_{U}\nabla^{\varepsilon}_{X_{i}}V$ $=-\dfrac{u_{2n+1}v_{2n+1}}{4\varepsilon^{2}}X_{i}+\dfrac{v_{i+n}}{4\varepsilon^{2}}\sum_{j=1}^{n}u_{j+n}X_{j}-\dfrac{v_{i+n}}{4\varepsilon^{2}}\sum_{j=1}^{n}u_{j}Y_{j}-\dfrac{v_{2n+1}u_{i}}{4\varepsilon^{2}}T_{\varepsilon},\quad\textrm{for every }i=1,\ldots,n,$
$\nabla^{\varepsilon}_{X_{i}}\nabla^{\varepsilon}_{U}V$ $=-\sum_{j=1}^{n}\left(\dfrac{u_{j}v_{j+n}-v_{j}u_{j+n}}{4\varepsilon^{2}}\right)Y_{i}-\left(\dfrac{u_{i}v_{2n+1}+v_{i}u_{2n+1}}{4\varepsilon^{2}}\right)T_{\varepsilon},\quad\textrm{for every }i=1,\ldots,n,$
[$\nabla^{\varepsilon}_{[U,X_{i}]}V$ $=-\dfrac{u_{i+n}}{2\varepsilon^{2}}\sum_{j=1}^{n}v_{j+n}X_{j}+\dfrac{u_{i+n}}{2\varepsilon^{2}}\sum_{j=1}^{n}v_{j}Y_{j},\quad\textrm{for every }i=1,\ldots,n,$
and hence
(6.16) $\sum_{i=1}^{n}\langle R^{\varepsilon}(U,X_{i})V,X_{i}\rangle_{\varepsilon}=\dfrac{3}{4\varepsilon^{2}}\sum_{i=1}^{n}u_{i+n}v_{i+n}-\dfrac{n}{4\varepsilon^{2}}u_{2n+1}v_{2n+1}.$
Similarly,
$\nabla^{\varepsilon}_{Y_{i}}V$ $=\dfrac{v_{2n+1}}{2\varepsilon}X_{i}-\dfrac{v_{i}}{2\varepsilon}T_{\varepsilon},\quad\textrm{for every }i=1,\ldots,n,$
$\nabla^{\varepsilon}_{U}\nabla^{\varepsilon}_{Y_{i}}V$ $=-\dfrac{v_{i}}{4\varepsilon^{2}}\sum_{j=1}^{n}u_{j+n}X_{j}-\dfrac{u_{2n+1}v_{2n+1}}{4\varepsilon^{2}}Y_{i}+\dfrac{v_{i}}{4\varepsilon^{2}}\sum_{j=1}^{n}u_{j}Y_{j}-\dfrac{v_{2n+1}u_{i+n}}{4\varepsilon^{2}}T_{\varepsilon},\quad\textrm{for every }i=1,\ldots,n,$
$\nabla^{\varepsilon}_{Y_{i}}\nabla^{\varepsilon}_{U}V$ $=\sum_{j=1}^{n}\left(\dfrac{u_{j}v_{j+n}-v_{j}u_{j+n}}{4\varepsilon^{2}}\right)X_{i}-\left(\dfrac{u_{i+n}v_{2n+1}+v_{i+n}u_{2n+1}}{4\varepsilon^{2}}\right)T_{\varepsilon},\quad\textrm{for every }i=1,\ldots,n,$
[$\nabla^{\varepsilon}_{[U,X_{i}]}V$ $=\dfrac{u_{i}}{2\varepsilon^{2}}\sum_{j=1}^{n}v_{j+n}X_{j}-\dfrac{u_{i}}{2\varepsilon^{2}}\sum_{j=1}^{n}v_{j}Y_{j},\quad\textrm{for every }i=1,\ldots,n,$
and hence
(6.17) $\sum_{i=1}^{n}\langle R^{\varepsilon}(U,Y_{i})V,Y_{i}\rangle_{\varepsilon}=\dfrac{3}{4\varepsilon^{2}}\sum_{i=1}^{n}u_{i}v_{i}-\dfrac{n}{4\varepsilon^{2}}u_{2n+1}v_{2n+1}.$
Finally,
$\nabla^{\varepsilon}_{T_{\varepsilon}}V$ $=\sum_{j=1}^{n}\dfrac{v_{j+n}}{2\varepsilon}X_{j}-\sum_{j=1}^{n}\dfrac{v_{j}}{2\varepsilon}Y_{j},$
$\nabla^{\varepsilon}_{U}\nabla^{\varepsilon}_{T_{\varepsilon}}V$ $=-\dfrac{u_{2n+1}}{4\varepsilon^{2}}\sum_{j=1}^{n}v_{j}X_{j}-\dfrac{u_{2n+1}}{4\varepsilon^{2}}\sum_{j=1}^{n}v_{j+n}Y_{j}-\sum_{j=1}^{n}\left(\dfrac{u_{j}v_{j}+u_{j+n}v_{j+n}}{4\varepsilon^{2}}\right)T_{\varepsilon},$
$\nabla^{\varepsilon}_{T_{\varepsilon}}\nabla^{\varepsilon}_{U}V$ $=-\sum_{j=1}^{n}\dfrac{u_{j}v_{2n+1}+u_{2n+1}v_{j}}{4\varepsilon^{2}}X_{j}-\sum_{j=1}^{n}\dfrac{u_{j+n}v_{2n+1}+u_{2n+1}v_{j+n}}{4\varepsilon^{2}}Y_{j},$
[$\nabla^{\varepsilon}_{[U,T_{\varepsilon}]}V=0,$
and hence
(6.18) $\langle R^{\varepsilon}(U,T_{\varepsilon})V,T_{\varepsilon}\rangle_{\varepsilon}=-\dfrac{1}{4\varepsilon^{2}}\sum_{i=1}^{n}(u_{i}v_{i}+u_{i+n}v_{i+n}).$
Combinign (6.16), (6.17) and (6.18), we finally get (6.14) as desired. ∎
Proof of Proposition 6.3.
As above we set for brevity $v=v_{p}^{\varepsilon}$ in what follows. First notice that the r.h.side for the claimed inequality is finite in view of Proposition 6.2. We argue by contradiction and suppose the claim is false. Note that, as for Proposition 6.4, in view of the $C^{1,\beta}$-regularity of $v$ in the whole $\overline{\Omega}$ and the assumption $|\nabla_{\varepsilon}v|_{\varepsilon}=o(v)$ as $\|g\|\to\infty$ the left hand side of (6.5) is by continuity and the contradiction assumption indeed an interior maximum denoted by $\hat{M}$ in what follows. More precisely, by contradiction assumption, there exists $\hat{g}\in\Omega$ such that
(6.19) $\hat{M}:=\frac{|\nabla_{\varepsilon}v|_{\varepsilon}^{2}}{v^{2}}(\hat{g})=\left\|\frac{|\nabla_{\varepsilon}v|_{\varepsilon}}{v}\right\|^{2}_{L^{\infty}(\Omega)}>(n+1)\left\|\frac{|\nabla_{\varepsilon}v|_{\varepsilon}}{v}\right\|^{2}_{L^{\infty}(\partial\Omega)}\,.$
In analogy with the proof of Lemma 6.4 we consider the auxiliary function
$\hat{w}=|\nabla_{\varepsilon}v|^{2}_{\varepsilon}-\hat{M}v^{2}\,,$
so that by the definition of $\hat{M}$ we have $\hat{w}\leq 0$ in $\Omega$, $\hat{w}(\hat{g})=0$ and $\nabla_{\varepsilon}v(\hat{g})\neq 0$ because $v>0$ in $\overline{\Omega}$ by the strong maximum principle. As above, the latter condition assures that $v$ is smooth near $\hat{g}$ applying elliptic regularity theory to (1.9a). In particular, stationarity condition yields $\nabla_{\varepsilon}w(\hat{g})=0$ which in turn implies at the point $\hat{g}$ the identity
(6.20) $|\,\nabla_{\varepsilon}|\nabla_{\varepsilon}v|_{\varepsilon}^{2}|_{\varepsilon}^{2}=4\hat{M}^{2}v^{2}|\nabla_{\varepsilon}v|_{\varepsilon}^{2}\,.$
Applying once more the elliptic operator $L_{v}$ defined in (6.8) we have at the maximum point $\hat{g}$ the inequality
(6.21) $L_{v}(\hat{w})(\hat{g})\leq 0\,.$
On the other hand, since $v$ is smooth near $\hat{g}$ we can apply Lemma 6.5, hence identity (6.11) holds near $\hat{g}$. Thus, recalling that $2|D^{2}_{\varepsilon}v|_{\varepsilon}^{2}\geq\tfrac{1}{2}\tfrac{|\nabla_{\varepsilon}|\nabla_{\varepsilon}v|_{\varepsilon}^{2}|_{\varepsilon}^{2}}{|\nabla_{\varepsilon}v|_{\varepsilon}^{2}}$ and arguing as in (6.10) near $\hat{g}$ we have,
$L_{v}(\hat{w})$ $=L_{v}(|\nabla_{\varepsilon}v|_{\varepsilon}^{2})-\hat{M}L_{v}(v^{2})$
$=|\nabla_{\varepsilon}v|_{\varepsilon}^{p-2}\left(2|D^{2}_{\varepsilon}v|_{\varepsilon}^{2}+2Ric_{\varepsilon}\left(\nabla_{\varepsilon}v\right)\right)+\dfrac{p-2}{2}|\nabla_{\varepsilon}v|_{\varepsilon}^{p-4}|\nabla_{\varepsilon}|\nabla_{\varepsilon}v|_{\varepsilon}^{2}|_{\varepsilon}^{2}-2\hat{M}(p-1)|\nabla_{\varepsilon}v|_{\varepsilon}^{p}$
$\geq 2|\nabla_{\varepsilon}v|_{\varepsilon}^{p-2}Ric_{\varepsilon}(\nabla_{\varepsilon}v)+\dfrac{(p-1)}{2}|\nabla_{\varepsilon}v|_{\varepsilon}^{p-4}|\nabla_{\varepsilon}|\nabla_{\varepsilon}v|_{\varepsilon}^{2}|_{\varepsilon}^{2}-2\hat{M}(p-1)|\nabla_{\varepsilon}v|_{\varepsilon}^{p}\,.$
In particular, evaluating the last inequality at the point $\hat{g}$ and taking into account (6.20) and (6.21), since $\hat{w}(\hat{g})=0$ we get
$0\geq L_{v}(\hat{w})(\hat{g})\geq 2|\nabla_{\varepsilon}v(\hat{g})|_{\varepsilon}^{p-2}Ric_{\varepsilon}\left(\nabla_{\varepsilon}v(\hat{g})\right)\,,$
which in view of the explicit structure of $Ric_{\varepsilon}$ (see (6.14) with $U=V=\nabla_{\varepsilon}v$) finally rewrites as
(6.22) $|\nabla_{0}v(\hat{g})|_{0}^{2}\leq n\left(T_{\varepsilon}v(\hat{g})\right)^{2}\,.$
To finish the proof, it is enough to combine (6.22) with Lemma 6.4. Indeed, it follows from the definition of $\hat{M}$ in (6.19), inequality (6.22), the structure of $|\cdot|_{\varepsilon}^{2}$ and inequality (6.4) that
(6.23) $\left\|\frac{|\nabla_{\varepsilon}v|_{\varepsilon}}{v}\right\|^{2}_{L^{\infty}(\Omega)}$ $=\dfrac{|\nabla_{\varepsilon}v|_{\varepsilon}^{2}}{v^{2}}(\hat{g})\leq\left(1+n\right)\dfrac{(T_{\varepsilon}v)^{2}}{v^{2}}(\hat{g})\leq\left(1+n\right)\left\|\frac{\left(T_{\varepsilon}v\right)^{2}}{v^{2}}\right\|_{L^{\infty}(\partial\Omega)}$
$\leq\left(1+n\right)\left\|\dfrac{|\nabla_{\varepsilon}v|_{\varepsilon}}{v}\right\|^{2}_{L^{\infty}(\partial\Omega)}\,.$
Finally, since (6.23) contradicts (6.19) the conclusion follows. ∎
7. Proof of the main results
In order to prove Theorem 1.2 we start with the following auxiliary result.
Proposition 7.1.
Let $1<p<Q$, $\varepsilon\in[0,1]$, and let $\Omega\subset\mathbb{H}^{n}$ be an open set with bounded complement and smooth boundary $\partial\Omega$. If $v^{\varepsilon}_{p}$ are the minimizers of (1.10) in the set $\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ then $v^{\varepsilon}_{p}\to v^{0}_{p}$ strongly in $\dot{W}^{1,p}_{1,0}(\Omega)$ and, up to subsequences, $|\nabla_{\varepsilon}v^{\varepsilon}_{p}|_{\varepsilon}\to|\nabla_{0}v^{0}_{p}|_{0}$ a.e. as $\varepsilon\to 0$.
Proof.
First we fix $R>0$ so large that $\mathbb{H}^{n}\setminus\Omega\subset B_{R}(0)$. In view of Lemma 3.2, we have $\tilde{\Phi}^{(0,R)}_{\alpha}\in\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ with finite norms for each $\varepsilon\in(0,1]$, and therefore for each $\varepsilon\in(0,1]$ the function $\tilde{\Phi}^{(0,R)}_{\alpha}$ is admissible competitor for each $v^{\varepsilon}_{p}$. Hence, it follows from energy minimality that
$E^{0}_{p}\left(v^{\varepsilon}_{p}\right)\leq E^{\varepsilon}_{p}\left(v^{\varepsilon}_{p}\right)\leq E^{\varepsilon}_{p}\left(\tilde{\Phi}^{(0,R)}_{\alpha}\right)\leq E^{1}_{p}\left(\tilde{\Phi}^{(0,R)}_{\alpha}\right)<\infty$
for each $\varepsilon\in(0,1]$, i.e., the sequence $\{v^{\varepsilon}_{p}\}_{\varepsilon}\subset\dot{W}^{1,p}_{1,0}(\Omega)$ is bounded.
Next, for any sequence $\varepsilon_{j}\to 0$ as $j\to+\infty$, possibly passing to a subsequence, there exists $v_{*}\in H\dot{W}^{1,p}(\mathbb{H}^{n})$ such that $v^{\varepsilon_{j}}_{p}$ converges to $v_{*}$ weakly in $H\dot{W}^{1,p}(\mathbb{H}^{n})$, strongly in $L^{1}_{\mathrm{loc}}$ and a.e. in $\mathbb{H}^{n}$, so that in particular we deduce that $v_{*}\in\dot{W}^{1,p}_{1,0}(\Omega)$. It is now enough to prove the convergence in norm together with the energy minimality of $v_{*}$ on $\dot{W}^{1,p}_{1,0}(\Omega)$ , whence $v_{*}=v_{p}^{\varepsilon}$ by strict convexity of $E^{0}_{p}(\cdot)$ and in turn the full convergence holds as $\varepsilon\downarrow 0$.
Note that for every $\hat{v}\in\dot{W}^{1,p}_{1,0}(\Omega)$, since $\Omega$ is smooth with bounded complement and $\hat{v}-\tilde{\Phi}^{(0,R)}_{\alpha}=0$ a.e. on $\mathbb{H}^{n}\setminus\Omega$, there exists a sequence $\varphi_{j}\subset C^{\infty}_{0}(\Omega)$ such that $\varphi_{j}\to\hat{v}-\overline{\Phi}^{(0,R_{0})}_{\alpha}$ in $H\dot{W}^{1,p}(\mathbb{H}^{n})$ as $j\to+\infty$. Since $\varphi_{m}+\tilde{\Phi}^{(0,R)}_{\alpha}\in\dot{W}^{1,p}_{1,0}(\Omega)$, the weak lower semicontinuity of $E^{0}_{p}(\cdot)$, the monotonicity of $\varepsilon\to E^{\varepsilon}_{p}(\cdot)$ and the energy minimality of each $v_{p}^{\varepsilon}$ yield for each $m\geq 1$ and in turn as $m\to\infty$ that
(7.1) $E^{0}_{p}(v_{*})\leq\liminf_{j\to+\infty}E_{p}^{0}(v_{p}^{\varepsilon_{j}})\leq\limsup_{j\to+\infty}E_{p}^{0}(v_{p}^{\varepsilon_{j}})\leq\limsup_{j\to+\infty}E_{p}^{\varepsilon_{j}}(v_{p}^{\varepsilon_{j}})$
$\leq$ $\limsup_{m\to\infty}\limsup_{j\to+\infty}E_{p}^{\varepsilon_{j}}\left(\varphi_{m}+\tilde{\Phi}^{(0,R)}_{\alpha}\right)=\lim_{m\to\infty}E_{p}^{0}\left(\varphi_{m}+\tilde{\Phi}^{(0,R)}_{\alpha}\right)=E^{0}_{p}(\hat{v}).$
Since $\hat{v}\in\dot{W}_{1,0}^{1,p}(\Omega)$ is arbitrary, the latter shows that $v_{*}$ is actually the minimizer and hence $v_{*}=v_{p}^{0}$. Moreover, chosing $\hat{v}=v^{0}_{p}$ in (7.1) we infer that $T_{\varepsilon}v_{p}^{\varepsilon}\to 0$ in $L^{p}(\mathbb{H}^{n})$ as $\varepsilon\to 0$ and $E^{0}_{p}(v_{p}^{\varepsilon})\to E^{0}_{p}(v^{0}_{p})$ as $\varepsilon\to 0$, from which the desired strong convergence follows. Finally, once $\nabla_{0}v^{\varepsilon}_{p}\to\nabla_{0}v^{0}_{p}$ in $L^{p}$, passing to a subsequence we also get $|\nabla_{\varepsilon}v^{\varepsilon}_{p}|_{\varepsilon}\to|\nabla_{0}v^{0}_{p}|_{0}$ a.e. as $\varepsilon\to 0$. ∎
Now we are ready to prove the first result of the Introduction.
Proof of Theorem 1.2.
Let $1<p<Q$ be fixed and for $\varepsilon\in(0,1]$ let us consider $v^{\varepsilon}_{p}\in\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ the minimizer of (1.10) in the set (1.11) which is the unique finite energy solution to (1.9a). As already recalled in the introduction, since $\Omega$ has $C^{2}$-smooth boundary by the interior and boundary regularity theory for $p$-harmonic function we have $v^{\varepsilon}_{p}\in C^{1}(\overline{\Omega})$ (the gradient $\nabla_{\varepsilon}v^{\varepsilon}_{p}$ being actually Hölder continuous up to the boundary) and it is strictly positive in $\overline{\Omega}$ by the strong maximum principle.
Next, we are going to prove that for $\varepsilon\in(0,1]$ when the domain $\Omega$ satisfies the uniform gauge-ball condition $(\bf{HP_{\Omega}})$ with parameter $R_{0}>0$ each solution $v^{\varepsilon}_{p}$ satisfies (1.12) for a constant $C>0$ depending on $R_{0}$ and $Q$. To do this, first notice that the inequality holds on $\partial\Omega$ in view of Proposition 6.2 with some constant $\widetilde{C}>0$ depending only on $R_{0}$ and $Q$. Moreover, in view of Proposition 6.1 we have $|\nabla_{\varepsilon}v^{\varepsilon}_{p}|_{\varepsilon}=o(v^{\varepsilon}_{p})$ as $\|g\|\to\infty$, therefore we can apply Proposition 6.3 and the inequality (1.12) follows just by taking $C=\sqrt{2}\widetilde{C}$.
Finally, if $\varepsilon=0$ and $v^{0}_{p}\in\dot{W}^{1,p}_{1,0}(\Omega)$ is the unique finite energy solution to (1.9a) corresponding to the minimizer of (1.10) then the interior regularity theory for $p$-harmonic function in the sub-Riemannian Heisenberg group proved in [37] yields $v^{0}_{p}\in C(\Omega)$ and $\nabla_{0}v^{0}_{p}\in C(\Omega)$. Moreover $v^{\varepsilon}_{p}\to v^{0}_{p}$ as $\varepsilon\to 0$ strongly in $\dot{W}^{1,p}_{1,0}(\Omega)$ and $|\nabla_{\varepsilon}v^{\varepsilon}_{p}|_{\varepsilon}\to|\nabla_{0}v^{\varepsilon}_{p}|_{0}$ a.e. in $\Omega$ because of Proposition 7.1. Since $d_{0}\leq d_{\varepsilon}$ for any $\varepsilon>0$, the functions $\{v^{\varepsilon}_{p}\}_{\varepsilon\in(0,1]}$ are $d_{0}$-equiLipschitz in $\mathbb{H}^{n}$ due to the bound (1.12) and the fact that $0<v^{\varepsilon}_{p}\leq 1$ on $\overline{\Omega}$. As a consequence $v^{\varepsilon}_{p}\to v^{0}_{p}$ locally uniformly on $\mathbb{H}^{n}$ by the Ascoli-Arzelà Theorem, hence $v^{0}_{p}\in C(\overline{\Omega})$, and (1.12) follows for $v^{0}_{p}$ from the same inequality for each $v^{\varepsilon}_{p}$ as $\varepsilon\to 0$. ∎
In order to prove the main result of the paper, i.e., Theorem 1.4, we need three auxiliary results. The first one relates weak solutions of (1.9a) and (1.6a).
Lemma 7.2.
Let $1<p<Q$, $\varepsilon\in[0,1]$ and $v^{\varepsilon}_{p}\in W^{1,p}_{\varepsilon}(\Omega)$ nonegative and nonconstant. If $v^{\varepsilon}_{p}$ is a weak solution of (1.9a), i.e., for test functions $\psi\in W^{1,p}_{\varepsilon}(\Omega)$ with compact support in $\Omega$, then $v^{\varepsilon}_{p}$ is continuous in $\Omega$, it is strictly positive and $\nabla_{\varepsilon}v^{\varepsilon}_{p}$ is continuous in $\Omega$. As a consequence, if $v^{\varepsilon}_{p}\not\equiv 0$ then the function $u^{\varepsilon}_{p}:=(1-p)\log(v^{\varepsilon}_{p})$ is well-defined, it is $d_{\varepsilon}$-locally Lipschitz and it is a weak solution of (1.6a), i.e., for test functions $\varphi\in W^{1,p}_{\varepsilon}(\Omega)\cap L^{\infty}(\Omega)$ with compact support in $\Omega$.
Proof.
First, to simplify the notation we just write throghout the proof $v=v^{\varepsilon}_{p}$ and $u=u^{\varepsilon}_{p}$. Note that by regularity theory recalled in the introduction both $v$ and $\nabla_{\varepsilon}v$ are continuous in $\Omega$, moreover $v>0$ in $\overline{\Omega}$ by the strong maximum principle.
Under this assumption we take $\varphi\in W^{1,p}_{\varepsilon}(\Omega)\cap L^{\infty}(\Omega)$ with compact support and we consider $\psi=\dfrac{\varphi}{v^{p-1}}$. Since $v>0$ in $\Omega$, $\psi\in W^{1,p}_{\varepsilon}(\Omega)$ and it is an admissible test function for (1.9a), so that
(7.2) $0=\int_{\Omega}|\nabla_{\varepsilon}v|_{\varepsilon}^{p-2}\langle\nabla_{\varepsilon}v,\nabla_{\varepsilon}\left(\dfrac{\varphi}{v^{p-1}}\right)\rangle_{\varepsilon}=\int_{\Omega}\dfrac{|\nabla_{\varepsilon}v|_{\varepsilon}^{p-2}}{v^{p-1}}\langle\nabla_{\varepsilon}v,\nabla_{\varepsilon}\varphi\rangle_{\varepsilon}-(p-1)\int_{\Omega}\dfrac{|\nabla_{\varepsilon}v|_{\varepsilon}^{p}}{v^{p}}\varphi\,.$
Clearly $u=(1-p)\log v$ is $d_{\varepsilon}$-locally Lipschitz by construction, because $\nabla_{\varepsilon}u=(1-p)\tfrac{\nabla_{\varepsilon}v}{v}$, and $u$ being a solution of (1.6a) amounts to
$-\int_{\Omega}|\nabla_{\varepsilon}u|^{p-2}\langle\nabla_{\varepsilon}u,\nabla_{\varepsilon}\varphi\rangle_{\varepsilon}=\int_{\Omega}|\nabla u_{\varepsilon}|_{\varepsilon}^{p}\varphi\,,$
for every test function $\varphi\in W^{1,p}_{\varepsilon}(\Omega)\cap L^{\infty}(\Omega)$ with compact support in $\Omega$. Finally, since the latter equation can be inferred from (7.2) up to multiply it by $(p-1)^{p-1}$ and $\varphi$ is arbitrary the proof is complete. ∎
The next lemma describes solutions to (1.6a) in terms of the minimality property for the associated energy functional (1.7).
Lemma 7.3.
Let $1<p<Q$, $\varepsilon\in[0,1]$ and $\Omega\subset\mathbb{H}^{n}$ an open set. For a function $u^{\varepsilon}_{p}\in W^{1,p}_{\varepsilon,loc}(\Omega)\cap L^{\infty}_{loc}(\Omega)$ the following are equivalent
  • (i)
    the function $u^{\varepsilon}_{p}$ is a weak solution of (1.6a) with respect to test functions in $W^{1,p}_{\varepsilon}(\Omega)\cap L^{\infty}(\Omega)$ with compact support in $\Omega$;
  • (ii)
    for every compact set $K\subset\Omega$ the function $u^{\varepsilon}_{p}$ satisfies $J_{u_{p}^{\varepsilon}}^{p,\varepsilon}(u_{p}^{\varepsilon};K)\leq J_{u_{p}^{\varepsilon}}^{p,\varepsilon}(w;K)$ for every competitor $w\in W^{1,p}_{\varepsilon,loc}(\Omega)\cap L^{\infty}_{loc}(\Omega)$ such that $w=u_{p}^{\varepsilon}$ a.e. in $\Omega\setminus K$.
Proof.
As for the previous Lemma, we simply write $u=u^{\varepsilon}_{p}$ wherever it appear. We start showing that $(i)\Rightarrow(ii)$ arguing as in [34, 15]. Since $u$ is a weak solution of (1.6a), for every compact set $K\subset\Omega$ it holds that
(7.3) $\int_{K}|\nabla_{\varepsilon}u|_{\varepsilon}^{p-2}\langle\nabla_{\varepsilon}u,\nabla_{\varepsilon}\varphi\rangle_{\varepsilon}=-\int_{K}|\nabla_{\varepsilon}u|_{\varepsilon}^{p}\varphi$
for every $\varphi\in W^{1,p}_{\varepsilon}(\Omega)\cap L^{\infty}(\Omega)$ such that $\varphi=0$ a.e. in $\Omega\setminus K$. Now, for $w\in W^{1,p}_{\varepsilon,loc}(\Omega)\cap L^{\infty}_{loc}(\Omega)$ such that $w=u_{p}^{\varepsilon}$ a.e. in $\Omega\setminus K$ and $\varphi=w-u$ equation (7.3) yields
$\int_{K}|\nabla_{\varepsilon}u|_{\varepsilon}^{p}(u-w)=\int_{K}|\nabla_{\varepsilon}u|_{\varepsilon}^{p-2}\langle\nabla_{\varepsilon}u,\nabla_{\varepsilon}w\rangle_{\varepsilon}-\int_{K}|\nabla_{\varepsilon}u|_{\varepsilon}^{p}.$
We now estimate from above the first term on the right hand side, first by using Cauchy-Schwarz inequality and then Young inequality (with conjugate exponents $\tfrac{p}{p-1}$ and $p$), getting
(7.4) $\int_{K}|\nabla_{\varepsilon}u|_{\varepsilon}^{p}(u-w)\leq\left(1-\frac{1}{p}\right)\int_{K}|\nabla_{\varepsilon}u|_{\varepsilon}^{p}+\int_{K}\dfrac{|\nabla_{\varepsilon}w|_{\varepsilon}^{p}}{p}-\int_{K}|\nabla_{\varepsilon}u|_{\varepsilon}^{p}=\int_{K}\dfrac{|\nabla_{\varepsilon}w|_{\varepsilon}^{p}}{p}-\int_{K}\dfrac{|\nabla_{\varepsilon}u|_{\varepsilon}^{p}}{p}.$
Finally, a simple rearrangement of the latter gives the desired conclusion.
Let us now prove that $(ii)\Rightarrow(i)$. For every compact set $K\subset\Omega$, we can fix any $\varphi\in W^{1,p}_{\varepsilon}(\Omega)\cap L^{\infty}(\Omega)$ such that $\varphi=0$ a.e. in $\Omega\setminus K$ and for any $t\in\mathbb{R}$ the function $u+t\varphi$ is now an admissible competitor for the minimality property of the functional $J^{p,\varepsilon}_{u}(\cdot,K)$. Therefore, differentiating under integral sign gives
$0=\dfrac{d}{dt}\Big{|}_{t=0}J^{p,\varepsilon}_{u}(u+t\varphi,K)=\int_{K}|\nabla_{\varepsilon}u|_{\varepsilon}^{p-2}\langle\nabla_{\varepsilon}u,\nabla_{\varepsilon}\varphi\rangle_{\varepsilon}+\int_{K}|\nabla_{\varepsilon}u|_{\varepsilon}^{p}\varphi,$
and this concludes the proof. ∎
The last auxiliary result describes how the minimality property for the energy functional (1.7) stated in the previous lemma passes to the limit.
Proposition 7.4.
Let $1<p<Q$, $\varepsilon\in[0,1]$, and $\Omega\subset\mathbb{H}^{n}$ be an open set with bounded complement and smooth boundary $\partial\Omega$. For $v^{\varepsilon}_{p}:\mathbb{H}^{n}\to(0,1]$ the minimizer of (1.10) in the set $\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ let $u^{\varepsilon}_{p}:\mathbb{H}^{n}\to[0,\infty)$ the corresponding functions given by (1.8), so that they minimize the functionals $J^{p,\varepsilon}_{u^{\varepsilon}_{p}}$ defined in (1.7) in the sense that for every compact set $K\subset\Omega$ it holds that
(7.5) $J_{u_{p}^{\varepsilon}}^{p,\varepsilon}(u_{p}^{\varepsilon};K)\leq J_{u_{p}^{\varepsilon}}^{p,\varepsilon}(w;K),$
for every locally $d_{\varepsilon}$-Lipschitz function $w:\Omega\to\mathbb{R}$ such that $w\equiv u_{p}^{\varepsilon}$ in $\Omega\setminus K$.
Assume that $\{u^{\varepsilon}_{p}\}$ satisfy the uniform bound (1.13), so that, up to subsequences, $u_{p}^{\varepsilon}\to u$ as $(\varepsilon,p)\to(0,1)$ locally uniformly in $\mathbb{H}^{n}$, where $u$ is a $d_{0}$-Lipschitz function. Then $u:\mathbb{H}^{n}\to[0,\infty)$ is a weak solution of (1.1), i.e., for every compact set $K\subset\Omega$, it holds
(7.6) $J^{0}_{u}(u;K)\leq J^{0}_{u}(w;K),$
for every locally $d_{0}$-Lipschitz function $w:\Omega\to\mathbb{R}$ such that $w\equiv u$ in $\Omega\setminus K$. Analogous conclusion holds for $\varepsilon\equiv const$ (resp., for $p\equiv 1$) as $p\to 1$ (resp., as $\varepsilon\to 0$).
Proof.
We give the prooof only in the case $(\varepsilon,p)\to(0,1)$, as the two others are similar and even simpler, adapting the argument in [34] for the Euclidean case (see also [15] for the case $\mathbb{H}^{1}$ when $\varepsilon=0$).
We consider a nonnegative test function $\eta\in C^{\infty}_{0}(\Omega)$ such that $\rm{supp}(\eta)\subset\subset\Omega$ satisfying the normalization condition $0\leq\eta\leq 1$ in $\Omega$, together with a bounded open set $V\subset\subset\Omega$ such that $\rm{supp}(\eta)\subset\subset V$. Then, for a compact set $K\subset\Omega$ we fix a locally $d_{0}$-Lipschitz function $w:\Omega\to\mathbb{R}$ such that $w\equiv u$ in $\Omega\setminus K$.
Note that, according to [14, Theorem 2.15], there exists a sequence of smooth functions $(w_{k})_{k\in\mathbb{N}}\subset C^{\infty}(V)$ such that
  • i)
    $w_{k}\to w$ uniformly on compact subsets of $V$;
  • ii)
    $\|\nabla_{0}w_{k}\|_{L^{\infty}(W)}$ is bounded for each $W\subset\subset V$ and $k$ sufficiently large;
  • iii)
    $\nabla_{0}w_{k}\to\nabla_{0}w$ a.e. in $V$.
Thus, for each fixed $k$ we use the function $\eta w_{k}+(1-\eta)u_{p}^{\varepsilon}$ as test function in (7.5) for the functional $J^{p,\varepsilon}_{u^{\varepsilon}_{p}}(\cdot;\rm{supp}(\eta))$ and we find that
$\int_{\rm{supp}(\eta)}$ $\left(\dfrac{1}{p}|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}+u_{p}^{\varepsilon}|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}\right)$
$\leq\int_{\rm{supp}(\eta)}\left(\dfrac{1}{p}|\nabla_{\varepsilon}(\eta w_{k}+(1-\eta)u_{p}^{\varepsilon})|_{\varepsilon}^{p}+(\eta w_{k}+(1-\eta)u_{p}^{\varepsilon})|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}\right),$
which readily yields
(7.7) $\int_{\rm{supp}(\eta)}$ $\left(\dfrac{1}{p}|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}+\eta(u_{p}^{\varepsilon}-w_{k})|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}\right)$
$\leq\dfrac{1}{p}\int_{\rm{supp}(\eta)}\left|\eta\nabla_{\varepsilon}w_{k}+(1-\eta)\nabla_{\varepsilon}u_{p}^{\varepsilon}+(w_{k}-u_{p}^{\varepsilon})\nabla_{\varepsilon}\eta\right|_{\varepsilon}^{p}$
$\leq\dfrac{3^{p-1}}{p}\int_{\rm{supp}(\eta)}\left(\eta^{p}|\nabla_{\varepsilon}w_{k}|_{\varepsilon}^{p}+(1-\eta)^{p}|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}+|w_{k}-u_{p}^{\varepsilon}|^{p}|\nabla_{\varepsilon}\eta|_{\varepsilon}^{p}\right).$
Since $0\leq 1-\eta\leq 1$ we can rewrite (7.7) as
$\int_{\rm{supp}(\eta)}$ $\left(\dfrac{\eta+(1-\eta)}{p}|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}+\eta(u_{p}^{\varepsilon}-w_{k})|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}\right)$
$\leq\dfrac{3^{p-1}}{p}\int_{\rm{supp}(\eta)}\left(\eta^{p}|\nabla_{\varepsilon}w_{k}|_{\varepsilon}^{p}+(1-\eta)|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}+|w_{k}-u_{p}^{\varepsilon}|^{p}|\nabla_{\varepsilon}\eta|_{\varepsilon}^{p}\right),$
which in turn is equivalent to
(7.8) $\int_{\rm{supp}(\eta)}$ $\left(\dfrac{1}{p}\eta|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}+\eta(u_{p}^{\varepsilon}-w_{k})|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}\right)$
$\leq\dfrac{3^{p-1}}{p}\int_{\rm{supp}(\eta)}\left(\eta^{p}|\nabla_{\varepsilon}w_{k}|_{\varepsilon}^{p}+|w_{k}-u_{p}^{\varepsilon}|^{p}|\nabla_{\varepsilon}\eta|_{\varepsilon}^{p}\right)$
$\qquad+\dfrac{3^{p-1}-1}{p}\int_{\rm{supp}(\eta)}(1-\eta)|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}.$
Now we estimate the right hand side of (7.8). Since $0\leq\eta\leq 1$ and each $w_{k}$ is smooth on $supp(\eta)$ we obtain
(7.9) $\dfrac{3^{p-1}}{p}\int_{\rm{supp}(\eta)}\eta^{p}|\nabla_{\varepsilon}w_{k}|_{\varepsilon}^{p}\leq\dfrac{3^{p-1}}{p}\int_{\rm{supp}(\eta)}\eta|\nabla_{\varepsilon}w_{k}|_{\varepsilon}^{p}\longrightarrow\int_{\rm{supp}(\eta)}\eta|\nabla_{0}w_{k}|_{0},\quad\textrm{as }(\varepsilon,p)\to(0,1).$
Concerning the second term in the right hand side of (7.8), we have $|w_{k}-u_{p}^{\varepsilon}|^{p}\to|w_{k}-u|$ and $|\nabla_{\varepsilon}\eta|_{\varepsilon}^{p}\to|\nabla_{0}\eta|_{0}$ uniformly on $supp(\eta)$ as $(\varepsilon,p)\to(0,1)$, hence
(7.10) $\dfrac{3^{p-1}}{p}\int_{\rm{supp}(\eta)}|w_{k}-u_{p}^{\varepsilon}|^{p}|\nabla_{\varepsilon}\eta|_{\varepsilon}^{p}\longrightarrow\int_{\rm{supp}(\eta)}|w_{k}-u||\nabla_{0}\eta|_{0}\quad\textrm{as }(\varepsilon,p)\to(0,1).$
Finally, since $\tfrac{3^{p-1}-1}{p}\to 0$ as $p\to 1$ then for the last term of (7.8) it follows from (1.13) that
(7.11) $\dfrac{3^{p-1}-1}{p}\int_{\rm{supp}(\eta)}(1-\eta)|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}\to 0,\quad\textrm{as }(\varepsilon,p)\to(0,1),$
Concerning the left hand side of (7.8), chosing $w=u$ and combining (1.13), the uniform convergence $u_{p}^{\varepsilon}\to u$ and (i) we infer that
[…]
whence (7.8)-(7.11) together with (ii) and (iii) and dominated convergence yield
(7.12) $\limsup_{(\varepsilon,p)\to(0,1)}\int_{\rm{supp}(\eta)}\dfrac{1}{p}\eta|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}$ $\leq\limsup_{k}\limsup_{(\varepsilon,p)\to(0,1)}\int_{\rm{supp}(\eta)}\left(\dfrac{1}{p}|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}+\eta(u_{p}^{\varepsilon}-w_{k})|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}\right)$
$\leq\limsup_{k\to\infty}\int_{\rm{supp}(\eta)}\eta|\nabla_{0}w_{k}|_{0}+|w_{k}-u||\nabla_{0}\eta|_{0}=\int_{\rm{supp}(\eta)}\eta|\nabla_{0}u|_{0}\,.$
On the other hand, note that $\nabla_{0}u_{p}^{\varepsilon}\to\nabla_{0}u$-weakly in $L^{\infty}(\mathbb{H}^{n})$ because (1.13) holds uniformly and $u^{\varepsilon}_{p}\to u$ locally uniformly, whence weak lower semicontinuity and Young’s inequality give
$\int_{\rm{supp}(\eta)}\eta|\nabla_{0}u|_{0}\leq\liminf_{(\varepsilon,p)\to(0,1)}\int_{\rm{supp}(\eta)}\eta|\nabla_{0}u_{p}^{\varepsilon}|_{0}\leq\liminf_{(\varepsilon,p)\to(0,1)}\int_{\rm{supp}(\eta)}\dfrac{1}{p}\eta|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}\,,$
because $\frac{p-1}{p}\int_{\rm{supp}(\eta)}\eta\to 0$ as $p\to 1$.
Combining the last inequality with (7.12) we conclude that
(7.13) $\int_{\rm{supp}(\eta)}\dfrac{1}{p}\eta|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}\to\int_{\rm{supp}(\eta)}\eta|\nabla_{0}u|_{0},\quad\textrm{as }(\varepsilon,p)\to(0,1),$
for any nonnegative test function $\eta\in C^{\infty}_{0}(\Omega)$ such that $\rm{supp}(\eta)\subset\subset\Omega$ and such that $0\leq\eta\leq 1$, whence the same holds if $\eta$ has the same properties but it is merely continuous, by approximation and (1.13). In turn, (7.13) holds for any test function $\eta\in C^{\infty}_{0}(\Omega)$ such that $\rm{supp}(\eta)\subset\subset\Omega$, by decomposition in positive and negative part and normalization.
Now for $w\neq u$ and $K$ as above we fix $\eta\in C^{\infty}_{0}(\Omega)$ such that $\rm{supp}(\eta)\subset\subset\Omega$, $0\leq\eta\leq 1$ and $\eta\equiv 1$ in an open neighborhood of $K$. Once more we take a bounded open set $V\subset\subset\Omega$ such that $\rm{supp}(\eta)\subset\subset V$ and a sequence $\{w_{k}\}$ satisfying properties (i)-(ii)-(iii) above. It follows from (7.13) that
$\int_{\rm{supp}(\eta)}\eta(u_{p}^{\varepsilon}-w_{k})|\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}\to\int_{\rm{supp}(\eta)}\eta(u-w_{k})|\nabla_{0}u|_{0}\quad\textrm{as }(\varepsilon,p)\to(0,1),$
where for fixed $k$ we used the uniform convergence of $\eta(u_{p}^{\varepsilon}-w_{k})$ to $\eta(u-w_{k})$ as a test function, so that (7.13) together with (7.8)-(7.11) imply
$\int_{\rm{supp}(\eta)}\eta|\nabla_{0}u|_{0}+\eta(u-w_{k})|\nabla_{0}u|_{0}\leq\int_{\rm{supp}(\eta)}\eta|\nabla_{0}w_{k}|_{0}+|w_{k}-u||\nabla_{0}\eta|_{0}.$
Finally, taking into account (i), (ii), and (iii), and passing to the limit as $k\to+\infty$ on both sides by dominated convergence, we get
$\int_{\rm{supp}(\eta)}\eta|\nabla_{0}u|_{0}+\eta(u-w)|\nabla_{0}u|_{0}\leq\int_{\rm{supp}(\eta)}\eta|\nabla_{0}w|_{0}+|w-u||\nabla_{0}\eta|_{0}\,,$
hence $\eta\equiv 1$ near $K$ and $w\equiv u$ in $\Omega\setminus K$ yield
$\int_{K}\left(|\nabla_{0}u|_{0}+u|\nabla_{0}u|_{0}\right)\leq\int_{K}\left(|\nabla_{0}w|_{0}+w|\nabla_{0}u|_{0}\right)\,,$
i.e., (7.6) holds and the proof is complete. ∎
Finally we are in the position to prove the main result of the paper.
Proof of Theorem 1.4.
Let us fix $\varepsilon\in[0,1]$ and for any $1<p<Q$ let $v^{\varepsilon}_{p}$ be the minimizer of (1.10) in the set $\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ and let $u^{\varepsilon}_{p}$ the corresponding functions given by (1.8). According to Theorem 1.2 we know that (1.12) holds for $\{v^{\varepsilon}_{p}\}$ for a constant $C>0$ independent of $p$ as well as of $\varepsilon$. Hence, in view of Corollary 1.3 functions in $\{u^{\varepsilon}_{p}\}$ satisfy the uniform bound (1.13), so that $\{u^{\varepsilon}_{p}\}_{p\in(1,Q)}$ are equi-Lipschitz with respect to $d_{\varepsilon}$, strictly positive in $\Omega$ and vanishing identically on $\mathbb{H}^{n}\setminus\Omega$. Moreover, they are weak solutions to (1.6a) due to Lemma 7.2, so that they minimize the functionals $J^{p,\varepsilon}_{u^{\varepsilon}_{p}}$ defined in (1.7) with respect to competitors $w\in\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ such that $w-u^{\varepsilon}_{p}$ is compactly supported because of Lemma 7.3.
According to Proposition 7.4, up to subsequences, for each $\varepsilon\in[0,1]$ there exist functions $u^{\varepsilon}:\mathbb{H}^{n}\to[0,\infty)$ which are $d_{\varepsilon}$-Lipschitz on $\mathbb{H}^{n}$, with bound on Lipschitz constant independent on $\varepsilon\in[0,1]$, such that $u_{p}^{\varepsilon}\to u^{\varepsilon}$ as $p\to 1$ when $\varepsilon>0$ or $u_{p}^{\bar{\varepsilon}}\to u^{0}$ as $(\bar{\varepsilon},p)\to(0,1)$ if $\varepsilon=0$ locally uniformly in $\mathbb{H}^{n}$. Alternatively, for $\varepsilon=0$ we can let first $\bar{\varepsilon}\to 0$ and then $p\to 1$. In addition, $\{u^{\varepsilon}\}_{\varepsilon\in[0,1]}$ vanish identically on $\mathbb{H}^{n}\setminus\Omega$ and they are weak solution of (1.1) and (1.3) in the sense of Definition 1.1.
In order to conclude the proof it is enough to establish (1.14) and we treat first the case $\varepsilon\equiv 0$ because it is much simpler. Indeed, combing weak comparison principle for sub-Riemannian $p$-harmonic function applied on the open sets $\Omega$ and $\mathbb{H}^{n}\setminus\overline{B_{\bar{R}}(g_{0})}$ with the explicit form of $p$-capacitary potentials from Lemma 3.1 we have
(7.14) $\left(\frac{\|g_{0}^{-1}*g\|}{R_{0}}\right)^{-\gamma}\leq v^{0}_{p}(g)\leq\left(\frac{\|g_{0}^{-1}*g\|}{\bar{R}}\right)^{-\gamma}\quad\hbox{on }\quad\overline{\Omega}\,,\qquad\hbox{with}\quad\gamma=\frac{Q-p}{p-1}\,,$
where the second inequality trivially holds in $\overline{\Omega}\cap B_{\bar{R}}(g_{0})$ because $v^{0}_{p}\leq 1$ in $\mathbb{H}^{n}$. Taking logarithm and multiplying by $01-p<$ 0, as $p\to 1$ inequality (1.14) follows in the form stated in the case $\varepsilon=0$.
As for the case $\varepsilon>0$, concerning the upper bound for each $u^{\varepsilon}=\lim u^{\varepsilon}_{p}$ we rely on Proposition 5.1. Thus, for $\varepsilon\in(0,1]$ and $1<p\leq 2$, since $\Omega\subset\mathbb{H}^{n}$ satisfies assumption ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}$ and $B_{R_{0}}(g_{0})\subset\Omega^{c}$ then for each $\bar{R}>2>R_{0}$ such that $\mathbb{H}^{n}\setminus\Omega\subset B_{\bar{R}}(g_{0})$ there exists $\widehat{C}>1$ depending only on $R_{0}$ and $\bar{R}$ such that for $v^{\varepsilon}_{p}$, $u^{\varepsilon}_{p}=(1-p)\log v^{\varepsilon}_{p}$, we have
(7.15) $v^{\varepsilon}_{p}(g)\geq\widehat{C}^{-\frac{\varepsilon^{4}}{p-1}}\left(\frac{\|g_{0}^{-1}*g\|}{R_{0}}\right)^{-\gamma}\qquad\hbox{on }\quad\overline{\Omega}\,,\qquad\hbox{with}\quad\gamma=\frac{Q-p}{p-1}\,.$
Thus, taking logarithm of both sides in (7.15) as $p\to 1$ the upper bound follows because $u^{\varepsilon}=\lim_{p}u^{\varepsilon}_{p}$.
When dealing with the lower bound in (1.14) for $\varepsilon\in(0,1]$ we rely instead on Proposition 5.5 so that inequality (5.11) holds in $\overline{\Omega}$ for some $C_{0}>1$ depending only on $R_{0}$, $\bar{R}$, $\partial\Omega$. Taking logarithm of both sides for $\gamma=\frac{Q-p}{p-1}$ and $g\in\overline{\Omega}$ we obtain
$u^{\varepsilon}(g)=\lim_{p\to 1}u^{\varepsilon}_{p}(g)=-\lim_{p\to 1}\log\left(v^{\varepsilon}_{p}(g)\right)^{p-1}$
$\geq-\lim_{p\to 1}\log\left({C_{0}}\left(\frac{\|g_{0}^{-1}*g\|}{\bar{R}}\right)^{p-Q}\right)=(Q-1)\log\frac{\|g_{0}^{-1}*g\|}{\bar{R}}-\log C_{0}\,,$
which completes the proof of the lower bound. Note that the previous argument still applies in the case of the joint limit $(\varepsilon,p)\to(0,1)$, hence inequality (5.11) holds no matter how $u=\lim_{\varepsilon,p}u^{\varepsilon}_{p}$ is constructed.